### v1

# Credal Prediction based on Relative Likelihood

Abstract
Predictions in the form of sets of probability distributions, so-called credal sets, provide a suitable means to represent a learner’s epistemic uncertainty. In this paper, we propose a theoretically grounded approach to credal prediction based on the statistical notion of relative likelihood: The target of prediction is the set of all (conditional) probability distributions produced by the collection of plausible models, namely those models whose relative likelihood exceeds a specified threshold. This threshold has an intuitive interpretation and allows for controlling the trade-off between correctness and precision of credal predictions. We tackle the problem of approximating credal sets defined in this way by means of suitably modified ensemble learning techniques. To validate our approach, we illustrate its effectiveness by experiments on benchmark datasets demonstrating superior uncertainty representation without compromising predictive performance. We also compare our method against several state-of-the-art baselines in credal prediction.
1 Introduction
The distinction between two types of uncertainty, referred to as aleatoric and epistemic, is receiving increasing interest in machine learning (Hüllermeier and Waegeman, 2021). Roughly speaking, the aleatoric uncertainty of a predictive model is caused by the inherent randomness of the data-generating process, whereas epistemic uncertainty is caused by the learner’s lack of knowledge about the true (or best) predictive model. While aleatoric uncertainty is irreducible, epistemic uncertainty can be reduced on the basis of additional information, e.g., by collecting more training data.
Aleatoric uncertainty can be captured adequately in terms of (conditional) probability distributions, whereas the representation of epistemic uncertainty requires “second-order” formalisms more general than (single) probability distributions. In the Bayesian approach to machine learning, second-order probability distributions are used for this purpose (Depeweg et al., 2018; Kendall and Gal, 2017): The learner maintains a probability distribution over the model space, which, if each model is a probabilistic predictor, results in predictions in the form of probability distributions of probability distributions (on outcomes). The Bayesian approach is theoretically appealing but computationally demanding, and commonly criticized for the need to specify a prior distribution (strongly influencing prediction and uncertainty quantification) (Gawlikowski et al., 2023).
Figure 1: Overview of our method. Language inference example from the ChaosNLI dataset, showing a premise-hypothesis pair with three possible labels: Entailed, Neutral and Contradicted. Middle: Illustration of the learning mechanism of our Credal Relative Likelihood (CreRLα) framework, showing ensemble member training up to different $\alpha$ thresholds. Right: Unlike baselines, which are centered around the MLE and remain fixed, our method allows to adapt credal set size to improve coverage of the ground-truth distribution.
An alternative second-order formalism is offered by credal sets (Walley, 1991), i.e., sets (instead of distributions) of probability distributions that are commonly assumed to be closed and convex (Cozman, 2000). Although a set appears to provide weaker information than a distribution, a set-based representation also has advantages. In particular, it can be argued that sets are more apt at representing ignorance in the sense of a lack of knowledge (Dubois et al., 1996), because distributions always involve additional assumptions beyond the mere distinction between plausible and implausible candidate models. Decision-making on the basis of credal sets has been proposed as a reasonable alternative to Bayesian decision-making (Levi, 1978; Girón and Ríos, 1980). The learning of credal predictors (i.e., models making predictions in the form of credal sets) has been considered in machine learning in the past (Zaffalon, 2001; Corani and Zaffalon, 2008) and, in light of the quest for proper representation of epistemic uncertainty, received renewed interest more recently (Wang et al., 2024a; Nguyen et al., 2025; Cella and Martin, 2024).
A major question to be addressed in this regard concerns the distinction between plausible and implausible models. In this paper, we adopt a theoretically grounded approach for constructing credal sets based on the notion of relative likelihood. Our approach explicitly defines the prediction target as the set of all probability distributions generated by plausible learners, namely those whose relative likelihood exceeds a specified threshold (cf. Section 3). That said, realizing this in complex machine learning settings comes with several practical challenges, notably the questions of representation, approximation, and inference: How to formally represent subsets of a high-dimensional model space, how to approximate such sets algorithmically, and how to infer credal predictions from them? We propose methods to address these questions based on suitably modified ensemble techniques (cf. Section 4). We highlight the effectiveness of our approach on real-world datasets, evaluating its performance in terms of coverage and efficiency as key criteria of credal prediction. Also, we analyze its performance on downstream tasks such as Out-of-Distribution (OoD) detection (cf. Section 5).
Contributions. In summary our contributions are as follows:
  • 1.
    We develop a theoretically sound approach to learning credal predictors, which is grounded in the statistical notion of relative likelihood.
  • 2.
    We cast the learning task as optimizing multi-objective generalization performance, namely as finding a compromise between coverage and efficiency of credal predictions.
  • 3.
    We propose an adaptive and conceptually intuitive ensemble-based method for approximating sets of plausible models and the induced credal predictions.
  • 4.
    We empirically compare our method to representative baselines and, for the first time, compare state-of-the-art credal predictors by coverage and efficiency. Our approach achieves superior uncertainty representation and strong OoD performance.
Related Work. In machine learning, uncertainty is often represented by (approximate) Bayesian methods (Mackay, 1992; Blundell et al., 2015; Lakshminarayanan et al., 2017; Gal and Ghahramani, 2016; Daxberger et al., 2021). An important characteristic of such representations, especially for uncertainty tasks, is diversity (D’Angelo and Fortuin, 2021; Wood et al., 2023), which can be enforced by regularization (de Mathelin et al., 2023), varying hyper-parameters (Wenzel et al., 2020), or increasing diversity in representations (Lopes et al., 2022). Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty (Zaffalon, 2001; Corani and Zaffalon, 2008; Corani and Mignatti, 2015). Such sets can be generated based on the relative likelihood, also referred to as normalized likelihood (Antonucci et al., 2012a) and have been used in machine learning with simple model classes (Senge et al., 2014; Cella and Martin, 2024). In this work, we build upon these approaches and address the challenges that emerge when adapting the relative likelihood to a setting with complex predictors. Recently, credal sets have also been applied in the context of deep learning. Some approaches use ensemble learning to derive class-wise lower and upper probabilities (Wang et al., 2024b; Nguyen et al., 2025). Others train models to directly predict probability intervals (Wang et al., 2024a). Hybrid methods, combining multiple uncertainty frameworks, have also been proposed. Caprio et al. (2023) combine Bayesian deep learning and credal sets by considering sets of priors. Another approach leverages conformal prediction to construct credal sets with validity guarantees (Javanmardi et al., 2024). A more detailed discussion of the related work is provided in Appendix A.
2 Problem Statement
We consider classification in a supervised learning setting. We have access to training data $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the instance space and $\mathcal{Y}=\{y_{1},\dots,y_{K}\}$ is the label space with $K\in\mathbb{N}$ classes. The training data are realizations of random variables that are independently and identically distributed according to a probability measure on $\mathcal{X}\times\mathcal{Y}$. We consider a hypothesis space $\mathcal{H}$ with probabilistic predictors $h:\mathcal{X}\to\Delta_{K}$, where $\Delta_{K}$ denotes the $(K-1)$-simplex (that we will also refer to as the probability simplex). Given an instance $\bm{x}$, a predictor $h$ assigns a probability distribution $p(\cdot\,|\,\bm{x},h)=h(\bm{x})$, which is an estimate of the ground-truth probability $p(\cdot\,|\,\bm{x},h^{*})$ generated by the ground-truth model $h^{*}$.
Predictive models $h$ are often evaluated in terms of their likelihood $L(h)=\prod_{i=1}^{N}p(y_{i}\,|\,\bm{x}_{i},h)$. Then, adopting the established principle of maximum likelihood inference, the model of choice is the maximum likelihood estimator (MLE) $h^{ML}$, i.e., the model whose likelihood is highest. By predicting probability distributions $h^{ML}(\bm{x})$ on $\mathcal{Y}$, this model is able to represent aleatoric uncertainty. However, it cannot capture information about the uncertainty of $h^{ML}$ itself, i.e., about how much it possibly deviates from the ground-truth model $h^{*}$.
In order to capture this epistemic uncertainty, we consider a second-order uncertainty representation in the form of credal sets: Instead of relying on a single predictor $h$, the uncertainty about the true underlying model $h^{*}$ is represented by the set of plausible predictors $\mathcal{C}\subseteq\mathcal{H}$. For a given query instance $\bm{x}\in\mathcal{X}$, this set induces a prediction in the form of a credal set of distributions:
$\mathcal{Q}_{\bm{x}}=\{p(\cdot\,|\,\bm{x},h):h\in\mathcal{C}\}\subseteq\Delta_{K}.$ (1)
We define the problem of learning a credal predictor as a generalization of the standard setting of supervised learning as introduced above: Given training data $\mathcal{D}$, the task is to induce a model $H:\mathcal{X}\rightarrow 2^{\Delta_{K}}$ that delivers predictions in the form of credal sets $\mathcal{Q}_{\bm{x}}=H(\bm{x})\subseteq\Delta_{K}$. In our setup, $H(\bm{x})$ represents the mapping of an instance $\bm{x}\in\mathcal{X}$ through a set of plausible models $\mathcal{C}$ to the corresponding credal set $\mathcal{Q}_{\bm{x}}$. In general, $H$ could also be realized differently. Inspired by other set-valued prediction methods such as conformal prediction (Vovk et al., 2005), we evaluate such a predictor in terms of its coverage and efficiency. Coverage (of the ground-truth distribution by the cedal set) is defined as follows:
$C(H)=\mathbb{E}\left[\llbracket p(\cdot\,|\,\bm{x},h^{*})\in H(\bm{x})\rrbracket\right],$ (2)
[…]
$E(H)=1-\mathbb{E}\left[\frac{1}{K}\sum_{k=1}^{K}\overline{p}(y_{k}\,|\,\bm{x})-\underline{p}(y_{k}\,|\,\bm{x})\right],$ (3)
where $\overline{p}(y_{k}\,|\,\bm{x})=\sup\{p(y_{k}\,|\,\bm{x}):p\in H(\bm{x})\}$ (and $\underline{p}(y_{k}\,|\,\bm{x})$ is defined analogously). We assume efficiency to be positively oriented, meaning that higher efficiency corresponds to smaller credal sets. As opposed to the volume of a credal set, which, besides being challenging to compute, is not intuitive in high dimensions, this measure of efficiency is particularly interpretable. Specifically, it describes the (complement of the) average interval length for each class. Importantly, we use efficiency strictly as an indicator of set size, not as a measure of epistemic uncertainty (cf. Section 2).
The empirical coverage and efficiency is determined in terms of the corresponding averages on a finite set of (test) data $\bm{x}_{1},\ldots,\bm{x}_{N}$. Note that ground-truth (first-order) distributions $p(\cdot\,|\,\bm{x}_{i},h^{*})$ are typically unavailable during both training and testing, which often necessitates approximating coverage through alternative means. In this work, however, we make use of data that provides access to ground-truth distributions.
Evaluating a credal predictor in terms of its coverage and efficiency means that two predictors $H$ and $H^{\prime}$ are not necessarily comparable in the sense that one of them is “better” than the other one. Instead, such predictors are only comparable in a Pareto sense: $H$ is better than $H^{\prime}$ if $C(H)\geq C(H^{\prime})$ and $E(H)\geq E(H^{\prime})$ (and one of the inequalities is strict). The task can then be specified as learning Pareto-optimal credal predictors.
3 Relative Likelihood-Based Credal Sets
To learn a credal predictor $H:\mathcal{X}\rightarrow 2^{\Delta_{K}}$ from training data, we adopt the representation given in (1), defining $\mathcal{C}$ as a set of plausible (first-order) predictors $h:\mathcal{X}\rightarrow\Delta_{K}$ drawn from an underlying hypothesis space $\mathcal{H}$. To characterize plausibility, we employ the concept of relative likelihood:
$\gamma(h)=\frac{L(h)}{L(h^{ML})}=\frac{L(h)}{\sup\limits_{h^{\prime}\in\mathcal{H}}L(h^{\prime})}.$
Here, $h^{ML}$ represents the model in $\mathcal{H}$ with the highest likelihood given the training data $\mathcal{D}$. This notion, also called extended or normalized likelihood, has been proposed by Birnbaum (1962) and used for statistical inference (Wasserman, 1990; Walley and Moral, 1999). It offers an attractive alternative to Bayesian or frequentist reasoning eliminating the need to specify priors or perform resampling (Giang and Shenoy, 2002). In machine learning this is especially useful as specifying a meaningful prior on $\mathcal{H}$ is a non-trivial problem and resampling from the data distribution is usually not possible either, as we only have access to the given samples $\mathcal{D}$. An axiomatic justification of the relative likelihood (in the context of evidence theory) has been given by Denoeux (2014), who derives it from (i) the likelihood principle, (ii) compatibility with the Bayes rule, and (iii) the minimal commitment principle.
On the basis of the relative likelihood, a set of models can be constructed, analogous to a confidence region (Aitkin, 1982), by including models that are plausible in the sense of surpassing a threshold $\alpha\in[0,1]$. This set of models is also referred to as an $\alpha$-cut (Antonucci et al., 2012a):
Figure 2: Relative Likelihood curve.
Based on 10 samples from a Bernoulli distribution with parameter 0.5. The credal set of $\alpha=0.9$ around the MLE at $0.6$ does not cover the ground-truth. Lowering $\alpha$ increases coverage at the cost of efficiency.
$\mathcal{C}_{\alpha}=\{h:\gamma(h)\geq\alpha\}\subseteq\mathcal{H}.$ (4)
According to this definition, a model $h$ is considered implausible (and hence ignored) if its likelihood is too small compared to the likelihood of the (presumably) best model, namely if its likelihood is less than $\alpha$ times the likelihood of the best model. Thus, the parameter $\alpha$ has a quite intuitive meaning. The concept of $\alpha$-cuts is further illustrated in Figure 2.
Given an instance $\bm{x}\in\mathcal{X}$, the set of predictors maps to a set of (conditional) distributions:
$\mathcal{Q}_{\bm{x},\alpha}=\{p(\cdot\,|\,\bm{x},h):h\in\mathcal{C}_{\alpha}\}\subseteq\Delta_{K}.$
This credal set can then be used for predictive tasks and to quantify uncertainty. In essence, it can be used to express the amount of uncertainty one has about the best predictor that can be obtained from the given data sample.
Considering that we aim to simultaneously optimize coverage and efficiency, it is important to highlight the role of the parameter $\alpha$ as a means to move along the Pareto front and trade coverage against efficiency: The lower $\alpha$, the higher coverage tends to be. At the same time, however, efficiency will deteriorate (cf. again Figure 2). As $\alpha$ increases, the intervals for given data points shrink since only models with high relative likelihood for the given data sample are included. Conversely, lowering $\alpha$ allows for the inclusion of models with lower likelihood, resulting in larger credal intervals. This controllability induced by $\alpha$ emphasizes the appeal of the relative likelihood as a tool for (plausible) model selection and, thereby, credal set construction.
4 Approximating Relative Likelihood-Based Credal Sets
Building on the theoretical framework for constructing credal sets using relative likelihood, it becomes evident that, in practice, we must approximate the set of models $\mathcal{C}_{\alpha}$, and consequently the resulting credal set $\mathcal{Q}_{\bm{x},\alpha}$. In this section, we examine how the formal notion of relative likelihood can be applied to construct credal sets in high-dimensional machine learning settings. Up to this point, we made no assumptions about the specific model class. Obviously, to approximate the $\alpha$-cut (4), the model class must allow to define and optimize toward a target relative likelihood during training. As a concrete and widely used example of complex models in machine learning, we focus on neural networks throughout the remainder of this work.
Estimating the Maximum Likelihood Estimator. Several strategies can be employed to obtain $h^{ML}$, including training a model with parameters known to perform well, utilizing a pre-trained model, or leveraging AutoML techniques to determine optimal parameter settings during training. Throughout this work, we adopt the first approach and train the maximum likelihood estimator $h^{ML}$ using parameter configurations that have been identified in the literature as effective. Further details and configurations of experiments are provided in Appendix B.
Given an estimated maximum likelihood model, the estimated relative likelihood is obtained as $\hat{\gamma}(h)=\frac{L(h)}{L(h^{ML})}.$
Algorithm 1 Train Credal Relative Likelihood ensemble.
0:  $\alpha,M$
1:  Step 1: Approximate maximum likelihood model
2:  $h^{ML}\leftarrow\arg\min_{h}\mathcal{L}(h)$
3:  Step 2: Train ensemble members
4:  $\Delta\tau=\frac{1-\alpha}{M}$
5:  for $i\in\{0,\ldots,M-1\}$ do
6:     $\tau_{i}=\alpha+i\cdot\Delta\tau$
7:     $h_{i}\leftarrow$ ToBias initialization
8:     Train $h_{i}$ such that $\hat{\gamma}(h_{i})\approx\tau_{i}$
9:  end for
Approximating the $\bm{\alpha}$-cut. Having found a suitable candidate model to compute the maximum likelihood, the next step is to construct the $\alpha$-cut. A straightforward approach is to train an ensemble of hypotheses and construct the $\alpha$-cut based on these hypotheses. In essence, approaches based on regular ensemble training (Wang et al., 2024b; Nguyen et al., 2025) form an example of this strategy. However, an issue is that all hypotheses $h$ typically tend to cluster around $h^{ML}$, hence the obtained hypotheses do not accurately approximate the $\alpha$-cut (unless $\alpha\approx 1$, see Figure 1).
To obtain better coverage of $\mathcal{C}_{\alpha}$, we propose to train $M$ hypotheses using an early stopping strategy, namely until a specific relative likelihood value $\tau_{i}$ is reached. Given $\alpha$, we consider thresholds $\tau=\{\tau_{i}\,|\,\tau_{i}=\alpha+i\cdot\Delta\tau\}$, where $\Delta\tau=\frac{1-\alpha}{M-1}$ and $i\in\{0,\ldots,M-2\}$. This guarantees a broad range of hypotheses in terms of relative likelihoods. The influence of the ensemble size $M$ is analyzed in an ablation study in Section C.2.
Although this should cover the $\alpha$-cut well, there might still be a lack of diversity in the predictions of the resulting hypotheses. In the literature, it is known that training ensembles without explicitly encouraging diversity actually minimizes diversity (Abe et al., 2022). To encourage diversity, we introduce a novel initialization strategy called ToBias. The idea is to make sure that the initial state of the ensemble, i.e., prior to training, represents a state of full uncertainty (or no knowledge). Hence, the resulting credal prediction should entail the entire probability simplex. As the predictors in the ensemble are trained, the amount of knowledge increases, and the predicted credal set should shrink. To enforce this in a finite predictor scenario, we ensure that the initial predictions of the learners correspond to degenerate probability distributions at vertices of the $(K-1)$-simplex, effectively covering the entire simplex in the initially predicted credal set. Figure 1 illustrates the mechanism of ToBias initialization and the overall learning process.
Technically, the initialization is performed by assigning a large constant to one of the biases in the final layer of each predictor. For example, given the biases $\bm{b}_{i}=[b_{i,1},...,b_{i,K}]$ of the last layer of predictor $h_{i}$, we set $b_{i,\ i\ \text{mod}\ K}=\beta$, where $\beta$ is large constant, following the regular initialization. In the remainder of this work, we use $\beta=100$. An ablation study examining the influence of ToBias initialization is presented in Section C.2.
Figure 3: Effect of ToBias. Example from ChaosNLI. a) Standard initialization initially predicts close to uniform distributions and converges to a small credal set. b) ToBias initializes at the vertices resulting in diverse predictions.
We illustrate the difference between regular initialization and ToBias initialization in Figure 3. Observe that without ToBias initialization, the credal set initially represents a state of low epistemic uncertainty, as all predictions are concentrated around the barycenter of the probability simplex, and increases its uncertainty as training progresses. The opposite can be observed with ToBias initialization. The credal set starts representing a state of full uncertainty and decreases its uncertainty as the learners acquire more knowledge.
Having trained $M$ models until the respective thresholds, the approximate $\alpha$-cut $\tilde{\mathcal{C}}_{\alpha}$ can be constructed as follows:
$\tilde{\mathcal{C}}_{\alpha}=\{h_{i}\ :\hat{\gamma}(h_{i})\approx\tau_{i}\}_{i=0}^{M-1}\subseteq\mathcal{H},$
where $\hat{\gamma}(h_{i})\approx\tau_{i}$ is enforced by the training process described above. The full algorithm to train Credal Relative Likelihood (CreRLα) ensembles is presented in Algorithm 1.
Credal Set Predictions. Given a query instance $\bm{x}\in\mathcal{X}$, the set $\tilde{\mathcal{C}}_{\alpha}$ induces a finite approximation of the credal set:
$\tilde{\mathcal{Q}}_{\bm{x},\alpha}=\{p(\cdot\,|\,\bm{x},h):h\in\tilde{\mathcal{C}}_{\alpha}\}\subseteq\Delta_{K}.$
There are several principled ways of transforming this finite set into a credal set. Two common approaches are to derive (i) the convex hull of the set $\tilde{\mathcal{Q}}_{\bm{x},\alpha}$ and (ii) the intervals (lower and upper probabilities) for individual classes obtained from this set. Note that the convex hull is contained in the “box” credal set induced by the intervals. In the remainder of this work, we opt for the latter approach and define the approximation of $\mathcal{Q}_{\bm{x},\alpha}$ as follows (de Campos et al., 1994):
$\hat{\mathcal{Q}}_{\bm{x},\alpha}=\{p(\cdot\,|\,\bm{x}):p(y_{k}\,|\,\bm{x})\in[\underline{p}(y_{k}\,|\,\bm{x}),\overline{p}(y_{k}\,|\,\bm{x})],\forall y_{k}\in\mathcal{Y}\}$
5 Empirical Results
To empirically evaluate our method, we measure its performance in terms of coverage and efficiency, as defined in Section 2, along with the corresponding Pareto criterion. Additionally, we evaluate our approach on the downstream task of OoD detection, which is often used to assess the quality of (epistemic) uncertainty representations. We compare our approach (CreRLα) to suitable baselines by implementing the following approaches: Credal Wrapper (CreWra) (Wang et al., 2024b), Credal Ensembling (CreEnsα) (Nguyen et al., 2025), Credal Bayesian Deep Learning (CreBNN) (Caprio et al., 2023), and Credal Deep Ensembles (CreNet) (Wang et al., 2024a). These methods represent the current state-of-art for credal prediction and to the best of our knowledge we are the first to do a systematic comparison of these approaches in a unified benchmark. Since the coverage and efficiency evaluation metrics are not well-defined for Bayesian methods, a direct comparison to such methods is not possible.
The code for all implementations and experiments is published in a Github repository [^1] https://github.com/timoverse/credal-prediction-relative-likelihood. Further experimental details of our method and the implementation of the baselines are provided in Appendix B.
Figure 4: Pareto front between coverage and efficiency for ChaosNLI, CIFAR-10, and QualityMRI. CreRLα (Ours) and CreEnsα allow trading off efficiency and coverage by varying $\alpha$. The other baselines do not allow direct adaptation of the credal set and hence result in a single point.
5.1 Predictive Performance
We use our method to train an ensemble of fully-connected neural networks on the embeddings of the ChaosNLI dataset (Nie et al., 2020). This dataset features premise-hypothesis pairs in textual form with the goal to classify each instance as one of three classes: entailed, neutral, and contradicted. Each instance in ChaosNLI has 100 annotations, which can be seen as a ground-truth distribution. We train an ensemble of ResNet18 models (He et al., 2016) based on our method on the CIFAR-10 dataset (Krizhevsky et al., 2009). This dataset consists of 10 classes of images, commonly used as a benchmark for image classification tasks. We use the human annotations provided by the CIFAR-10H dataset (Peterson et al., 2019) as ground-truth distributions. In addition, we train an ensemble of ResNet18 models on the QualityMRI dataset from Schmarje et al. (2022). This dataset comprises MRI images with multiple annotations from radiologists, each providing a quality assessment of the image (Obuchowicz et al., 2020). These multi-class annotations were merged into annotations for two classes (Schmarje et al., 2022). The uncertainty in these labels, as represented by the high average entropy reported in Appendix C, confirms the importance of uncertainty quantification in medical settings (Löhr et al., 2024).
To train the neural networks, the annotations are converted into a probability distribution, which we consider to be the ground-truth. The neural networks are then trained on targets that are sampled from the ground-truth probability distribution. Thus, our method does not require ground-truth probability distributions, or soft labels, for training.
We evaluate our method using the coverage and efficiency as defined in (2) and (3), where the expectation is approximated by averaging over the the instances in the test sets of the respective datasets. The results can be found in Figure 4. In Appendix C, we report the accuracy of the ensemble members for each implementation.
Trade-off between Efficiency and Coverage. In general, both our method (CreRLα) and the CreEnsα approach are flexible and allow to trade off efficiency and coverage of the credal predictor by varying $\alpha$. However, CreEnsα can only attain credal sets in the low coverage region, whereas our approach can predict credal sets with a wide range of coverage values, including high coverage. The other methods only deliver fixed credal sets, meaning there is no direct way to to increase (or decrease) the coverage and efficiency. The CreWra approach and $\text{CreEns}_{0}$ (with $\alpha=0$) will generally have the same efficiency, but a different coverage, due to the different credal set constructions (interval vs. convex hull). The CreNet converges to a point prediction, because its loss function is constructed such that the lower probability converges to the upper probability, and hence is generally in the top left region of the Pareto figure. We refer to Appendix A for more details about these methods.
For the ChaosNLI dataset, our approach dominates the CreBNN approach and the CreWra resembles a point on the Pareto front obtained by our method. As mentioned, the CreEnsα method can only reach the low coverage region and the CreNet is in the top left.
On CIFAR-10, our method generates a Pareto front in a region of high efficiency and high coverage. The CreWra baseline is close to our Pareto front. Since all ensemble members have high performance, as illustrated by their accuracies in Appendix C, and there is no explicit diversity promotion in CreEnsα, it produces small sets for every $\alpha$. CreNet, again, produces a (close to) point prediction with low coverage. We refer to Appendix B for more details.
For QualityMRI, the story is largely consistent with that of the ChaosNLI dataset. The difference is that our method is only positioned in the high coverage half of the Pareto figure.
The ability of our method to target the high coverage area can be explained by the fact that we promote diversity, by means of our ToBias initialization, in order to promote high coverage, at the cost of efficiency. In addition, one may argue that the low coverage region represents credal sets that are of limited benefit for practitioners, as they are unlikely to demand a coverage of, say, 0.3.
Furthermore, Figure 4 shows that the shape and position of the Pareto front varies across datasets suggesting that the coverage-efficiency trade-off is inherently data-dependent. In particular, the better the maximum likelihood model, the more the Pareto front should move towards the top right (high coverage, high efficiency), which we also observe in the case of CIFAR-10.
To further study the behavior of our method, we perform ablation experiments, varying the number of ensemble members and the ToBias initialization constant. The results are presented in Appendix C.
Table 1: Out-of-Distribution detection based on epistemic uncertainty using CIFAR-10 as in-Distribution dataset. Mean and standard deviation over 3 runs are reported. Best performance is bold.
Method SVHN Places CIFAR-100 FMNIST ImageNet
CreWra $\mathbf{0.957\scriptstyle{\pm 0.003}}$ $0.916\scriptstyle{\pm 0.001}$ $\mathbf{0.916\scriptstyle{\pm 0.000}}$ $0.952\scriptstyle{\pm 0.000}$ $\mathbf{0.890\scriptstyle{\pm 0.001}}$
$\text{CreEns}_{0.0}$ $0.955\scriptstyle{\pm 0.001}$ $0.913\scriptstyle{\pm 0.000}$ $0.914\scriptstyle{\pm 0.001}$ $0.949\scriptstyle{\pm 0.001}$ $0.888\scriptstyle{\pm 0.000}$
CreBNN $0.907\scriptstyle{\pm 0.006}$ $0.885\scriptstyle{\pm 0.002}$ $0.880\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.002}$ $0.859\scriptstyle{\pm 0.002}$
CreNet $0.943\scriptstyle{\pm 0.003}$ $\mathbf{0.918\scriptstyle{\pm 0.000}}$ $0.912\scriptstyle{\pm 0.000}$ $0.951\scriptstyle{\pm 0.002}$ $0.884\scriptstyle{\pm 0.001}$
$\text{CreRL}_{1.0}$ $0.948\scriptstyle{\pm 0.003}$ $\mathbf{0.918\scriptstyle{\pm 0.002}}$ $\mathbf{0.916\scriptstyle{\pm 0.001}}$ $\mathbf{0.957\scriptstyle{\pm 0.002}}$ $\mathbf{0.889\scriptstyle{\pm 0.002}}$
$\text{CreRL}_{0.95}$ $0.917\scriptstyle{\pm 0.013}$ $0.910\scriptstyle{\pm 0.001}$ $0.901\scriptstyle{\pm 0.000}$ $0.945\scriptstyle{\pm 0.004}$ $0.878\scriptstyle{\pm 0.002}$
$\text{CreRL}_{0.9}$ $0.918\scriptstyle{\pm 0.011}$ $0.907\scriptstyle{\pm 0.001}$ $0.896\scriptstyle{\pm 0.001}$ $0.944\scriptstyle{\pm 0.004}$ $0.874\scriptstyle{\pm 0.001}$
$\text{CreRL}_{0.8}$ $0.906\scriptstyle{\pm 0.008}$ $0.894\scriptstyle{\pm 0.001}$ $0.884\scriptstyle{\pm 0.003}$ $0.936\scriptstyle{\pm 0.009}$ $0.865\scriptstyle{\pm 0.002}$
$\text{CreRL}_{0.6}$ $0.862\scriptstyle{\pm 0.035}$ $0.874\scriptstyle{\pm 0.003}$ $0.852\scriptstyle{\pm 0.002}$ $0.893\scriptstyle{\pm 0.005}$ $0.837\scriptstyle{\pm 0.003}$
$\text{CreRL}_{0.4}$ $0.739\scriptstyle{\pm 0.029}$ $0.821\scriptstyle{\pm 0.007}$ $0.796\scriptstyle{\pm 0.007}$ $0.815\scriptstyle{\pm 0.020}$ $0.788\scriptstyle{\pm 0.010}$
$\text{CreRL}_{0.2}$ $0.582\scriptstyle{\pm 0.041}$ $0.736\scriptstyle{\pm 0.010}$ $0.700\scriptstyle{\pm 0.013}$ $0.676\scriptstyle{\pm 0.046}$ $0.698\scriptstyle{\pm 0.013}$
5.2 Out-of-Distribution Detection
Due to a lack of ground-truth uncertainties for most datasets, uncertainty methods are usually evaluated on downstream tasks such as OoD detection. We perform the OoD task to assess the (epistemic) uncertainty representation of our method. The predictor is trained on a dataset called the in-Distribution (iD) dataset and at test time introduced to instances from both the iD dataset and another OoD dataset that it has not seen before. An effective epistemic uncertainty representation should assign higher epistemic uncertainty to OoD instances than to iD instances. In the literature, this is commonly evaluated by computing epistemic uncertainty and using it to distinguish between iD and OoD instances. We adopt the epistemic uncertainty measure based on the additive decomposition proposed (and axiomatically justified) by Abellán et al. (2006):
$\operatorname{EU}(\hat{\mathcal{Q}}_{\bm{x},\alpha})=\overline{S}(\hat{\mathcal{Q}}_{\bm{x},\alpha})-\underline{S}(\hat{\mathcal{Q}}_{\bm{x},\alpha}),$ (5)
where $\overline{S}(\hat{\mathcal{Q}}_{\bm{x},\alpha})$ denotes the maximum Shannon entropy of all distributions in $\hat{\mathcal{Q}}_{\bm{x},\alpha}$ and, likewise, $\underline{S}(\hat{\mathcal{Q}}_{\bm{x},\alpha})$ the minimum entropy. [^2] The bounds of Shannon entropy are computed numerically using SciPy (Virtanen et al., 2020). The performance is then measured using the AUROC. Section B.4 provides details about the implementation of the optimization.
We utilize the same CIFAR-10 ensembles of ResNet18 models and compare to the same baselines as in Section 5.1. We introduce SVHN Netzer et al. (2011), Places365 Zhou et al. (2018), CIFAR-100 Krizhevsky et al. (2009), FashionMNIST Xiao et al. (2017), and ImageNet Deng et al. (2009) as out-of-distribution datasets, and compute epistemic uncertainty as in (5). Our method ($\text{CreRL}_{\alpha}$) and the Credal Ensembling ($\text{CreEns}_{\alpha}$) approach make use of a hyper-parameter $\alpha$. The results with the best performing $\alpha$ are presented here. The results, summarized in Table 1, present the OoD detection performance. Additionally, we provide results for both methods with different $\alpha$ values in Appendix C.
Our method (CreRLα) with $\alpha=1$ performs either the best or is on par with the best methods in OoD detection, except for the SVHN dataset, where the CreWra approach has a small advantage. This shows the good (epistemic) uncertainty representation of our method.
Influence of $\bm{\alpha}$. Furthermore, the value of the $\alpha$ parameter of our method has a significant impact on the behavior of credal sets for both iD and OoD data. We expect greater epistemic uncertainty for instances from a distribution that differs from the training data, as predictions may diverge for these instances. When $\alpha$ is close to 1, the resulting set consists of models with high relative likelihood, leading to smaller credal sets and thus lower epistemic uncertainty for iD data. Conversely, a smaller $\alpha$ includes low relative likelihood models, resulting in higher epistemic uncertainty. Given that $\alpha$ strongly influences epistemic uncertainty for iD instances, we expect a larger $\alpha$ to increase the discrepancy in uncertainty between iD and OoD instances. As shown in Table 1, across all OoD datasets, a larger $\alpha$ indeed leads to higher AUROC scores, confirming that larger $\alpha$ values lead to better separation of iD and OoD instances.
6 Conclusion
We proposed a new approach to generating credal predictors in a machine learning setting with complex learners on the basis of relative likelihood, which is used to identify a set of “plausible” models. The relative likelihood is an intuitive notion of model plausibility, grounded in classical statistics. Given a query instance for which a prediction is sought, a credal set is obtained as the collection of probability distributions predicted by all plausible models. Specifically, we introduced a novel method for training neural networks to approximate the credal set generated by an $\alpha$-cut of models. In order to obtain a good approximation, the $\alpha$-cut needs to be covered well, i.e., the included hypotheses should be sufficiently spread over the hypothesis space. To this end, we proposed ToBias initialization as a simple but effective diversification strategy. Furthermore, the parameter $\alpha$, which specifies the likelihood ratio between the maximum likelihood model and the “weakest” model that is still included in the $\alpha$-cut, can be adjusted depending on the task at hand and used to control the trade-off between the coverage and efficiency of credal predictions.
Experimentally, we have shown that this workflow performs well in terms of providing coverage of the ground-truth conditional distributions, while maintaining a competitive efficiency, as well as performing competitively in downstream tasks such as OoD detection. The relation between $\alpha$ and the performance on different tasks sheds further light on the trade-off between diversity and prediction performance.
Limitations and Future Work. Our method has so far been evaluated exclusively on neural networks. However, the general concept of relative likelihood is applicable to any model class. Certain aspects of our method — such as ToBias initialization — are specifically tailored to neural networks and would have to be adapted to other learners. A promising direction for future work would be to extend this framework to other model classes.
When generating a convex credal set from our finite approximation, we include all probability distributions such that the respective class probabilities are bounded by their lower and upper probabilities. This makes the optimization problem of quantifying uncertainty easier, but it also generates sets that are larger than necessary. An alternative would be to take the convex hull of the probability distributions as the credal set, which would generate more efficient sets. Future work may explore the effect of this on metrics such as coverage or OoD performance.
Another interesting question that arises relates to the approximation quality of our approach. Having specified a threshold for the relative likelihood, there exists some ground-truth credal set for an instance. It could be interesting to explore how many probability distributions, and thus predictors, are required to accurately approximate this credal set.
Broader Impacts. This work contributes to the development of more reliable machine learning models by improving uncertainty quantification. We do not foresee any direct negative broader impacts arising from this.
Appendix A Extended Related Work
Uncertainty Representation. In machine learning, uncertainty is often represented by Bayesian methods [Mackay, 1992, Blundell et al., 2015]. Frequently, the Bayesian posterior is approximated by ensembles [Lakshminarayanan et al., 2017] or methods such as Dropout [Gal and Ghahramani, 2016] or Laplace approximation [Daxberger et al., 2021]. An important characteristic of such representations, especially for uncertainty tasks, is diversity [D’Angelo and Fortuin, 2021, Wood et al., 2023]. Some works enforce this by means of regularization [de Mathelin et al., 2023], whereas others vary hyperparameters across models within the ensemble [Wenzel et al., 2020] or enforce diversity in the representations [Lopes et al., 2022].
Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty [Zaffalon, 2001, Corani and Zaffalon, 2008, Corani and Mignatti, 2015]. Antonucci et al. [2012a] proposed to generate such sets based on relative likelihoods. Relative likelihoods, also referred to as normalized likelihoods, have also been used in machine learning with simple model classes such as logistic regression [Senge et al., 2014, Cella and Martin, 2024]. In this work, we build upon these approaches and address the challenges that emerge when adapting the relative likelihood to a setting with complex predictors.
Recently, credal sets have been applied in the context of machine learning. Wang et al. [2024b] take multiple samples from a Bayesian posterior or ensembles and derive class-wise lower and upper probabilities from this. Based on these samples or ensemble predictions, a credal set is constructed by including all probability distributions such that the individual class probabilities are in the respective lower to upper class probability interval. Nguyen et al. [2025] also use ensemble predictions, but they additionally allow for the discarding of potential outliers to ensure that the resulting credal set too large. This is done by comparing all ensemble predictions to a representative prediction, e.g. the mean prediction, using some distance between distributions, and keeping only $(1-\alpha)\cdot 100\%$ closest predictions. The credal set is then constructed by taking the convex hull of the remaining probability distributions. Other methods directly train neural networks to predict intervals by explicitly having the neural network predict a lower and upper probability for every class [Wang et al., 2024a]. In combination with a custom loss function, consisting of regular cross-entropy for the upper probabilities and cross-entropy computed on the highest loss subset of a batch for the lower probabilities, the claim is that this approach encourages both “optimistic” and “pessimistic” predictions. Credal sets are constructed by taking the same interval-based approach as [Wang et al., 2024b]. Besides this, hybrid methods, combining multiple uncertainty frameworks, have also been proposed. Caprio et al. [2023] combine Bayesian deep learning and credal sets by considering sets of priors over weights of neural networks. Training these neural network by variational inference then results in a set of posteriors. Based on this, a credal set is constructed by sampling from the Bayesian neural networks and taking the convex hull of the sampled probability distributions. Another approach leverages conformal prediction to construct credal sets with validity guarantees [Javanmardi et al., 2024]. However, this work uses distributions over classes to perform the conformal calibration step. Since our method does not require such data, we do not use this method as a baseline. Our work distinguishes itself from aforementioned works by making use of relative likelihood cuts which allow for an intuitive and adaptable construction of the credal set.
Uncertainty Quantification. Given a credal representation of uncertainty, there are many ways to quantify uncertainty. Some measures consider only the epistemic uncertainty [Abellán and Moral, 2000], whereas others use entropy to reason about the total uncertainty [Abellán and Moral, 2003]. In this line of work, Abellán et al. [2006] also proposed measures based on entropy that decompose the total uncertainty into an aleatoric and epistemic component. Conversely, Antonucci et al. [2012b] quantify uncertainty by measuring the lack of dominance of one class over others in the predictive distributions represented by the credal set. Hüllermeier et al. [2022] offer a critical analysis of these measures and propose an alternative for the dominance-based measure. Recently, Hofman et al. [2024] proposed to quantify credal uncertainty based on a decomposition of scoring rules. In the following, we consider the uncertainty measures by Abellán et al. [2006] in order to ensure a fair comparison with previous works.
Appendix B Experimental Details
B.1 Datasets
ChaosNLI The ChaosNLI dataset, introduced by Nie et al. [2020], is a large-scale dataset designed to study human disagreement in natural language inference (NLI) tasks. It comprises 100 human annotations per example for 3,113 examples from the SNLI and MNLI datasets, and 1,532 examples from the $\alpha$NLI dataset, totaling approximately 464,500 annotations. In line with Javanmardi et al. [2024], we use only the SNLI and MNLI subsets of the ChaosNLI dataset, but for simplicity, we will refer to this as the ChaosNLI dataset. Each example includes metadata such as the unique identifier, counts of each label assigned by annotators, the majority label, label distribution, entropy of the label distribution, the original example text, and the original label from the source dataset. This dataset enables a detailed analysis of the distribution of human opinions in NLI tasks, highlighting instances of high disagreement and questioning the validity of using majority labels as the sole ground truth. ChaosNLI is publicly available under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license. We train our models on the 768-dimensional embeddings of the ChaosNLI dataset retrieved from https://github.com/alireza-javanmardi/conformal-credal-sets. We refer to [Javanmardi et al., 2024] for more details on the generation of the embeddings.
CIFAR-10 The CIFAR-10 dataset is a widely used benchmark in machine learning and computer vision, introduced by Krizhevsky et al. [2009], and Geoffrey Hinton in 2009. It comprises 60,000 color images at a resolution of 32×32 pixels, evenly distributed across 10 distinct classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. The dataset is partitioned into 50,000 training images and 10,000 test images, organized into five training batches and one test batch, each containing 10,000 images. The dataset is publicly available and has been utilized extensively for developing and benchmarking machine learning models. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
CIFAR-10H The CIFAR-10H dataset provides human-derived soft labels for the 10,000 images in the CIFAR-10 test set, capturing the variability in human annotation during image classification tasks. Developed by Peterson et al. [2019], the dataset comprises 511,400 annotations collected from 2,571 Amazon Mechanical Turk workers, with each image receiving approximately 51 labels. Annotators classified images into one of the ten CIFAR-10 categories, enabling the construction of probability distributions over labels for each image. CIFAR-10H is publicly available under the Creative Commons BY-NC-SA 4.0 license.
CIFAR-100 The CIFAR-100 dataset, introduced by Krizhevsky et al. [2009], comprises 60,000 color images at 32×32 resolution, divided into 100 classes with 600 images each. Each image has a “fine” label (specific class) and a “coarse” label (superclass), with the 100 classes grouped into 20 superclasses. The dataset is split into 50,000 training and 10,000 test images. It is a subset of the Tiny Images dataset and is commonly used for evaluating image classification algorithms. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
QualityMRI The QualityMRI dataset, introduced by Obuchowicz et al. [2020], is part of the Data-Centric Image Classification (DCIC) Benchmark, which aims to evaluate the impact of dataset curation on model performance. The dataset contains 310 magnetic resonance (MRI) images spanning various quality levels and is designed to assess the MRI image quality. The dataset is publicly available under the Creative Commons BY-SA 4.0 license.
SVHN The SVHN dataset, introduced by Netzer et al. [2011], consists of over 600,000 32×32 RGB images of digits (0–9) obtained from real-world house number images in Google Street View. It includes three subsets: 73,257 training images, 26,032 test images, and 531,131 additional images for extra training. The dataset is designed for digit recognition tasks with minimal preprocessing. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
Places365 Places365, introduced by Zhou et al. [2018], is a large-scale scene recognition dataset containing 1.8 million training images across 365 scene categories. The validation set includes 50 images per category, and the test set has 900 images per category. An extended version, Places365-Challenge-2016, adds 6.2 million images and 69 new scene classes, totaling 8 million images over 434 categories. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
FMNIST Fashion-MNIST (FMNIST), introduced by Xiao et al. [2017], is a dataset of Zalando’s article images, comprising 70,000 28×28 grayscale images labeled across 10 classes, such as T-shirt/top, Trouser, and Sneaker. It includes 60,000 training and 10,000 test images and serves as a direct replacement for the original MNIST dataset for benchmarking machine learning algorithms. FMNIST is publicly available under the MIT License.
ImageNet ImageNet, introduced by Deng et al. [2009], is a large-scale image database organized according to the WordNet hierarchy, containing over 14 million images across more than 20,000 categories. The ILSVRC subset (ImageNet-1K) includes 1,281,167 training images, 50,000 validation images, and 100,000 test images across 1,000 classes. The dataset is available for free to researchers for non-commercial use.
B.2 Models
Fully-Connected Network We train fully-connected neural networks on the ChaosNLI dataset. The network consists of 4 linear layers with $[768-256-64-16-3]$ units with ReLU activations, except for the last layer, which has Softmax transformation to transform the logits into probabilities. We use the hyperparameters (cf. Table 2) similar to the optimal parameters found in Javanmardi et al. [2024].
ResNet18 For experiments on the CIFAR-10 dataset, we use the PyTorch ResNet-18 implementation and hyperparameters provided by https://github.com/kuangliu/pytorch-cifar. This model is specifically optimized for CIFAR-10 and is trained from scratch, without any pretraining on ImageNet. For experiments on the QualityMRI dataset, we use the ResNet18 implementation from the PyTorch torchvision package with random initialization, i.e. no pretrained weights.
Hyperparameters Each dataset is trained using a dedicated set of hyperparameters as presented in Table 2. We evaluated multiple configurations and selected the best-performing ones for each dataset. To ensure fair and consistent comparisons, all models trained on a given dataset, both our approach and the baselines, use the same hyperparameter settings. The only exception is the CreBNN, which requires a KL-divergence penalty of $71e-$ 7 and zero weight decay when using the Adam optimizer [Kingma and Ba, 2015]. When we apply the SGD optimizer with a learning rate scheduler, namely Cosine Annealing [Loshchilov and Hutter, 2017], CreBNN requires additionally a momentum of $0.9$ to enable effective learning.
Table 2: Hyperparameters used for each dataset.
Hyperparameter ChaosNLI CIFAR-10 QualityMRI
Model FCNet ResNet18 ResNet18
Epochs 300 200 200
Learning rate 0.01 0.1 0.01
Weight decay 0.0 0.0005 0.0005
Optimizer Adam SGD SGD
Ensemble members 20 20 20
LR scheduler - CosineAnnealing CosineAnnealing
Tobias value 100 100 100
B.3 Out-of-Distribution Detection
We use the SVHN [Netzer et al., 2011], Places365 [Zhou et al., 2018], CIFAR-100 [Krizhevsky et al., 2009], Fashion-MNIST [Xiao et al., 2017], and ImageNet [Le and Yang, 2015] datasets for Out-of-Distribution detection.
The Out-of-Distribution task is treated as a binary classification task where the epistemic uncertainty is used as the classification criterion. In order to balance the data, we sample 10000 instances from the test set of the respective datasets. On both the in-Distribution data (CIFAR-10) and the Out-of-Distribution data, the same transforms — normalization and resizing to 32 by 32 pixels — are applied. After computing the epistemic uncertainty the area under the receiver operating characteristics curve (AUROC) is computed and used as the comparison metric.
B.4 Computing Uncertainty
Computing the total and aleatoric uncertainty involves solving a per instance optimization problem, after which the epistemic uncertainty is obtained as their difference (see Equation 5). This optimization is performed using SciPy’s minimize function with the SLSQP solver and the default parameters.
Interval-Based For the interval-based credal sets, the optimization is initialized with the mean of the predicted distributions, bounded between the lower and upper probabilities for each class, and constrained to ensure that the solution forms a valid probability distribution, namely, the class probabilities must sum to $1$.
Convex Hull For the credal sets based on the convex hull, the optimization is done on the weights of convex combination, instead of the probability distribution. Uniform weights are used as the initial value, the weights are bounded between 00 and $1$, and constrained to sum to $1$.
Estimated Computing Time Here, we provide an estimated upper bound of the computation time for computing the lower and upper entropy of a credal set based on the example of the largest possible credal set, namely the full probability simplex for a $10$-class problem, such as CIFAR-10. For this largest set, computing the upper entropy using an interval-based credal set takes on average $0.03$ seconds, while computing the lower entropy takes about $0.02$ seconds. For the convex hull-based credal set of the same size, the average computation time is $0.07$ seconds for upper entropy and $0.03$ seconds for lower entropy. To simplify our estimation, we average these times and assume that for a single instance it takes $0.03$ seconds to optimize the lower entropy and $0.04$ seconds to optimize upper entropy, summing up to $0.07$ seconds of computing time per instance.
In the Out-of-Distribution (OoD) detection experiments, we need to compute both lower and upper entropy for each instance in both the in-Distribution (iD) and OoD datasets. Each dataset contains $10,000$ instances, resulting in roughly $23$ minutes of computation time per model to obtain the epistemic uncertainty across all instances. Since we run three seeds per model and evaluate $12$ different models, the total runtime for a single OoD experiment on one dataset is approximately $14$ hours. As we evaluate OoD detection across five datasets, the total computing time amounts to around $70$ hours excluding any time needed to train models beforehand.
B.5 Computing Coverage
Coverage is evaluated by checking whether the ground-truth distribution lies within the predicted credal set. For the interval-based approach, this involves verifying that each class probability of the ground-truth distribution falls between the corresponding lower and upper bounds of the credal set. For the convex hull approach, we assess whether the ground-truth distribution can be expressed as a convex combination of the extreme points defining the credal set. This optimization is done using the SciPy linprog function.
B.6 Baselines
We list all details regarding the implementations of the baselines that were used in the paper. In general, all baselines were implemented in our own code base.
Credal Wrapper (CreWra) The Credal Wrapper was initially implemented in TensorFlow, but we reimplemented it in PyTorch to ensure compatibility with our framework. It follows a standard ensemble learning approach, training multiple models independently. Like our method, the Credal Wrapper constructs credal sets using class-wise upper and lower probability bounds, making it well-aligned with our implementation. Overall, we closely follow both the original paper and their available implementation [Wang et al., 2024b].
Credal Ensembling ($\text{CreEns}_{\alpha}$) Since no official code was available for neural network implementations of Credal Ensembling, we reimplemented the method ourselves. Our implementation closely follows all details provided in Nguyen et al. [2025]. The approach builds on standard ensemble training procedures, with inference adapted according to their proposed method of sorting predictions based on a distance measure and selecting only $\alpha\%$ closest predictions to construct credal sets. In our experiments, we use the Euclidean distance measure and evaluate several values of $\alpha$.
Credal Deep Ensembles (CreNet) As the official implementation of Credal Deep Ensembles is only available in TensorFlow, we reimplemented the method in PyTorch to ensure compatibility with our codebase. Our implementation closely mirrors the original TensorFlow code, particularly in adapting the model architecture and loss function. Specifically, we replace each model’s final linear layer with a head comprising a linear layer outputting $2\times\text{classes}$ (representing upper and lower probability bounds), followed by a batch normalization layer and the custom IntSoftmax layer. We also reimplemented the proposed loss function, which computes a cross-entropy loss for the upper bounds and selectively backpropagates the lower-bound loss only for the $\delta\%$ of samples with the highest loss values, as described in Wang et al. [2024a]. In our experiments, we use $\delta=0.5$, as suggested in Wang et al. [2024a].
Credal Bayesian Deep Learning (CreBNN) No code or implementation details for Credal Bayesian Deep Learning (CreBNN) were made publicly available, and despite multiple attempts to contact the authors, we received no response or further clarification. As a result, we reimplemented the method ourselves based solely on the high-level description provided in the paper. In our implementation, each ensemble member is a Bayesian neural network (BNN) trained with variational inference using different priors, with prior means $\mu$ sampled from $[-1,1]$ and standard deviations $\sigma$ from $[0.1,2]$ to form a diverse prior set. During inference, we draw one sample from each BNN to obtain a finite set of probability distributions, and construct the credal set as the convex hull of these predictions.
B.7 Computing Resources
To run the experiments presented in this work, we utilized the computing resources detailed in Table 3. The total estimated GPU usage amounts to approximately $750$ hours.
Table 3: Specifications of Computing Resources
Component Specification
CPU AMD EPYC MILAN 7413 Processor, 24C/48T 2.65GHz 128MB L3 Cache
GPU 2 × NVIDIA A40 (48 GB GDDR each)
RAM 128 GB (4x 32GB) DDR4-3200MHz ECC DIMM
Storage 2 × 480GB Samsung Datacenter SSD PM893
Appendix C Additional Experiments
This section presents additional results, including ablation studies and alternative hyperparameter configurations, that complement the findings reported in the main paper.
C.1 Out-of-Distribution detection
In addition to the OoD experiments in Section 5, we evaluate a broad range of values $\alpha$ for the CreEnsα approach.
Table 4: Out-of-Distribution detection based on epistemic uncertainty. CIFAR-10 is used as the in-Distribution dataset. The mean and standard deviation over 3 runs are reported. Best performance is in bold.
Method SVHN Places CIFAR-100 FMNIST ImageNet
CreWra $\mathbf{0.957\scriptstyle{\pm 0.003}}$ $0.916\scriptstyle{\pm 0.001}$ $\mathbf{0.916\scriptstyle{\pm 0.000}}$ $0.952\scriptstyle{\pm 0.000}$ $\mathbf{0.890\scriptstyle{\pm 0.001}}$
$\text{CreEns}_{0.95}$ $0.500\scriptstyle{\pm 0.000}$ $0.500\scriptstyle{\pm 0.000}$ $0.500\scriptstyle{\pm 0.000}$ $0.500\scriptstyle{\pm 0.000}$ $0.500\scriptstyle{\pm 0.000}$
$\text{CreEns}_{0.9}$ $0.921\scriptstyle{\pm 0.002}$ $0.879\scriptstyle{\pm 0.002}$ $0.883\scriptstyle{\pm 0.001}$ $0.915\scriptstyle{\pm 0.001}$ $0.857\scriptstyle{\pm 0.002}$
$\text{CreEns}_{0.8}$ $0.937\scriptstyle{\pm 0.001}$ $0.896\scriptstyle{\pm 0.001}$ $0.900\scriptstyle{\pm 0.001}$ $0.929\scriptstyle{\pm 0.001}$ $0.875\scriptstyle{\pm 0.001}$
$\text{CreEns}_{0.6}$ $0.944\scriptstyle{\pm 0.002}$ $0.902\scriptstyle{\pm 0.001}$ $0.906\scriptstyle{\pm 0.000}$ $0.935\scriptstyle{\pm 0.001}$ $0.881\scriptstyle{\pm 0.001}$
$\text{CreEns}_{0.4}$ $0.947\scriptstyle{\pm 0.001}$ $0.906\scriptstyle{\pm 0.001}$ $0.908\scriptstyle{\pm 0.000}$ $0.940\scriptstyle{\pm 0.001}$ $0.883\scriptstyle{\pm 0.001}$
$\text{CreEns}_{0.2}$ $0.950\scriptstyle{\pm 0.001}$ $0.909\scriptstyle{\pm 0.001}$ $0.911\scriptstyle{\pm 0.000}$ $0.946\scriptstyle{\pm 0.001}$ $0.885\scriptstyle{\pm 0.001}$
$\text{CreEns}_{0.0}$ $0.955\scriptstyle{\pm 0.001}$ $0.913\scriptstyle{\pm 0.000}$ $0.914\scriptstyle{\pm 0.001}$ $0.949\scriptstyle{\pm 0.001}$ $0.888\scriptstyle{\pm 0.000}$
CreNet $0.943\scriptstyle{\pm 0.003}$ $0.918\scriptstyle{\pm 0.000}$ $0.912\scriptstyle{\pm 0.000}$ $0.951\scriptstyle{\pm 0.002}$ $0.884\scriptstyle{\pm 0.001}$
CreBNN $0.907\scriptstyle{\pm 0.006}$ $0.885\scriptstyle{\pm 0.002}$ $0.880\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.002}$ $0.859\scriptstyle{\pm 0.002}$
$\text{CreRL}_{1.0}$ $0.948\scriptstyle{\pm 0.003}$ $\mathbf{0.918\scriptstyle{\pm 0.002}}$ $\mathbf{0.916\scriptstyle{\pm 0.001}}$ $\mathbf{0.957\scriptstyle{\pm 0.002}}$ $\mathbf{0.889\scriptstyle{\pm 0.002}}$
$\text{CreRL}_{0.95}$ $0.917\scriptstyle{\pm 0.013}$ $0.910\scriptstyle{\pm 0.001}$ $0.901\scriptstyle{\pm 0.000}$ $0.945\scriptstyle{\pm 0.004}$ $0.878\scriptstyle{\pm 0.002}$
$\text{CreRL}_{0.9}$ $0.918\scriptstyle{\pm 0.011}$ $0.907\scriptstyle{\pm 0.001}$ $0.896\scriptstyle{\pm 0.001}$ $0.944\scriptstyle{\pm 0.004}$ $0.874\scriptstyle{\pm 0.001}$
$\text{CreRL}_{0.8}$ $0.906\scriptstyle{\pm 0.008}$ $0.894\scriptstyle{\pm 0.001}$ $0.884\scriptstyle{\pm 0.003}$ $0.936\scriptstyle{\pm 0.009}$ $0.865\scriptstyle{\pm 0.002}$
$\text{CreRL}_{0.6}$ $0.862\scriptstyle{\pm 0.035}$ $0.874\scriptstyle{\pm 0.003}$ $0.852\scriptstyle{\pm 0.002}$ $0.893\scriptstyle{\pm 0.005}$ $0.837\scriptstyle{\pm 0.003}$
$\text{CreRL}_{0.4}$ $0.739\scriptstyle{\pm 0.029}$ $0.821\scriptstyle{\pm 0.007}$ $0.796\scriptstyle{\pm 0.007}$ $0.815\scriptstyle{\pm 0.020}$ $0.788\scriptstyle{\pm 0.010}$
$\text{CreRL}_{0.2}$ $0.582\scriptstyle{\pm 0.041}$ $0.736\scriptstyle{\pm 0.010}$ $0.700\scriptstyle{\pm 0.013}$ $0.676\scriptstyle{\pm 0.046}$ $0.698\scriptstyle{\pm 0.013}$
C.2 Ablations
In the following ablation study, we evaluate two factors: the impact of varying the initialization constant in our proposed ToBias initialization, and the effect of changing the number of ensemble members. The study is conducted on the ChaosNLI dataset.
C.2.1 Tobias Initialization
We study the impact of the ToBias initialization constant $\beta$ on the coverage and efficiency of the resulting credal predictor on the ChaosNLI dataset. To do so, we vary the constant, taking values $\beta\in\{5,10,20,30,50,80,100,200,500\}$. The ensemble is then trained as proposed in Algorithm 1. The coverage and efficiency Pareto front is shown in Figure 5 with the mean and standard deviation over three runs. We split the results into three Figures for better readability.
Figure 5: Pareto front between coverage and efficiency for different values of ToBias constant $\beta$. Left: low values of $\beta$, middle: high values of $\beta$, right: all values of $\beta$.
For low values of $\beta$, the coverage of the ensembles is rather low and as $\beta$ increases, the coverage also increases. For large values of $\beta$, e.g. $\beta=200$ and $\beta=500$, the ensembles with large $\alpha$ values are no longer able to reach into the low coverage, high efficiency region. In particular, $\beta=500$ has a higher coverage and lower efficiency for larger values of $\alpha$ than for smaller values of $\alpha$. This may caused by converge problems due to having a very large bias for a particular class. This causes the individual ensemble members to not always converge to their desired threshold, hence resulting in large credal sets (with high coverage and low efficiency), because the border of the credal set is not reached. In essence, the $\beta$ value provides the ability to slightly shift the Pareto front to reach the desired coverage, efficiency region (in addition to $\alpha$).
C.2.2 Number of Ensemble Members
We study the impact of the numbers of ensemble members $M$ on the coverage and efficiency of the resulting credal predictor on the ChaosNLI dataset. To do so, we vary the constant, taking values $M\in\{1,2,3,4,5,10,20,30,50\}$. The ensemble is then trained as proposed in Algorithm 1. The coverage and efficiency Pareto front is shown in Figure 6 with the mean and standard deviation over three runs. We split the results into three Figures for better readability.
Figure 6: Pareto front between coverage and efficiency for different number of ensemble members $M$. Left: low values of $M$, middle: high values of $M$, right: all values of $M$.
Naturally, when $M=1$, the coverage is 0 and efficiency 1, because the credal set reduces to a point prediction. With $M=2$ and $M=3$, the $\alpha=0$ ensemble does not have coverage 1 and efficiency 0, because there are not enough ensemble members, and hence probability distribution, to span the whole probability simplex for this 3 class problem. As $M$ increases, the Pareto front starts to span more of the Pareto figure, until stabilizing around $M=20$. After this, the number of ensemble members does not have a significant impact on the resulting Pareto front anymore.
Table 5: Average entropy of ground-truth distribution in the test sets of ChaosNLI, CIFAR-10 and QualityMRI.
Dataset Avg. Entropy
ChaosNLI $0.932$
CIFAR-10 $0.223$
QualityMRI $0.782$
C.3 Data Uncertainty
The datasets used in our experiments, namely ChaosNLI, CIFAR-10, and QualityMRI, differ in their inherent levels of (aleatoric) uncertainty. This is reflected, for example, in the degree of disagreement among the collected annotations, which serve as the ground-truth distributions in our evaluations. To provide further insight, we report the average entropy of these ground-truth distributions in Table 5.
C.4 Performance of Ensemble Members
We visualized the trade-off between coverage and efficiency in Figure 4, and provide the corresponding numerical values in Table 6. We also report the accuracies of the individual predictors within each ensemble for both the baselines and our proposed method to provide insight into their standalone performance (cf. Tables 7 and 8). For CreNet, whose ensemble members directly predict probability intervals, we use their intersection probability technique to derive pointwise predictions [Wang et al., 2024a].
Table 6: Coverage and efficiency for different methods across datasets.
ChaosNLI CIFAR-10 QualityMRI
Approach Coverage Efficiency Coverage Efficiency Coverage Efficiency
CreRL0.0 $1.000\scriptstyle{\pm 0.000}$ $0.000\scriptstyle{\pm 0.000}$ $1.000\scriptstyle{\pm 0.000}$ $0.000\scriptstyle{\pm 0.000}$ $1.000\scriptstyle{\pm 0.000}$ $0.000\scriptstyle{\pm 0.000}$
CreRL0.2 $0.851\scriptstyle{\pm 0.011}$ $0.312\scriptstyle{\pm 0.025}$ $0.754\scriptstyle{\pm 0.005}$ $0.766\scriptstyle{\pm 0.003}$ $0.796\scriptstyle{\pm 0.068}$ $0.261\scriptstyle{\pm 0.053}$
CreRL0.4 $0.707\scriptstyle{\pm 0.007}$ $0.358\scriptstyle{\pm 0.013}$ $0.680\scriptstyle{\pm 0.004}$ $0.824\scriptstyle{\pm 0.004}$ $0.747\scriptstyle{\pm 0.033}$ $0.293\scriptstyle{\pm 0.010}$
CreRL0.6 $0.559\scriptstyle{\pm 0.034}$ $0.488\scriptstyle{\pm 0.012}$ $0.634\scriptstyle{\pm 0.004}$ $0.862\scriptstyle{\pm 0.004}$ $0.677\scriptstyle{\pm 0.023}$ $0.360\scriptstyle{\pm 0.016}$
CreRL0.8 $0.410\scriptstyle{\pm 0.021}$ $0.631\scriptstyle{\pm 0.022}$ $0.581\scriptstyle{\pm 0.005}$ $0.894\scriptstyle{\pm 0.003}$ $0.613\scriptstyle{\pm 0.035}$ $0.417\scriptstyle{\pm 0.025}$
CreRL0.9 $0.294\scriptstyle{\pm 0.007}$ $0.716\scriptstyle{\pm 0.010}$ $0.556\scriptstyle{\pm 0.001}$ $0.911\scriptstyle{\pm 0.001}$ $0.548\scriptstyle{\pm 0.046}$ $0.475\scriptstyle{\pm 0.038}$
CreRL0.95 $0.260\scriptstyle{\pm 0.022}$ $0.743\scriptstyle{\pm 0.012}$ $0.543\scriptstyle{\pm 0.002}$ $0.920\scriptstyle{\pm 0.001}$ $0.511\scriptstyle{\pm 0.046}$ $0.508\scriptstyle{\pm 0.037}$
CreRL1.0 $0.246\scriptstyle{\pm 0.012}$ $0.757\scriptstyle{\pm 0.008}$ $0.498\scriptstyle{\pm 0.001}$ $0.950\scriptstyle{\pm 0.000}$ $0.500\scriptstyle{\pm 0.086}$ $0.526\scriptstyle{\pm 0.036}$
CreWra $0.453\scriptstyle{\pm 0.050}$ $0.607\scriptstyle{\pm 0.037}$ $0.450\scriptstyle{\pm 0.001}$ $0.963\scriptstyle{\pm 0.000}$ $0.355\scriptstyle{\pm 0.057}$ $0.608\scriptstyle{\pm 0.021}$
CreNet $0.001\scriptstyle{\pm 0.002}$ $0.978\scriptstyle{\pm 0.008}$ $0.094\scriptstyle{\pm 0.010}$ $0.999\scriptstyle{\pm 0.000}$ $0.188\scriptstyle{\pm 0.020}$ $0.792\scriptstyle{\pm 0.003}$
CreBNN $0.195\scriptstyle{\pm 0.019}$ $0.648\scriptstyle{\pm 0.019}$ $0.331\scriptstyle{\pm 0.004}$ $0.871\scriptstyle{\pm 0.003}$ $0.780\scriptstyle{\pm 0.312}$ $0.184\scriptstyle{\pm 0.255}$
CreEns0.0 $0.298\scriptstyle{\pm 0.041}$ $0.607\scriptstyle{\pm 0.037}$ $0.000\scriptstyle{\pm 0.000}$ $0.963\scriptstyle{\pm 0.000}$ $0.329\scriptstyle{\pm 0.040}$ $0.611\scriptstyle{\pm 0.023}$
CreEns0.2 $0.165\scriptstyle{\pm 0.007}$ $0.769\scriptstyle{\pm 0.010}$ $0.000\scriptstyle{\pm 0.000}$ $0.983\scriptstyle{\pm 0.000}$ $0.177\scriptstyle{\pm 0.060}$ $0.787\scriptstyle{\pm 0.020}$
CreEns0.4 $0.096\scriptstyle{\pm 0.008}$ $0.834\scriptstyle{\pm 0.006}$ $0.000\scriptstyle{\pm 0.000}$ $0.986\scriptstyle{\pm 0.000}$ $0.118\scriptstyle{\pm 0.059}$ $0.844\scriptstyle{\pm 0.021}$
CreEns0.6 $0.043\scriptstyle{\pm 0.003}$ $0.886\scriptstyle{\pm 0.005}$ $0.000\scriptstyle{\pm 0.000}$ $0.988\scriptstyle{\pm 0.000}$ $0.081\scriptstyle{\pm 0.057}$ $0.899\scriptstyle{\pm 0.012}$
CreEns0.8 $0.012\scriptstyle{\pm 0.002}$ $0.941\scriptstyle{\pm 0.002}$ $0.000\scriptstyle{\pm 0.000}$ $0.993\scriptstyle{\pm 0.000}$ $0.054\scriptstyle{\pm 0.033}$ $0.950\scriptstyle{\pm 0.008}$
CreEns0.9 $0.000\scriptstyle{\pm 0.000}$ $0.976\scriptstyle{\pm 0.001}$ $0.000\scriptstyle{\pm 0.000}$ $0.997\scriptstyle{\pm 0.000}$ $0.016\scriptstyle{\pm 0.023}$ $0.983\scriptstyle{\pm 0.002}$
CreEns0.95 $0.000\scriptstyle{\pm 0.000}$ $1.000\scriptstyle{\pm 0.000}$ $0.000\scriptstyle{\pm 0.000}$ $1.000\scriptstyle{\pm 0.000}$ $0.000\scriptstyle{\pm 0.000}$ $1.000\scriptstyle{\pm 0.000}$
Table 7: Accuracy per ensemble member of CreRLα for varying $\alpha$ values on ChaosNLI, CIFAR-10, and QualityMRI.
ChaosNLI $\bm{\alpha}$ $\bm{0.0}$ $\bm{0.2}$ $\bm{0.4}$ $\bm{0.6}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{0.95}$ $\bm{1.0}$ Member $\bm{m}$ 1 $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ 2 $0.370\scriptstyle{\pm 0.012}$ $0.526\scriptstyle{\pm 0.075}$ $0.512\scriptstyle{\pm 0.063}$ $0.577\scriptstyle{\pm 0.011}$ $0.608\scriptstyle{\pm 0.055}$ $0.664\scriptstyle{\pm 0.028}$ $0.682\scriptstyle{\pm 0.016}$ $0.678\scriptstyle{\pm 0.019}$ 3 $0.453\scriptstyle{\pm 0.010}$ $0.549\scriptstyle{\pm 0.008}$ $0.407\scriptstyle{\pm 0.020}$ $0.604\scriptstyle{\pm 0.045}$ $0.634\scriptstyle{\pm 0.016}$ $0.668\scriptstyle{\pm 0.003}$ $0.663\scriptstyle{\pm 0.027}$ $0.657\scriptstyle{\pm 0.020}$ 4 $0.178\scriptstyle{\pm 0.014}$ $0.468\scriptstyle{\pm 0.022}$ $0.483\scriptstyle{\pm 0.045}$ $0.483\scriptstyle{\pm 0.046}$ $0.651\scriptstyle{\pm 0.016}$ $0.660\scriptstyle{\pm 0.012}$ $0.661\scriptstyle{\pm 0.010}$ $0.667\scriptstyle{\pm 0.006}$ 5 $0.370\scriptstyle{\pm 0.012}$ $0.568\scriptstyle{\pm 0.018}$ $0.425\scriptstyle{\pm 0.046}$ $0.560\scriptstyle{\pm 0.042}$ $0.624\scriptstyle{\pm 0.018}$ $0.656\scriptstyle{\pm 0.016}$ $0.677\scriptstyle{\pm 0.017}$ $0.686\scriptstyle{\pm 0.020}$ 6 $0.453\scriptstyle{\pm 0.010}$ $0.535\scriptstyle{\pm 0.015}$ $0.530\scriptstyle{\pm 0.023}$ $0.585\scriptstyle{\pm 0.025}$ $0.647\scriptstyle{\pm 0.017}$ $0.663\scriptstyle{\pm 0.012}$ $0.674\scriptstyle{\pm 0.007}$ $0.676\scriptstyle{\pm 0.006}$ 7 $0.178\scriptstyle{\pm 0.014}$ $0.453\scriptstyle{\pm 0.015}$ $0.506\scriptstyle{\pm 0.041}$ $0.594\scriptstyle{\pm 0.044}$ $0.637\scriptstyle{\pm 0.020}$ $0.647\scriptstyle{\pm 0.003}$ $0.664\scriptstyle{\pm 0.005}$ $0.675\scriptstyle{\pm 0.006}$ 8 $0.370\scriptstyle{\pm 0.012}$ $0.585\scriptstyle{\pm 0.027}$ $0.494\scriptstyle{\pm 0.119}$ $0.585\scriptstyle{\pm 0.039}$ $0.646\scriptstyle{\pm 0.038}$ $0.673\scriptstyle{\pm 0.022}$ $0.677\scriptstyle{\pm 0.024}$ $0.682\scriptstyle{\pm 0.014}$ 9 $0.453\scriptstyle{\pm 0.010}$ $0.493\scriptstyle{\pm 0.074}$ $0.595\scriptstyle{\pm 0.042}$ $0.604\scriptstyle{\pm 0.029}$ $0.663\scriptstyle{\pm 0.007}$ $0.668\scriptstyle{\pm 0.002}$ $0.661\scriptstyle{\pm 0.023}$ $0.670\scriptstyle{\pm 0.009}$ 10 $0.178\scriptstyle{\pm 0.014}$ $0.476\scriptstyle{\pm 0.022}$ $0.514\scriptstyle{\pm 0.050}$ $0.552\scriptstyle{\pm 0.054}$ $0.650\scriptstyle{\pm 0.004}$ $0.648\scriptstyle{\pm 0.008}$ $0.664\scriptstyle{\pm 0.013}$ $0.675\scriptstyle{\pm 0.008}$ 11 $0.370\scriptstyle{\pm 0.012}$ $0.533\scriptstyle{\pm 0.041}$ $0.584\scriptstyle{\pm 0.046}$ $0.587\scriptstyle{\pm 0.035}$ $0.676\scriptstyle{\pm 0.029}$ $0.684\scriptstyle{\pm 0.018}$ $0.684\scriptstyle{\pm 0.012}$ $0.683\scriptstyle{\pm 0.017}$ 12 $0.453\scriptstyle{\pm 0.010}$ $0.644\scriptstyle{\pm 0.011}$ $0.638\scriptstyle{\pm 0.015}$ $0.650\scriptstyle{\pm 0.011}$ $0.659\scriptstyle{\pm 0.011}$ $0.670\scriptstyle{\pm 0.013}$ $0.672\scriptstyle{\pm 0.005}$ $0.665\scriptstyle{\pm 0.019}$ 13 $0.178\scriptstyle{\pm 0.014}$ $0.510\scriptstyle{\pm 0.065}$ $0.584\scriptstyle{\pm 0.079}$ $0.633\scriptstyle{\pm 0.021}$ $0.652\scriptstyle{\pm 0.009}$ $0.664\scriptstyle{\pm 0.006}$ $0.668\scriptstyle{\pm 0.010}$ $0.670\scriptstyle{\pm 0.014}$ 14 $0.370\scriptstyle{\pm 0.012}$ $0.560\scriptstyle{\pm 0.040}$ $0.603\scriptstyle{\pm 0.015}$ $0.632\scriptstyle{\pm 0.031}$ $0.668\scriptstyle{\pm 0.020}$ $0.675\scriptstyle{\pm 0.014}$ $0.684\scriptstyle{\pm 0.014}$ $0.676\scriptstyle{\pm 0.011}$ 15 $0.453\scriptstyle{\pm 0.010}$ $0.644\scriptstyle{\pm 0.019}$ $0.643\scriptstyle{\pm 0.013}$ $0.662\scriptstyle{\pm 0.017}$ $0.679\scriptstyle{\pm 0.014}$ $0.685\scriptstyle{\pm 0.013}$ $0.677\scriptstyle{\pm 0.007}$ $0.670\scriptstyle{\pm 0.019}$ 16 $0.178\scriptstyle{\pm 0.014}$ $0.630\scriptstyle{\pm 0.021}$ $0.637\scriptstyle{\pm 0.001}$ $0.647\scriptstyle{\pm 0.013}$ $0.651\scriptstyle{\pm 0.005}$ $0.668\scriptstyle{\pm 0.017}$ $0.658\scriptstyle{\pm 0.007}$ $0.671\scriptstyle{\pm 0.002}$ 17 $0.370\scriptstyle{\pm 0.012}$ $0.623\scriptstyle{\pm 0.045}$ $0.647\scriptstyle{\pm 0.017}$ $0.655\scriptstyle{\pm 0.010}$ $0.684\scriptstyle{\pm 0.014}$ $0.688\scriptstyle{\pm 0.024}$ $0.685\scriptstyle{\pm 0.020}$ $0.687\scriptstyle{\pm 0.010}$ 18 $0.453\scriptstyle{\pm 0.010}$ $0.652\scriptstyle{\pm 0.005}$ $0.655\scriptstyle{\pm 0.009}$ $0.664\scriptstyle{\pm 0.014}$ $0.684\scriptstyle{\pm 0.008}$ $0.669\scriptstyle{\pm 0.011}$ $0.663\scriptstyle{\pm 0.009}$ $0.663\scriptstyle{\pm 0.005}$ 19 $0.178\scriptstyle{\pm 0.014}$ $0.650\scriptstyle{\pm 0.008}$ $0.659\scriptstyle{\pm 0.003}$ $0.663\scriptstyle{\pm 0.015}$ $0.673\scriptstyle{\pm 0.006}$ $0.676\scriptstyle{\pm 0.005}$ $0.667\scriptstyle{\pm 0.001}$ $0.667\scriptstyle{\pm 0.010}$ 20 $0.370\scriptstyle{\pm 0.012}$ $0.689\scriptstyle{\pm 0.017}$ $0.677\scriptstyle{\pm 0.014}$ $0.679\scriptstyle{\pm 0.024}$ $0.684\scriptstyle{\pm 0.029}$ $0.688\scriptstyle{\pm 0.024}$ $0.671\scriptstyle{\pm 0.026}$ $0.673\scriptstyle{\pm 0.033}$ CIFAR-10 $\bm{\alpha}$ $\bm{0.0}$ $\bm{0.2}$ $\bm{0.4}$ $\bm{0.6}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{0.95}$ $\bm{1.0}$ Member $\bm{m}$ 1 $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.943\scriptstyle{\pm 0.001}$ 2 $0.100\scriptstyle{\pm 0.000}$ $0.490\scriptstyle{\pm 0.027}$ $0.711\scriptstyle{\pm 0.030}$ $0.792\scriptstyle{\pm 0.014}$ $0.860\scriptstyle{\pm 0.003}$ $0.889\scriptstyle{\pm 0.003}$ $0.898\scriptstyle{\pm 0.005}$ $0.934\scriptstyle{\pm 0.001}$ 3 $0.100\scriptstyle{\pm 0.000}$ $0.582\scriptstyle{\pm 0.043}$ $0.725\scriptstyle{\pm 0.027}$ $0.809\scriptstyle{\pm 0.004}$ $0.873\scriptstyle{\pm 0.008}$ $0.891\scriptstyle{\pm 0.002}$ $0.898\scriptstyle{\pm 0.004}$ $0.934\scriptstyle{\pm 0.002}$ 4 $0.100\scriptstyle{\pm 0.000}$ $0.591\scriptstyle{\pm 0.009}$ $0.722\scriptstyle{\pm 0.008}$ $0.816\scriptstyle{\pm 0.009}$ $0.871\scriptstyle{\pm 0.008}$ $0.891\scriptstyle{\pm 0.000}$ $0.902\scriptstyle{\pm 0.001}$ $0.936\scriptstyle{\pm 0.001}$ 5 $0.099\scriptstyle{\pm 0.000}$ $0.645\scriptstyle{\pm 0.036}$ $0.754\scriptstyle{\pm 0.012}$ $0.837\scriptstyle{\pm 0.011}$ $0.874\scriptstyle{\pm 0.002}$ $0.886\scriptstyle{\pm 0.004}$ $0.906\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.002}$ 6 $0.098\scriptstyle{\pm 0.000}$ $0.704\scriptstyle{\pm 0.010}$ $0.775\scriptstyle{\pm 0.002}$ $0.843\scriptstyle{\pm 0.012}$ $0.879\scriptstyle{\pm 0.006}$ $0.891\scriptstyle{\pm 0.003}$ $0.903\scriptstyle{\pm 0.003}$ $0.934\scriptstyle{\pm 0.002}$ 7 $0.100\scriptstyle{\pm 0.000}$ $0.697\scriptstyle{\pm 0.012}$ $0.782\scriptstyle{\pm 0.004}$ $0.851\scriptstyle{\pm 0.008}$ $0.881\scriptstyle{\pm 0.003}$ $0.893\scriptstyle{\pm 0.006}$ $0.902\scriptstyle{\pm 0.002}$ $0.936\scriptstyle{\pm 0.001}$ 8 $0.100\scriptstyle{\pm 0.000}$ $0.748\scriptstyle{\pm 0.015}$ $0.801\scriptstyle{\pm 0.011}$ $0.850\scriptstyle{\pm 0.006}$ $0.880\scriptstyle{\pm 0.006}$ $0.897\scriptstyle{\pm 0.003}$ $0.906\scriptstyle{\pm 0.004}$ $0.935\scriptstyle{\pm 0.001}$ 9 $0.101\scriptstyle{\pm 0.000}$ $0.765\scriptstyle{\pm 0.025}$ $0.813\scriptstyle{\pm 0.020}$ $0.848\scriptstyle{\pm 0.004}$ $0.890\scriptstyle{\pm 0.004}$ $0.896\scriptstyle{\pm 0.005}$ $0.908\scriptstyle{\pm 0.001}$ $0.936\scriptstyle{\pm 0.002}$ 10 $0.100\scriptstyle{\pm 0.000}$ $0.779\scriptstyle{\pm 0.007}$ $0.830\scriptstyle{\pm 0.007}$ $0.862\scriptstyle{\pm 0.011}$ $0.886\scriptstyle{\pm 0.012}$ $0.895\scriptstyle{\pm 0.004}$ $0.904\scriptstyle{\pm 0.003}$ $0.936\scriptstyle{\pm 0.001}$ 11 $0.100\scriptstyle{\pm 0.000}$ $0.790\scriptstyle{\pm 0.003}$ $0.831\scriptstyle{\pm 0.003}$ $0.865\scriptstyle{\pm 0.005}$ $0.889\scriptstyle{\pm 0.007}$ $0.904\scriptstyle{\pm 0.001}$ $0.908\scriptstyle{\pm 0.004}$ $0.936\scriptstyle{\pm 0.001}$ 12 $0.100\scriptstyle{\pm 0.000}$ $0.821\scriptstyle{\pm 0.009}$ $0.854\scriptstyle{\pm 0.013}$ $0.863\scriptstyle{\pm 0.008}$ $0.890\scriptstyle{\pm 0.002}$ $0.902\scriptstyle{\pm 0.004}$ $0.907\scriptstyle{\pm 0.003}$ $0.934\scriptstyle{\pm 0.002}$ 13 $0.100\scriptstyle{\pm 0.000}$ $0.824\scriptstyle{\pm 0.001}$ $0.850\scriptstyle{\pm 0.002}$ $0.875\scriptstyle{\pm 0.003}$ $0.884\scriptstyle{\pm 0.009}$ $0.904\scriptstyle{\pm 0.003}$ $0.907\scriptstyle{\pm 0.002}$ $0.937\scriptstyle{\pm 0.002}$ 14 $0.100\scriptstyle{\pm 0.000}$ $0.840\scriptstyle{\pm 0.002}$ $0.860\scriptstyle{\pm 0.008}$ $0.874\scriptstyle{\pm 0.009}$ $0.891\scriptstyle{\pm 0.002}$ $0.906\scriptstyle{\pm 0.006}$ $0.912\scriptstyle{\pm 0.004}$ $0.934\scriptstyle{\pm 0.002}$ 15 $0.099\scriptstyle{\pm 0.000}$ $0.849\scriptstyle{\pm 0.007}$ $0.868\scriptstyle{\pm 0.006}$ $0.879\scriptstyle{\pm 0.004}$ $0.898\scriptstyle{\pm 0.003}$ $0.905\scriptstyle{\pm 0.001}$ $0.911\scriptstyle{\pm 0.000}$ $0.933\scriptstyle{\pm 0.002}$ 16 $0.098\scriptstyle{\pm 0.000}$ $0.868\scriptstyle{\pm 0.009}$ $0.873\scriptstyle{\pm 0.006}$ $0.887\scriptstyle{\pm 0.004}$ $0.904\scriptstyle{\pm 0.004}$ $0.910\scriptstyle{\pm 0.003}$ $0.911\scriptstyle{\pm 0.002}$ $0.936\scriptstyle{\pm 0.002}$ 17 $0.100\scriptstyle{\pm 0.000}$ $0.863\scriptstyle{\pm 0.003}$ $0.879\scriptstyle{\pm 0.003}$ $0.888\scriptstyle{\pm 0.004}$ $0.901\scriptstyle{\pm 0.005}$ $0.908\scriptstyle{\pm 0.002}$ $0.909\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.001}$ 18 $0.100\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.004}$ $0.892\scriptstyle{\pm 0.005}$ $0.898\scriptstyle{\pm 0.002}$ $0.902\scriptstyle{\pm 0.005}$ $0.913\scriptstyle{\pm 0.005}$ $0.916\scriptstyle{\pm 0.001}$ $0.934\scriptstyle{\pm 0.001}$ 19 $0.101\scriptstyle{\pm 0.000}$ $0.887\scriptstyle{\pm 0.007}$ $0.897\scriptstyle{\pm 0.004}$ $0.902\scriptstyle{\pm 0.001}$ $0.909\scriptstyle{\pm 0.002}$ $0.915\scriptstyle{\pm 0.002}$ $0.915\scriptstyle{\pm 0.001}$ $0.936\scriptstyle{\pm 0.001}$ 20 $0.100\scriptstyle{\pm 0.000}$ $0.903\scriptstyle{\pm 0.005}$ $0.904\scriptstyle{\pm 0.003}$ $0.908\scriptstyle{\pm 0.002}$ $0.910\scriptstyle{\pm 0.004}$ $0.920\scriptstyle{\pm 0.002}$ $0.921\scriptstyle{\pm 0.001}$ $0.935\scriptstyle{\pm 0.002}$ QualityMRI $\bm{\alpha}$ $\bm{0.0}$ $\bm{0.2}$ $\bm{0.4}$ $\bm{0.6}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{0.95}$ $\bm{1.0}$ Member $\bm{m}$ 1 $0.624\scriptstyle{\pm 0.033}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.624\scriptstyle{\pm 0.033}$ $0.532\scriptstyle{\pm 0.105}$ 2 $0.376\scriptstyle{\pm 0.008}$ $0.511\scriptstyle{\pm 0.050}$ $0.532\scriptstyle{\pm 0.023}$ $0.570\scriptstyle{\pm 0.046}$ $0.575\scriptstyle{\pm 0.020}$ $0.597\scriptstyle{\pm 0.057}$ $0.586\scriptstyle{\pm 0.055}$ $0.495\scriptstyle{\pm 0.119}$ 3 $0.624\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.015}$ $0.672\scriptstyle{\pm 0.027}$ $0.629\scriptstyle{\pm 0.026}$ $0.618\scriptstyle{\pm 0.015}$ $0.624\scriptstyle{\pm 0.055}$ $0.656\scriptstyle{\pm 0.042}$ $0.570\scriptstyle{\pm 0.075}$ 4 $0.376\scriptstyle{\pm 0.008}$ $0.586\scriptstyle{\pm 0.015}$ $0.548\scriptstyle{\pm 0.103}$ $0.543\scriptstyle{\pm 0.020}$ $0.581\scriptstyle{\pm 0.013}$ $0.618\scriptstyle{\pm 0.027}$ $0.618\scriptstyle{\pm 0.053}$ $0.511\scriptstyle{\pm 0.112}$ 5 $0.624\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.027}$ $0.645\scriptstyle{\pm 0.035}$ $0.602\scriptstyle{\pm 0.042}$ $0.597\scriptstyle{\pm 0.070}$ $0.629\scriptstyle{\pm 0.013}$ $0.651\scriptstyle{\pm 0.038}$ $0.570\scriptstyle{\pm 0.101}$ 6 $0.376\scriptstyle{\pm 0.008}$ $0.543\scriptstyle{\pm 0.020}$ $0.586\scriptstyle{\pm 0.040}$ $0.581\scriptstyle{\pm 0.023}$ $0.591\scriptstyle{\pm 0.020}$ $0.570\scriptstyle{\pm 0.008}$ $0.597\scriptstyle{\pm 0.035}$ $0.532\scriptstyle{\pm 0.103}$ 7 $0.624\scriptstyle{\pm 0.008}$ $0.618\scriptstyle{\pm 0.042}$ $0.624\scriptstyle{\pm 0.020}$ $0.608\scriptstyle{\pm 0.020}$ $0.618\scriptstyle{\pm 0.030}$ $0.640\scriptstyle{\pm 0.042}$ $0.629\scriptstyle{\pm 0.013}$ $0.511\scriptstyle{\pm 0.124}$ 8 $0.376\scriptstyle{\pm 0.008}$ $0.597\scriptstyle{\pm 0.099}$ $0.586\scriptstyle{\pm 0.027}$ $0.575\scriptstyle{\pm 0.046}$ $0.597\scriptstyle{\pm 0.060}$ $0.591\scriptstyle{\pm 0.027}$ $0.602\scriptstyle{\pm 0.020}$ $0.500\scriptstyle{\pm 0.115}$ 9 $0.624\scriptstyle{\pm 0.008}$ $0.640\scriptstyle{\pm 0.020}$ $0.651\scriptstyle{\pm 0.055}$ $0.683\scriptstyle{\pm 0.046}$ $0.613\scriptstyle{\pm 0.023}$ $0.645\scriptstyle{\pm 0.013}$ $0.629\scriptstyle{\pm 0.013}$ $0.522\scriptstyle{\pm 0.107}$ 10 $0.376\scriptstyle{\pm 0.008}$ $0.565\scriptstyle{\pm 0.035}$ $0.602\scriptstyle{\pm 0.008}$ $0.548\scriptstyle{\pm 0.035}$ $0.602\scriptstyle{\pm 0.059}$ $0.591\scriptstyle{\pm 0.038}$ $0.570\scriptstyle{\pm 0.008}$ $0.505\scriptstyle{\pm 0.095}$ 11 $0.624\scriptstyle{\pm 0.008}$ $0.629\scriptstyle{\pm 0.026}$ $0.624\scriptstyle{\pm 0.020}$ $0.608\scriptstyle{\pm 0.027}$ $0.591\scriptstyle{\pm 0.033}$ $0.613\scriptstyle{\pm 0.066}$ $0.581\scriptstyle{\pm 0.013}$ $0.565\scriptstyle{\pm 0.068}$ 12 $0.376\scriptstyle{\pm 0.008}$ $0.538\scriptstyle{\pm 0.015}$ $0.570\scriptstyle{\pm 0.055}$ $0.640\scriptstyle{\pm 0.008}$ $0.591\scriptstyle{\pm 0.008}$ $0.602\scriptstyle{\pm 0.038}$ $0.570\scriptstyle{\pm 0.030}$ $0.538\scriptstyle{\pm 0.107}$ 13 $0.624\scriptstyle{\pm 0.008}$ $0.688\scriptstyle{\pm 0.053}$ $0.613\scriptstyle{\pm 0.013}$ $0.677\scriptstyle{\pm 0.023}$ $0.656\scriptstyle{\pm 0.008}$ $0.602\scriptstyle{\pm 0.042}$ $0.618\scriptstyle{\pm 0.042}$ $0.559\scriptstyle{\pm 0.100}$ 14 $0.376\scriptstyle{\pm 0.008}$ $0.532\scriptstyle{\pm 0.035}$ $0.581\scriptstyle{\pm 0.035}$ $0.570\scriptstyle{\pm 0.046}$ $0.586\scriptstyle{\pm 0.020}$ $0.591\scriptstyle{\pm 0.020}$ $0.554\scriptstyle{\pm 0.033}$ $0.538\scriptstyle{\pm 0.127}$ 15 $0.624\scriptstyle{\pm 0.008}$ $0.634\scriptstyle{\pm 0.040}$ $0.651\scriptstyle{\pm 0.008}$ $0.586\scriptstyle{\pm 0.050}$ $0.651\scriptstyle{\pm 0.059}$ $0.629\scriptstyle{\pm 0.023}$ $0.618\scriptstyle{\pm 0.050}$ $0.559\scriptstyle{\pm 0.112}$ 16 $0.376\scriptstyle{\pm 0.008}$ $0.565\scriptstyle{\pm 0.023}$ $0.618\scriptstyle{\pm 0.040}$ $0.586\scriptstyle{\pm 0.062}$ $0.570\scriptstyle{\pm 0.030}$ $0.597\scriptstyle{\pm 0.026}$ $0.554\scriptstyle{\pm 0.008}$ $0.516\scriptstyle{\pm 0.092}$ 17 $0.624\scriptstyle{\pm 0.008}$ $0.624\scriptstyle{\pm 0.040}$ $0.677\scriptstyle{\pm 0.035}$ $0.629\scriptstyle{\pm 0.047}$ $0.618\scriptstyle{\pm 0.027}$ $0.591\scriptstyle{\pm 0.053}$ $0.624\scriptstyle{\pm 0.040}$ $0.554\scriptstyle{\pm 0.066}$ 18 $0.376\scriptstyle{\pm 0.008}$ $0.575\scriptstyle{\pm 0.055}$ $0.586\scriptstyle{\pm 0.008}$ $0.597\scriptstyle{\pm 0.013}$ $0.591\scriptstyle{\pm 0.020}$ $0.570\scriptstyle{\pm 0.008}$ $0.602\scriptstyle{\pm 0.027}$ $0.522\scriptstyle{\pm 0.119}$ 19 $0.624\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.050}$ $0.656\scriptstyle{\pm 0.046}$ $0.634\scriptstyle{\pm 0.055}$ $0.629\scriptstyle{\pm 0.040}$ $0.624\scriptstyle{\pm 0.038}$ $0.618\scriptstyle{\pm 0.040}$ $0.527\scriptstyle{\pm 0.077}$ 20 $0.376\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.027}$ $0.570\scriptstyle{\pm 0.065}$ $0.608\scriptstyle{\pm 0.020}$ $0.597\scriptstyle{\pm 0.026}$ $0.608\scriptstyle{\pm 0.020}$ $0.554\scriptstyle{\pm 0.020}$ $0.548\scriptstyle{\pm 0.115}$
Table 8: Accuracy per ensemble member of baselines on ChaosNLI, CIFAR-10, and QualityMRI.
ChaosNLI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.658\scriptstyle{\pm 0.013}$ $0.671\scriptstyle{\pm 0.011}$ $0.512\scriptstyle{\pm 0.083}$ $0.525\scriptstyle{\pm 0.051}$ 2 $0.648\scriptstyle{\pm 0.012}$ $0.676\scriptstyle{\pm 0.013}$ $0.453\scriptstyle{\pm 0.010}$ $0.488\scriptstyle{\pm 0.098}$ 3 $0.641\scriptstyle{\pm 0.020}$ $0.676\scriptstyle{\pm 0.013}$ $0.506\scriptstyle{\pm 0.085}$ $0.542\scriptstyle{\pm 0.030}$ 4 $0.636\scriptstyle{\pm 0.012}$ $0.666\scriptstyle{\pm 0.008}$ $0.511\scriptstyle{\pm 0.092}$ $0.485\scriptstyle{\pm 0.037}$ 5 $0.651\scriptstyle{\pm 0.032}$ $0.667\scriptstyle{\pm 0.011}$ $0.532\scriptstyle{\pm 0.111}$ $0.554\scriptstyle{\pm 0.051}$ 6 $0.660\scriptstyle{\pm 0.016}$ $0.674\scriptstyle{\pm 0.021}$ $0.503\scriptstyle{\pm 0.081}$ $0.470\scriptstyle{\pm 0.115}$ 7 $0.637\scriptstyle{\pm 0.014}$ $0.677\scriptstyle{\pm 0.009}$ $0.453\scriptstyle{\pm 0.010}$ $0.518\scriptstyle{\pm 0.036}$ 8 $0.667\scriptstyle{\pm 0.016}$ $0.670\scriptstyle{\pm 0.010}$ $0.453\scriptstyle{\pm 0.010}$ $0.446\scriptstyle{\pm 0.030}$ 9 $0.584\scriptstyle{\pm 0.086}$ $0.663\scriptstyle{\pm 0.005}$ $0.453\scriptstyle{\pm 0.010}$ $0.462\scriptstyle{\pm 0.113}$ 10 $0.663\scriptstyle{\pm 0.035}$ $0.674\scriptstyle{\pm 0.029}$ $0.664\scriptstyle{\pm 0.012}$ $0.531\scriptstyle{\pm 0.064}$ 11 $0.648\scriptstyle{\pm 0.017}$ $0.653\scriptstyle{\pm 0.015}$ $0.591\scriptstyle{\pm 0.099}$ $0.484\scriptstyle{\pm 0.047}$ 12 $0.660\scriptstyle{\pm 0.016}$ $0.662\scriptstyle{\pm 0.016}$ $0.587\scriptstyle{\pm 0.097}$ $0.549\scriptstyle{\pm 0.071}$ 13 $0.664\scriptstyle{\pm 0.014}$ $0.655\scriptstyle{\pm 0.017}$ $0.581\scriptstyle{\pm 0.099}$ $0.487\scriptstyle{\pm 0.040}$ 14 $0.650\scriptstyle{\pm 0.011}$ $0.654\scriptstyle{\pm 0.006}$ $0.655\scriptstyle{\pm 0.007}$ $0.505\scriptstyle{\pm 0.039}$ 15 $0.635\scriptstyle{\pm 0.038}$ $0.651\scriptstyle{\pm 0.018}$ $0.602\scriptstyle{\pm 0.097}$ $0.458\scriptstyle{\pm 0.143}$ 16 $0.652\scriptstyle{\pm 0.026}$ $0.654\scriptstyle{\pm 0.007}$ $0.629\scriptstyle{\pm 0.046}$ $0.486\scriptstyle{\pm 0.012}$ 17 $0.628\scriptstyle{\pm 0.019}$ $0.633\scriptstyle{\pm 0.015}$ $0.587\scriptstyle{\pm 0.086}$ $0.518\scriptstyle{\pm 0.057}$ 18 $0.652\scriptstyle{\pm 0.009}$ $0.605\scriptstyle{\pm 0.021}$ $0.502\scriptstyle{\pm 0.079}$ $0.519\scriptstyle{\pm 0.031}$ 19 $0.660\scriptstyle{\pm 0.012}$ $0.581\scriptstyle{\pm 0.037}$ $0.584\scriptstyle{\pm 0.095}$ $0.487\scriptstyle{\pm 0.050}$ 20 $0.651\scriptstyle{\pm 0.013}$ $0.480\scriptstyle{\pm 0.074}$ $0.586\scriptstyle{\pm 0.095}$ $0.560\scriptstyle{\pm 0.043}$ CIFAR-10 CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.942\scriptstyle{\pm 0.001}$ $0.955\scriptstyle{\pm 0.000}$ $0.875\scriptstyle{\pm 0.004}$ $0.941\scriptstyle{\pm 0.000}$ 2 $0.944\scriptstyle{\pm 0.001}$ $0.954\scriptstyle{\pm 0.000}$ $0.877\scriptstyle{\pm 0.005}$ $0.943\scriptstyle{\pm 0.000}$ 3 $0.944\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.867\scriptstyle{\pm 0.015}$ $0.944\scriptstyle{\pm 0.001}$ 4 $0.943\scriptstyle{\pm 0.002}$ $0.952\scriptstyle{\pm 0.000}$ $0.881\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.001}$ 5 $0.943\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.004}$ $0.942\scriptstyle{\pm 0.001}$ 6 $0.943\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.000}$ $0.869\scriptstyle{\pm 0.006}$ $0.942\scriptstyle{\pm 0.002}$ 7 $0.941\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.877\scriptstyle{\pm 0.001}$ $0.943\scriptstyle{\pm 0.001}$ 8 $0.942\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.001}$ $0.880\scriptstyle{\pm 0.006}$ $0.943\scriptstyle{\pm 0.002}$ 9 $0.943\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.005}$ $0.944\scriptstyle{\pm 0.001}$ 10 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.873\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.003}$ 11 $0.942\scriptstyle{\pm 0.001}$ $0.953\scriptstyle{\pm 0.001}$ $0.882\scriptstyle{\pm 0.004}$ $0.943\scriptstyle{\pm 0.000}$ 12 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.000}$ 13 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.001}$ $0.874\scriptstyle{\pm 0.000}$ $0.943\scriptstyle{\pm 0.002}$ 14 $0.945\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.006}$ $0.941\scriptstyle{\pm 0.001}$ 15 $0.942\scriptstyle{\pm 0.002}$ $0.948\scriptstyle{\pm 0.000}$ $0.856\scriptstyle{\pm 0.019}$ $0.942\scriptstyle{\pm 0.001}$ 16 $0.942\scriptstyle{\pm 0.000}$ $0.941\scriptstyle{\pm 0.000}$ $0.870\scriptstyle{\pm 0.002}$ $0.942\scriptstyle{\pm 0.001}$ 17 $0.942\scriptstyle{\pm 0.001}$ $0.930\scriptstyle{\pm 0.001}$ $0.854\scriptstyle{\pm 0.035}$ $0.942\scriptstyle{\pm 0.002}$ 18 $0.944\scriptstyle{\pm 0.001}$ $0.921\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.002}$ 19 $0.943\scriptstyle{\pm 0.001}$ $0.906\scriptstyle{\pm 0.002}$ $0.873\scriptstyle{\pm 0.001}$ $0.942\scriptstyle{\pm 0.001}$ 20 $0.942\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.002}$ $0.860\scriptstyle{\pm 0.012}$ $0.942\scriptstyle{\pm 0.001}$ QualityMRI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.457\scriptstyle{\pm 0.099}$ $0.430\scriptstyle{\pm 0.170}$ $0.548\scriptstyle{\pm 0.115}$ $0.581\scriptstyle{\pm 0.137}$ 2 $0.452\scriptstyle{\pm 0.139}$ $0.409\scriptstyle{\pm 0.163}$ $0.484\scriptstyle{\pm 0.115}$ $0.559\scriptstyle{\pm 0.112}$ 3 $0.425\scriptstyle{\pm 0.124}$ $0.419\scriptstyle{\pm 0.160}$ $0.468\scriptstyle{\pm 0.126}$ $0.570\scriptstyle{\pm 0.110}$ 4 $0.419\scriptstyle{\pm 0.126}$ $0.446\scriptstyle{\pm 0.187}$ $0.548\scriptstyle{\pm 0.115}$ $0.575\scriptstyle{\pm 0.102}$ 5 $0.398\scriptstyle{\pm 0.135}$ $0.387\scriptstyle{\pm 0.139}$ $0.468\scriptstyle{\pm 0.126}$ $0.548\scriptstyle{\pm 0.126}$ 6 $0.403\scriptstyle{\pm 0.117}$ $0.398\scriptstyle{\pm 0.158}$ $0.554\scriptstyle{\pm 0.118}$ $0.554\scriptstyle{\pm 0.084}$ 7 $0.409\scriptstyle{\pm 0.175}$ $0.414\scriptstyle{\pm 0.161}$ $0.446\scriptstyle{\pm 0.107}$ $0.554\scriptstyle{\pm 0.095}$ 8 $0.430\scriptstyle{\pm 0.119}$ $0.441\scriptstyle{\pm 0.182}$ $0.554\scriptstyle{\pm 0.118}$ $0.559\scriptstyle{\pm 0.133}$ 9 $0.392\scriptstyle{\pm 0.137}$ $0.419\scriptstyle{\pm 0.181}$ $0.473\scriptstyle{\pm 0.110}$ $0.586\scriptstyle{\pm 0.141}$ 10 $0.446\scriptstyle{\pm 0.177}$ $0.430\scriptstyle{\pm 0.140}$ $0.462\scriptstyle{\pm 0.130}$ $0.570\scriptstyle{\pm 0.118}$ 11 $0.425\scriptstyle{\pm 0.137}$ $0.414\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.602\scriptstyle{\pm 0.107}$ 12 $0.398\scriptstyle{\pm 0.122}$ $0.430\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.103}$ 13 $0.398\scriptstyle{\pm 0.130}$ $0.430\scriptstyle{\pm 0.140}$ $0.446\scriptstyle{\pm 0.107}$ $0.575\scriptstyle{\pm 0.090}$ 14 $0.409\scriptstyle{\pm 0.066}$ $0.430\scriptstyle{\pm 0.142}$ $0.457\scriptstyle{\pm 0.122}$ $0.581\scriptstyle{\pm 0.070}$ 15 $0.409\scriptstyle{\pm 0.146}$ $0.414\scriptstyle{\pm 0.132}$ $0.468\scriptstyle{\pm 0.126}$ $0.624\scriptstyle{\pm 0.133}$ 16 $0.414\scriptstyle{\pm 0.107}$ $0.425\scriptstyle{\pm 0.167}$ $0.554\scriptstyle{\pm 0.118}$ $0.608\scriptstyle{\pm 0.100}$ 17 $0.419\scriptstyle{\pm 0.139}$ $0.414\scriptstyle{\pm 0.119}$ $0.554\scriptstyle{\pm 0.118}$ $0.581\scriptstyle{\pm 0.091}$ 18 $0.403\scriptstyle{\pm 0.168}$ $0.414\scriptstyle{\pm 0.129}$ $0.468\scriptstyle{\pm 0.126}$ $0.581\scriptstyle{\pm 0.117}$ 19 $0.425\scriptstyle{\pm 0.151}$ $0.414\scriptstyle{\pm 0.073}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.142}$ 20 $0.430\scriptstyle{\pm 0.154}$ $0.446\scriptstyle{\pm 0.068}$ $0.548\scriptstyle{\pm 0.117}$ $0.581\scriptstyle{\pm 0.103}$

### v2

@@
 1 Introduction
-The distinction between two types of uncertainty, referred to as aleatoric and epistemic, is receiving increasing interest in machine learning (Hüllermeier and Waegeman, 2021).
+The distinction between two types of uncertainty, referred to as aleatoric and epistemic, is receiving increasing interest in machine learning (hullermeier2021aleatoric).
 Roughly speaking, the aleatoric uncertainty of a predictive model is caused by the inherent randomness of the data-generating process, whereas epistemic uncertainty is caused by the learner’s lack of knowledge about the true (or best) predictive model.
 While aleatoric uncertainty is irreducible, epistemic uncertainty can be reduced on the basis of additional information, e.g., by collecting more training data.
+***equal contribution
 Aleatoric uncertainty can be captured adequately in terms of (conditional) probability distributions, whereas the representation of epistemic uncertainty requires “second-order” formalisms more general than (single) probability distributions.
-In the Bayesian approach to machine learning, second-order probability distributions are used for this purpose (Depeweg et al., 2018; Kendall and Gal, 2017): The learner maintains a probability distribution over the model space, which, if each model is a probabilistic predictor, results in predictions in the form of probability distributions of probability distributions (on outcomes).
-The Bayesian approach is theoretically appealing but computationally demanding, and commonly criticized for the need to specify a prior distribution (strongly influencing prediction and uncertainty quantification) (Gawlikowski et al., 2023).
+In the Bayesian approach to machine learning, second-order probability distributions are used for this purpose (depewegDecompositionUncertainty2018; kendall2017uncertainties): The learner maintains a probability distribution over the model space, which, if each model is a probabilistic predictor, results in predictions in the form of probability distributions of probability distributions (on outcomes).
+The Bayesian approach is theoretically appealing but computationally demanding, and commonly criticized for the need to specify a prior distribution (strongly influencing prediction and uncertainty quantification) (gawlikowskiSurveyUncertainty2023).
 Figure 1: Overview of our method.
 Language inference example from the ChaosNLI dataset, showing a premise-hypothesis pair with three possible labels: Entailed, Neutral and Contradicted.
-Middle: Illustration of the learning mechanism of our Credal Relative Likelihood (CreRLα) framework, showing ensemble member training up to different $\alpha$ thresholds.
+Middle: Illustration of the learning mechanism of our Credal Relative Likelihood (CreRL) framework, showing ensemble member training up to different $\alpha$ thresholds.
 Right: Unlike baselines, which are centered around the MLE and remain fixed, our method allows to adapt credal set size to improve coverage of the ground-truth distribution.
-An alternative second-order formalism is offered by credal sets (Walley, 1991), i.e., sets (instead of distributions) of probability distributions that are commonly assumed to be closed and convex (Cozman, 2000).
+An alternative second-order formalism is offered by credal sets (walleyStatisticalReasoning1991), i.e., sets (instead of distributions) of probability distributions that are commonly assumed to be closed and convex (cozman2000credal).
 Although a set appears to provide weaker information than a distribution, a set-based representation also has advantages.
-In particular, it can be argued that sets are more apt at representing ignorance in the sense of a lack of knowledge (Dubois et al., 1996), because distributions always involve additional assumptions beyond the mere distinction between plausible and implausible candidate models.
-Decision-making on the basis of credal sets has been proposed as a reasonable alternative to Bayesian decision-making (Levi, 1978; Girón and Ríos, 1980).
-The learning of credal predictors (i.e., models making predictions in the form of credal sets) has been considered in machine learning in the past (Zaffalon, 2001; Corani and Zaffalon, 2008) and, in light of the quest for proper representation of epistemic uncertainty, received renewed interest more recently (Wang et al., 2024a; Nguyen et al., 2025; Cella and Martin, 2024).
+In particular, it can be argued that sets are more apt at representing ignorance in the sense of a lack of knowledge (dubo_rp96), because distributions always involve additional assumptions beyond the mere distinction between plausible and implausible candidate models.
+Decision-making on the basis of credal sets has been proposed as a reasonable alternative to Bayesian decision-making (leviIndeterminateProbabilities1978; gironQuasiBayesian1980).
+The learning of credal predictors (i.e., models making predictions in the form of credal sets) has been considered in machine learning in the past (zaffalonStatisticalInference2001; coraniLearningReliable2008) and, in light of the quest for proper representation of epistemic uncertainty, received renewed interest more recently (wangCredalDeep2024; nguyenCredalEnsembling2025; cellaVariationalApproximations2024).
 A major question to be addressed in this regard concerns the distinction between plausible and implausible models.
@@
 Our approach explicitly defines the prediction target as the set of all probability distributions generated by plausible learners, namely those whose relative likelihood exceeds a specified threshold (cf.
-Section 3).
+Section˜3).
 That said, realizing this in complex machine learning settings comes with several practical challenges, notably the questions of representation, approximation, and inference: How to formally represent subsets of a high-dimensional model space, how to approximate such sets algorithmically, and how to infer credal predictions from them?
 We propose methods to address these questions based on suitably modified ensemble techniques (cf.
-Section 4).
+Section˜4).
 We highlight the effectiveness of our approach on real-world datasets, evaluating its performance in terms of coverage and efficiency as key criteria of credal prediction.
 Also, we analyze its performance on downstream tasks such as Out-of-Distribution (OoD) detection (cf.
-Section 5).
-Contributions. In summary our contributions are as follows:
+Section˜5).
+Contributions.
+In summary our contributions are as follows:
   • 1.
@@
 Related Work.
-In machine learning, uncertainty is often represented by (approximate) Bayesian methods (Mackay, 1992; Blundell et al., 2015; Lakshminarayanan et al., 2017; Gal and Ghahramani, 2016; Daxberger et al., 2021).
-An important characteristic of such representations, especially for uncertainty tasks, is diversity (D’Angelo and Fortuin, 2021; Wood et al., 2023), which can be enforced by regularization (de Mathelin et al., 2023), varying hyper-parameters (Wenzel et al., 2020), or increasing diversity in representations (Lopes et al., 2022).
-Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty (Zaffalon, 2001; Corani and Zaffalon, 2008; Corani and Mignatti, 2015).
-Such sets can be generated based on the relative likelihood, also referred to as normalized likelihood (Antonucci et al., 2012a) and have been used in machine learning with simple model classes (Senge et al., 2014; Cella and Martin, 2024).
+In machine learning, uncertainty is often represented by (approximate) Bayesian methods (mackayBayesianMethods1992; blundellWeightUncertainty2015; lakshminarayananDeepEnsembles2017; galDropoutAs2016; daxbergerLaplaceRedux2021).
+An important characteristic of such representations, especially for uncertainty tasks, is diversity (d2021repulsive; woodUnifiedTheory2023), which can be enforced by regularization (deMathelinDeepAnti2023), varying hyper-parameters (wenzelHyperparameterEnsembles2020), or increasing diversity in representations (lopesNoOne2022).
+Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty (zaffalonStatisticalInference2001; coraniLearningReliable2008; coraniCredalModel2015).
+Such sets can be generated based on the relative likelihood, also referred to as normalized likelihood (antonucciLikelihoodBased2012) and have been used in machine learning with simple model classes (sengeReliableClassification2014; cellaVariationalApproximations2024).
 In this work, we build upon these approaches and address the challenges that emerge when adapting the relative likelihood to a setting with complex predictors.
 Recently, credal sets have also been applied in the context of deep learning.
-Some approaches use ensemble learning to derive class-wise lower and upper probabilities (Wang et al., 2024b; Nguyen et al., 2025).
-Others train models to directly predict probability intervals (Wang et al., 2024a).
+Some approaches use ensemble learning to derive class-wise lower and upper probabilities (wangCredalWrapper2024; nguyenCredalEnsembling2025).
+Others train models to directly predict probability intervals (wangCredalDeep2024).
 Hybrid methods, combining multiple uncertainty frameworks, have also been proposed.
-Caprio et al.
-(2023) combine Bayesian deep learning and credal sets by considering sets of priors.
-Another approach leverages conformal prediction to construct credal sets with validity guarantees (Javanmardi et al., 2024).
-A more detailed discussion of the related work is provided in Appendix A.
+caprioImpreciseBayesian2023 combine Bayesian deep learning and credal sets by considering sets of priors.
+Another approach leverages conformal prediction to construct credal sets with validity guarantees (javanmardiConformalizedCredal2024).
+A more detailed discussion of the related work is provided in Appendix˜A.
 2 Problem Statement
@@
 The training data are realizations of random variables that are independently and identically distributed according to a probability measure on $\mathcal{X}\times\mathcal{Y}$.
-We consider a hypothesis space $\mathcal{H}$ with probabilistic predictors $h:\mathcal{X}\to\Delta_{K}$, where $\Delta_{K}$ denotes the $(K-1)$-simplex (that we will also refer to as the probability simplex).
+We consider a hypothesis space $\mathcal{H}$ with probabilistic predictors $h:\mathcal{X}\to\Delta_{K}$, where $\Delta_{K}$ denotes the (K1)(K-1)-simplex (that we will also refer to as the probability simplex).
 Given an instance $\bm{x}$, a predictor $h$ assigns a probability distribution $p(\cdot\,|\,\bm{x},h)=h(\bm{x})$, which is an estimate of the ground-truth probability $p(\cdot\,|\,\bm{x},h^{*})$ generated by the ground-truth model $h^{*}$.
@@
 In general, $H$ could also be realized differently.
-Inspired by other set-valued prediction methods such as conformal prediction (Vovk et al., 2005), we evaluate such a predictor in terms of its coverage and efficiency.
+Inspired by other set-valued prediction methods such as conformal prediction (vovk2005algorithmic), we evaluate such a predictor in terms of its coverage and efficiency.
 Coverage (of the ground-truth distribution by the cedal set) is defined as follows:
@@
 Importantly, we use efficiency strictly as an indicator of set size, not as a measure of epistemic uncertainty (cf.
-Section 2).
+Section˜2).
 The empirical coverage and efficiency is determined in terms of the corresponding averages on a finite set of (test) data $\bm{x}_{1},\ldots,\bm{x}_{N}$.
@@
 Here, $h^{ML}$ represents the model in $\mathcal{H}$ with the highest likelihood given the training data $\mathcal{D}$.
-This notion, also called extended or normalized likelihood, has been proposed by Birnbaum (1962) and used for statistical inference (Wasserman, 1990; Walley and Moral, 1999).
-It offers an attractive alternative to Bayesian or frequentist reasoning eliminating the need to specify priors or perform resampling (Giang and Shenoy, 2002).
+This notion, also called extended or normalized likelihood, has been proposed by birnbaumFoundationsStatistical1962 and used for statistical inference (wassermanBeliefFunctions1990; walleyUpperProbabilities1999).
+It offers an attractive alternative to Bayesian or frequentist reasoning eliminating the need to specify priors or perform resampling (giangStatisticalDecisions2002).
 In machine learning this is especially useful as specifying a meaningful prior on $\mathcal{H}$ is a non-trivial problem and resampling from the data distribution is usually not possible either, as we only have access to the given samples $\mathcal{D}$.
-An axiomatic justification of the relative likelihood (in the context of evidence theory) has been given by Denoeux (2014), who derives it from (i) the likelihood principle, (ii) compatibility with the Bayes rule, and (iii) the minimal commitment principle.
-On the basis of the relative likelihood, a set of models can be constructed, analogous to a confidence region (Aitkin, 1982), by including models that are plausible in the sense of surpassing a threshold $\alpha\in[0,1]$.
-This set of models is also referred to as an $\alpha$-cut (Antonucci et al., 2012a):
+An axiomatic justification of the relative likelihood (in the context of evidence theory) has been given by denoeuxLikelihoodBased2014, who derives it from (i) the likelihood principle, (ii) compatibility with the Bayes rule, and (iii) the minimal commitment principle.
 Figure 2: Relative Likelihood curve.
@@
 The credal set of $\alpha=0.9$ around the MLE at $0.6$ does not cover the ground-truth.
-Lowering $\alpha$ increases coverage at the cost of efficiency.
+Lowering $\alpha$ increases the coverage at the cost of efficiency.
+On the basis of the relative likelihood, a set of models can be constructed, analogous to a confidence region (aitkinDirectLikelihood1982), by including models that are plausible in the sense of surpassing a threshold $\alpha\in[0,1]$.
+This set of models is also referred to as an $\alpha$-cut (antonucciLikelihoodBased2012):
 $\mathcal{C}_{\alpha}=\{h:\gamma(h)\geq\alpha\}\subseteq\mathcal{H}.$ (4)
@@
 Thus, the parameter $\alpha$ has a quite intuitive meaning.
-The concept of $\alpha$-cuts is further illustrated in Figure 2.
+The concept of $\alpha$-cuts is illustrated in Figure˜2.
 Given an instance $\bm{x}\in\mathcal{X}$, the set of predictors maps to a set of (conditional) distributions:
@@
 At the same time, however, efficiency will deteriorate (cf.
-again Figure 2).
+again Figure˜2).
 As $\alpha$ increases, the intervals for given data points shrink since only models with high relative likelihood for the given data sample are included.
@@
 Throughout this work, we adopt the first approach and train the maximum likelihood estimator $h^{ML}$ using parameter configurations that have been identified in the literature as effective.
-Further details and configurations of experiments are provided in Appendix B.
+Further details and configurations of experiments are provided in Appendix˜B.
 Given an estimated maximum likelihood model, the estimated relative likelihood is obtained as $\hat{\gamma}(h)=\frac{L(h)}{L(h^{ML})}.$
 Algorithm 1 Train Credal Relative Likelihood ensemble.
-0:  $\alpha,M$
+[…]
 1:Step 1: Approximate maximum likelihood model
-2:  $h^{ML}\leftarrow\arg\min_{h}\mathcal{L}(h)$
+[…]
 3:Step 2: Train ensemble members
-4:  $\Delta\tau=\frac{1-\alpha}{M}$
+[…]
 5:for $i\in\{0,\ldots,M-1\}$ do
@@
 A straightforward approach is to train an ensemble of hypotheses and construct the $\alpha$-cut based on these hypotheses.
-In essence, approaches based on regular ensemble training (Wang et al., 2024b; Nguyen et al., 2025) form an example of this strategy.
-However, an issue is that all hypotheses $h$ typically tend to cluster around $h^{ML}$, hence the obtained hypotheses do not accurately approximate the $\alpha$-cut (unless $\alpha\approx 1$, see Figure 1).
+In essence, approaches based on regular ensemble training (wangCredalWrapper2024; nguyenCredalEnsembling2025) form an example of this strategy.
+However, an issue is that all hypotheses $h$ typically tend to cluster around $h^{ML}$, hence the obtained hypotheses do not accurately approximate the $\alpha$-cut (unless $\alpha\approx 1$, see Figure˜1).
 To obtain better coverage of $\mathcal{C}_{\alpha}$, we propose to train $M$ hypotheses using an early stopping strategy, namely until a specific relative likelihood value $\tau_{i}$ is reached.
@@
 This guarantees a broad range of hypotheses in terms of relative likelihoods.
-The influence of the ensemble size $M$ is analyzed in an ablation study in Section C.2.
+The influence of the ensemble size $M$ is analyzed in an ablation study in Section˜C.2.
 Although this should cover the $\alpha$-cut well, there might still be a lack of diversity in the predictions of the resulting hypotheses.
-In the literature, it is known that training ensembles without explicitly encouraging diversity actually minimizes diversity (Abe et al., 2022).
+In the literature, it is known that training ensembles without explicitly encouraging diversity actually minimizes diversity (abeBestDeep2022).
 To encourage diversity, we introduce a novel initialization strategy called ToBias.
@@
 As the predictors in the ensemble are trained, the amount of knowledge increases, and the predicted credal set should shrink.
-To enforce this in a finite predictor scenario, we ensure that the initial predictions of the learners correspond to degenerate probability distributions at vertices of the $(K-1)$-simplex, effectively covering the entire simplex in the initially predicted credal set.
-Figure 1 illustrates the mechanism of ToBias initialization and the overall learning process.
+To enforce this in a finite predictor scenario, we ensure that the initial predictions of the learners correspond to degenerate probability distributions at vertices of the (K1)(K-1)-simplex, effectively covering the entire simplex in the initially predicted credal set.
+Figure˜1 illustrates the mechanism of ToBias initialization and the overall learning process.
 Technically, the initialization is performed by assigning a large constant to one of the biases in the final layer of each predictor.
@@
 In the remainder of this work, we use $\beta=100$.
-An ablation study examining the influence of ToBias initialization is presented in Section C.2.
+An ablation study examining the influence of ToBias initialization is presented in Section˜C.2.
 Figure 3: Effect of ToBias.
@@
 b) ToBias initializes at the vertices resulting in diverse predictions.
-We illustrate the difference between regular initialization and ToBias initialization in Figure 3.
+We illustrate the difference between regular initialization and ToBias initialization in Figure˜3.
 Observe that without ToBias initialization, the credal set initially represents a state of low epistemic uncertainty, as all predictions are concentrated around the barycenter of the probability simplex, and increases its uncertainty as training progresses.
@@
 $\tilde{\mathcal{C}}_{\alpha}=\{h_{i}\ :\hat{\gamma}(h_{i})\approx\tau_{i}\}_{i=0}^{M-1}\subseteq\mathcal{H},$
-where $\hat{\gamma}(h_{i})\approx\tau_{i}$ is enforced by the training process described above. The full algorithm to train Credal Relative Likelihood (CreRLα) ensembles is presented in Algorithm 1.
-Credal Set Predictions. Given a query instance $\bm{x}\in\mathcal{X}$, the set $\tilde{\mathcal{C}}_{\alpha}$ induces a finite approximation of the credal set:
+where $\hat{\gamma}(h_{i})\approx\tau_{i}$ is enforced by the training process described above. The full algorithm to train Credal Relative Likelihood (CreRL) ensembles is presented in Algorithm˜1.
+Credal Set Predictions.
+Given a query instance $\bm{x}\in\mathcal{X}$, the set $\tilde{\mathcal{C}}_{\alpha}$ induces a finite approximation of the credal set:
 $\tilde{\mathcal{Q}}_{\bm{x},\alpha}=\{p(\cdot\,|\,\bm{x},h):h\in\tilde{\mathcal{C}}_{\alpha}\}\subseteq\Delta_{K}.$
@@
 Note that the convex hull is contained in the “box” credal set induced by the intervals.
-In the remainder of this work, we opt for the latter approach and define the approximation of $\mathcal{Q}_{\bm{x},\alpha}$ as follows (de Campos et al., 1994):
+In the remainder of this work, we opt for the latter approach and define the approximation of $\mathcal{Q}_{\bm{x},\alpha}$ as follows (camposProbabilityIntervals1994):
 $\hat{\mathcal{Q}}_{\bm{x},\alpha}=\{p(\cdot\,|\,\bm{x}):p(y_{k}\,|\,\bm{x})\in[\underline{p}(y_{k}\,|\,\bm{x}),\overline{p}(y_{k}\,|\,\bm{x})],\forall y_{k}\in\mathcal{Y}\}$
 5 Empirical Results
-To empirically evaluate our method, we measure its performance in terms of coverage and efficiency, as defined in Section 2, along with the corresponding Pareto criterion.
+To empirically evaluate our method, we measure its performance in terms of coverage and efficiency, as defined in Section˜2, along with the corresponding Pareto criterion.
 Additionally, we evaluate our approach on the downstream task of OoD detection, which is often used to assess the quality of (epistemic) uncertainty representations.
-We compare our approach (CreRLα) to suitable baselines by implementing the following approaches: Credal Wrapper (CreWra) (Wang et al., 2024b), Credal Ensembling (CreEnsα) (Nguyen et al., 2025), Credal Bayesian Deep Learning (CreBNN) (Caprio et al., 2023), and Credal Deep Ensembles (CreNet) (Wang et al., 2024a).
+We compare our approach (CreRL) to suitable baselines by implementing the following approaches: Credal Wrapper (CreWra) (wangCredalWrapper2024), Credal Ensembling (CreEns) (nguyenCredalEnsembling2025), Credal Bayesian Deep Learning (CreBNN) (caprioImpreciseBayesian2023), and Credal Deep Ensembles (CreNet) (wangCredalDeep2024).
 These methods represent the current state-of-art for credal prediction and to the best of our knowledge we are the first to do a systematic comparison of these approaches in a unified benchmark.
-Since the coverage and efficiency evaluation metrics are not well-defined for Bayesian methods, a direct comparison to such methods is not possible.
+Since coverage and efficiency are not well-defined for Bayesian methods, a direct comparison to such methods is not possible.
 The code for all implementations and experiments is published in a Github repository [^1] https://github.com/timoverse/credal-prediction-relative-likelihood.
-Further experimental details of our method and the implementation of the baselines are provided in Appendix B.
+Further experimental details of our method and the implementation of the baselines are provided in Appendix˜B.
 Figure 4: Pareto front between coverage and efficiency for ChaosNLI, CIFAR-10, and QualityMRI.
-CreRLα (Ours) and CreEnsα allow trading off efficiency and coverage by varying $\alpha$.
+CreRL (Ours) and CreEns allow trading off efficiency and coverage by varying $\alpha$.
 The other baselines do not allow direct adaptation of the credal set and hence result in a single point.
 5.1 Predictive Performance
-We use our method to train an ensemble of fully-connected neural networks on the embeddings of the ChaosNLI dataset (Nie et al., 2020).
+We use our method to train an ensemble of fully-connected neural networks on the embeddings of the ChaosNLI dataset (nieChaosNLI2020).
 This dataset features premise-hypothesis pairs in textual form with the goal to classify each instance as one of three classes: entailed, neutral, and contradicted.
 Each instance in ChaosNLI has 100 annotations, which can be seen as a ground-truth distribution.
-We train an ensemble of ResNet18 models (He et al., 2016) based on our method on the CIFAR-10 dataset (Krizhevsky et al., 2009).
+We train an ensemble of ResNet18 models (he2016deep) based on our method on the CIFAR-10 dataset (krizhevsky2009learning).
 This dataset consists of 10 classes of images, commonly used as a benchmark for image classification tasks.
-We use the human annotations provided by the CIFAR-10H dataset (Peterson et al., 2019) as ground-truth distributions.
-In addition, we train an ensemble of ResNet18 models on the QualityMRI dataset from Schmarje et al.
-(2022).
-This dataset comprises MRI images with multiple annotations from radiologists, each providing a quality assessment of the image (Obuchowicz et al., 2020).
-These multi-class annotations were merged into annotations for two classes (Schmarje et al., 2022).
-The uncertainty in these labels, as represented by the high average entropy reported in Appendix C, confirms the importance of uncertainty quantification in medical settings (Löhr et al., 2024).
+We use the human annotations provided by the CIFAR-10H dataset (petersonHumanUncertainty2019) as ground-truth distributions.
+In addition, we train an ensemble of ResNet18 models on the QualityMRI dataset from schmarjeIsOne2022.
+This dataset comprises MRI images with multiple annotations from radiologists, each providing a quality assessment of the image (qualityMRI).
+These multi-class annotations were merged into annotations for two classes (schmarjeIsOne2022).
+The uncertainty in these labels, as represented by the high average entropy reported in Appendix˜C, confirms the importance of uncertainty quantification in medical settings (lohr2024towards).
 To train the neural networks, the annotations are converted into a probability distribution, which we consider to be the ground-truth.
@@
 Thus, our method does not require ground-truth probability distributions, or soft labels, for training.
-We evaluate our method using the coverage and efficiency as defined in (2) and (3), where the expectation is approximated by averaging over the the instances in the test sets of the respective datasets.
-The results can be found in Figure 4.
-In Appendix C, we report the accuracy of the ensemble members for each implementation.
+We evaluate our method using the coverage and efficiency as defined in (2) and (3), where the expectation is approximated by averaging over the instances in the test set of the respective dataset.
+The results can be found in Figure˜4.
+Appendix˜C, we report the accuracy of the ensemble members for each implementation.
 Trade-off between Efficiency and Coverage.
 In general, both our method (CreRLα) and the CreEnsα approach are flexible and allow to trade off efficiency and coverage of the credal predictor by varying $\alpha$.
-However, CreEnsα can only attain credal sets in the low coverage region, whereas our approach can predict credal sets with a wide range of coverage values, including high coverage.
+However, CreEns can only attain credal sets in the low coverage region, whereas our approach can predict credal sets with a wide range of coverage values, including high coverage.
 The other methods only deliver fixed credal sets, meaning there is no direct way to to increase (or decrease) the coverage and efficiency.
@@
 The CreNet converges to a point prediction, because its loss function is constructed such that the lower probability converges to the upper probability, and hence is generally in the top left region of the Pareto figure.
-We refer to Appendix A for more details about these methods.
+We refer to Appendix˜A for more details about these methods.
 For the ChaosNLI dataset, our approach dominates the CreBNN approach and the CreWra resembles a point on the Pareto front obtained by our method.
@@
 The CreWra baseline is close to our Pareto front.
-Since all ensemble members have high performance, as illustrated by their accuracies in Appendix C, and there is no explicit diversity promotion in CreEnsα, it produces small sets for every $\alpha$.
+Since all ensemble members have high performance, as illustrated by their accuracies in Appendix˜C, and there is no explicit diversity promotion in CreEnsα, it produces small sets for every $\alpha$.
 CreNet, again, produces a (close to) point prediction with low coverage.
-We refer to Appendix B for more details.
+We refer to Appendix˜B for more details.
 For QualityMRI, the story is largely consistent with that of the ChaosNLI dataset. The difference is that our method is only positioned in the high coverage half of the Pareto figure.
@@
 In addition, one may argue that the low coverage region represents credal sets that are of limited benefit for practitioners, as they are unlikely to demand a coverage of, say, 0.3.
-Furthermore, Figure 4 shows that the shape and position of the Pareto front varies across datasets suggesting that the coverage-efficiency trade-off is inherently data-dependent.
+Furthermore, Figure˜4 shows that the shape and position of the Pareto front varies across datasets suggesting that the coverage-efficiency trade-off is inherently data-dependent.
 In particular, the better the maximum likelihood model, the more the Pareto front should move towards the top right (high coverage, high efficiency), which we also observe in the case of CIFAR-10.
-To further study the behavior of our method, we perform ablation experiments, varying the number of ensemble members and the ToBias initialization constant. The results are presented in Appendix C.
-Table 1: Out-of-Distribution detection based on epistemic uncertainty using CIFAR-10 as in-Distribution dataset. Mean and standard deviation over 3 runs are reported. Best performance is bold.
+To further study the behavior of our method, we present ablation experiments in Appendix˜C, varying the number of members in the ensemble and the ToBias initialization constant.
+Table 1: Out-of-Distribution detection based on epistemic uncertainty.
+CIFAR-10 is used as the in-Distribution dataset.
+The mean and standard deviation over 3 runs are reported.
+Best performance is in bold.
 Method SVHN Places CIFAR-100 FMNIST ImageNet
@@
 $\text{CreEns}_{0.0}$ $0.955\scriptstyle{\pm 0.001}$ $0.913\scriptstyle{\pm 0.000}$ $0.914\scriptstyle{\pm 0.001}$ $0.949\scriptstyle{\pm 0.001}$ $0.888\scriptstyle{\pm 0.000}$
-CreBNN $0.907\scriptstyle{\pm 0.006}$ $0.885\scriptstyle{\pm 0.002}$ $0.880\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.002}$ $0.859\scriptstyle{\pm 0.002}$
+CreBNN $0.907\scriptstyle{\pm 0.006}$ $85\scriptstyle{\pm 0.002} 0.880$$80\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.002}$ $0.859\scriptstyle{\pm 0.002}$
 CreNet $0.943\scriptstyle{\pm 0.003}$ $\mathbf{0.918\scriptstyle{\pm 0.000}}$ $0.912\scriptstyle{\pm 0.000}$ $0.951\scriptstyle{\pm 0.002}$ $0.884\scriptstyle{\pm 0.001}$
@@
 In the literature, this is commonly evaluated by computing epistemic uncertainty and using it to distinguish between iD and OoD instances.
-We adopt the epistemic uncertainty measure based on the additive decomposition proposed (and axiomatically justified) by Abellán et al.
-(2006):
+We adopt the epistemic uncertainty measure based on the additive decomposition proposed (and axiomatically justified) by abellanDisaggregatedTotal2006:
 $\operatorname{EU}(\hat{\mathcal{Q}}_{\bm{x},\alpha})=\overline{S}(\hat{\mathcal{Q}}_{\bm{x},\alpha})-\underline{S}(\hat{\mathcal{Q}}_{\bm{x},\alpha}),$ (5)
 where $\overline{S}(\hat{\mathcal{Q}}_{\bm{x},\alpha})$ denotes the maximum Shannon entropy of all distributions in $\hat{\mathcal{Q}}_{\bm{x},\alpha}$ and, likewise, $\underline{S}(\hat{\mathcal{Q}}_{\bm{x},\alpha})$ the minimum entropy.
-[^2] The bounds of Shannon entropy are computed numerically using SciPy (Virtanen et al., 2020).
+[^2] The bounds of Shannon entropy are computed numerically using SciPy (virtanenSciPy2020).
 The performance is then measured using the AUROC.
-Section B.4 provides details about the implementation of the optimization.
-We utilize the same CIFAR-10 ensembles of ResNet18 models and compare to the same baselines as in Section 5.1.
-We introduce SVHN Netzer et al.
-(2011), Places365 Zhou et al.
-(2018), CIFAR-100 Krizhevsky et al.
-(2009), FashionMNIST Xiao et al.
-(2017), and ImageNet Deng et al.
-(2009) as out-of-distribution datasets, and compute epistemic uncertainty as in (5).
+Section˜B.4 provides details about the implementation of the optimization.
+We utilize the same CIFAR-10 ensembles of ResNet18 models and compare to the same baselines as in Section˜5.1.
+We introduce SVHN netzerReading2011, Places365 zhouPlaces2018, CIFAR-100 krizhevsky2009learning, FashionMNIST xiaoFashionMNIST2017, and ImageNet deng2009imagenet as out-of-distribution datasets, and compute epistemic uncertainty as in (5).
 Our method ($\text{CreRL}_{\alpha}$) and the Credal Ensembling ($\text{CreEns}_{\alpha}$) approach make use of a hyper-parameter $\alpha$.
 The results with the best performing $\alpha$ are presented here.
-The results, summarized in Table 1, present the OoD detection performance.
-Additionally, we provide results for both methods with different $\alpha$ values in Appendix C.
+The results, summarized in Table˜1, present the OoD detection performance.
+We provide further results for both methods with different $\alpha$ values in Appendix˜C.
 Our method (CreRLα) with $\alpha=1$ performs either the best or is on par with the best methods in OoD detection, except for the SVHN dataset, where the CreWra approach has a small advantage.
@@
 Given that $\alpha$ strongly influences epistemic uncertainty for iD instances, we expect a larger $\alpha$ to increase the discrepancy in uncertainty between iD and OoD instances.
-As shown in Table 1, across all OoD datasets, a larger $\alpha$ indeed leads to higher AUROC scores, confirming that larger $\alpha$ values lead to better separation of iD and OoD instances.
+As shown in Table˜1, across all OoD datasets, a larger $\alpha$ indeed leads to higher AUROC scores, confirming that larger $\alpha$ values lead to better separation of iD and OoD instances.
 6 Conclusion
@@
 Furthermore, the parameter $\alpha$, which specifies the likelihood ratio between the maximum likelihood model and the “weakest” model that is still included in the $\alpha$-cut, can be adjusted depending on the task at hand and used to control the trade-off between the coverage and efficiency of credal predictions.
-Experimentally, we have shown that this workflow performs well in terms of providing coverage of the ground-truth conditional distributions, while maintaining a competitive efficiency, as well as performing competitively in downstream tasks such as OoD detection.
+Experimentally, we have shown that this workflow provides strong coverage of the ground-truth conditional distributions, while maintaining a competitive efficiency, as well as performing competitively in downstream tasks such as OoD detection.
 The relation between $\alpha$ and the performance on different tasks sheds further light on the trade-off between diversity and prediction performance.
@@
 Broader Impacts.
-This work contributes to the development of more reliable machine learning models by improving uncertainty quantification.
-We do not foresee any direct negative broader impacts arising from this.
+This work contributes to the development of reliable machine learning models by improving uncertainty quantification. We do not foresee any direct negative broader impacts arising from this.
+Acknowledgments and Disclosure of Funding
+We gratefully thank the anonymous reviewers for their valuable feedback, which helped improve this work.
+This work has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101073307.
+Felix Mohr participated through the project ING-312-2023 at Universidad de La Sabana.
+We acknowledge the Munich Center for Machine Learning (MCML) for their support of this work.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Our claims are supported by Sections˜3, 4 and 5.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The limitations are discussed in Section˜6.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: We do not provide theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: The full experimental details are discussed in Sections˜5 and B.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Code is provided in Section˜5 with run instructions in the respective repository. The data used is publicly available.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The full details are described in Appendix˜B.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We provide information about the error bars for every figure and table, see Sections˜5, C.2 and C.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We provide information on the compute resources in Appendix˜B.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We have read the NeurIPS Code of Ethics and confirm that our work complies with it.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We provide a broader impacts statements in Section˜6.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: We do not contribute any such models or datasets.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We provide references to the models used and additionally include licenses for the datasets in Appendix˜B.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: We do not introduce any new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The method proposed in this work does involve LLMs.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix A Extended Related Work
 Uncertainty Representation.
-In machine learning, uncertainty is often represented by Bayesian methods [Mackay, 1992, Blundell et al., 2015].
-Frequently, the Bayesian posterior is approximated by ensembles [Lakshminarayanan et al., 2017] or methods such as Dropout [Gal and Ghahramani, 2016] or Laplace approximation [Daxberger et al., 2021].
-An important characteristic of such representations, especially for uncertainty tasks, is diversity [D’Angelo and Fortuin, 2021, Wood et al., 2023].
-Some works enforce this by means of regularization [de Mathelin et al., 2023], whereas others vary hyperparameters across models within the ensemble [Wenzel et al., 2020] or enforce diversity in the representations [Lopes et al., 2022].
-Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty [Zaffalon, 2001, Corani and Zaffalon, 2008, Corani and Mignatti, 2015].
-Antonucci et al.
-[2012a] proposed to generate such sets based on relative likelihoods.
-Relative likelihoods, also referred to as normalized likelihoods, have also been used in machine learning with simple model classes such as logistic regression [Senge et al., 2014, Cella and Martin, 2024].
+In machine learning, uncertainty is often represented by Bayesian methods [mackayBayesianMethods1992, blundellWeightUncertainty2015].
+Frequently, the Bayesian posterior is approximated by ensembles [lakshminarayananDeepEnsembles2017] or methods such as Dropout [galDropoutAs2016] or Laplace approximation [daxbergerLaplaceRedux2021].
+An important characteristic of such representations, especially for uncertainty tasks, is diversity [d2021repulsive, woodUnifiedTheory2023].
+Some works enforce this by means of regularization [deMathelinDeepAnti2023], whereas others vary hyperparameters across models within the ensemble [wenzelHyperparameterEnsembles2020] or enforce diversity in the representations [lopesNoOne2022].
+Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty [zaffalonStatisticalInference2001, coraniLearningReliable2008, coraniCredalModel2015].
+antonucciLikelihoodBased2012 proposed to generate such sets based on relative likelihoods.
+Relative likelihoods, also referred to as normalized likelihoods, have also been used in machine learning with simple model classes such as logistic regression [sengeReliableClassification2014, cellaVariationalApproximations2024].
 In this work, we build upon these approaches and address the challenges that emerge when adapting the relative likelihood to a setting with complex predictors.
 Recently, credal sets have been applied in the context of machine learning.
-Wang et al.
-[2024b] take multiple samples from a Bayesian posterior or ensembles and derive class-wise lower and upper probabilities from this.
-Based on these samples or ensemble predictions, a credal set is constructed by including all probability distributions such that the individual class probabilities are in the respective lower to upper class probability interval.
-Nguyen et al.
-[2025] also use ensemble predictions, but they additionally allow for the discarding of potential outliers to ensure that the resulting credal set too large.
+wangCredalWrapper2024 take multiple samples from a Bayesian posterior or ensembles and derive class-wise lower and upper probabilities.
+Based on these samples, a credal set is constructed by including all probability distributions such that the individual predicted class probabilities are in the respective lower to upper class probability interval.
+nguyenCredalEnsembling2025 also construct credal sets from ensemble predictions, but with the additional option of discarding potential outliers to prevent the set from becoming too large.
 This is done by comparing all ensemble predictions to a representative prediction, e.g.
@@
 The credal set is then constructed by taking the convex hull of the remaining probability distributions.
-Other methods directly train neural networks to predict intervals by explicitly having the neural network predict a lower and upper probability for every class [Wang et al., 2024a].
+Other methods directly train neural networks to predict intervals by explicitly predicting a lower and upper probability for every class [wangCredalDeep2024].
 In combination with a custom loss function, consisting of regular cross-entropy for the upper probabilities and cross-entropy computed on the highest loss subset of a batch for the lower probabilities, the claim is that this approach encourages both “optimistic” and “pessimistic” predictions.
-Credal sets are constructed by taking the same interval-based approach as [Wang et al., 2024b].
+Credal sets are constructed by taking the same interval-based approach as [wangCredalWrapper2024].
 Besides this, hybrid methods, combining multiple uncertainty frameworks, have also been proposed.
-Caprio et al.
-[2023] combine Bayesian deep learning and credal sets by considering sets of priors over weights of neural networks.
+caprioImpreciseBayesian2023 combine Bayesian deep learning and credal sets by considering sets of priors over weights of neural networks.
 Training these neural network by variational inference then results in a set of posteriors.
 Based on this, a credal set is constructed by sampling from the Bayesian neural networks and taking the convex hull of the sampled probability distributions.
-Another approach leverages conformal prediction to construct credal sets with validity guarantees [Javanmardi et al., 2024].
+Another approach leverages conformal prediction to construct credal sets with validity guarantees [javanmardiConformalizedCredal2024].
 However, this work uses distributions over classes to perform the conformal calibration step.
-Since our method does not require such data, we do not use this method as a baseline.
-Our work distinguishes itself from aforementioned works by making use of relative likelihood cuts which allow for an intuitive and adaptable construction of the credal set.
+Since our method does not require such data, we exclude this method as a baseline.
+Our work distinguishes itself from aforementioned works by using relative likelihood cuts which allow for an intuitive and adaptable construction of the credal set.
+Moreover, our approach is conceptually related to Rashomon sets [semenova2022existence].
+Both characterize a collection of models whose performance exceeds a given threshold, thereby facing the same challenge in approximating this set of plausible models [donnelly2025rashomon].
+However, the objectives differ: Rashomon sets are primarily concerned with interpretability and (syntactic) model diversity, whereas our focus is on uncertainty quantification and predictive diversity.
 Uncertainty Quantification.
 Given a credal representation of uncertainty, there are many ways to quantify uncertainty.
-Some measures consider only the epistemic uncertainty [Abellán and Moral, 2000], whereas others use entropy to reason about the total uncertainty [Abellán and Moral, 2003].
-In this line of work, Abellán et al.
-[2006] also proposed measures based on entropy that decompose the total uncertainty into an aleatoric and epistemic component.
-Conversely, Antonucci et al.
-[2012b] quantify uncertainty by measuring the lack of dominance of one class over others in the predictive distributions represented by the credal set.
-Hüllermeier et al.
-[2022] offer a critical analysis of these measures and propose an alternative for the dominance-based measure.
-Recently, Hofman et al.
-[2024] proposed to quantify credal uncertainty based on a decomposition of scoring rules.
-In the following, we consider the uncertainty measures by Abellán et al.
-[2006] in order to ensure a fair comparison with previous works.
+Some measures consider only the epistemic uncertainty [abellanNonSpecificty2000], whereas others use entropy to reason about the total uncertainty [abellanMaximumOf2003].
+In this line of work, abellanDisaggregatedTotal2006 also proposed measures based on entropy that decompose the total uncertainty into an aleatoric and epistemic component.
+Conversely, antonucciActiveLearning2012 quantify uncertainty by measuring the lack of dominance of one class over others in the predictive distributions represented by the credal set.
+huellermeierQuantificationOf2022 offer a critical analysis of these measures and propose an alternative for the dominance-based measure.
+Recently, hofmanQuantifyingAleatoric2024 proposed to quantify credal uncertainty based on a decomposition of scoring rules.
+In the following, we consider the uncertainty measures by abellanDisaggregatedTotal2006 in order to ensure a fair comparison with previous works.
 Appendix B Experimental Details
 B.1 Datasets
-ChaosNLI The ChaosNLI dataset, introduced by Nie et al.
-[2020], is a large-scale dataset designed to study human disagreement in natural language inference (NLI) tasks.
-It comprises 100 human annotations per example for 3,113 examples from the SNLI and MNLI datasets, and 1,532 examples from the $\alpha$NLI dataset, totaling approximately 464,500 annotations.
-In line with Javanmardi et al.
-[2024], we use only the SNLI and MNLI subsets of the ChaosNLI dataset, but for simplicity, we will refer to this as the ChaosNLI dataset.
-Each example includes metadata such as the unique identifier, counts of each label assigned by annotators, the majority label, label distribution, entropy of the label distribution, the original example text, and the original label from the source dataset.
-This dataset enables a detailed analysis of the distribution of human opinions in NLI tasks, highlighting instances of high disagreement and questioning the validity of using majority labels as the sole ground truth.
-ChaosNLI is publicly available under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license.
-We train our models on the 768-dimensional embeddings of the ChaosNLI dataset retrieved from https://github.com/alireza-javanmardi/conformal-credal-sets.
-We refer to [Javanmardi et al., 2024] for more details on the generation of the embeddings.
-CIFAR-10 The CIFAR-10 dataset is a widely used benchmark in machine learning and computer vision, introduced by Krizhevsky et al.
-[2009], and Geoffrey Hinton in 2009.
+ChaosNLI
+[…]
+CIFAR-10
+The CIFAR-10 dataset is a widely used benchmark in machine learning and computer vision, introduced by krizhevsky2009learning, and Geoffrey Hinton in 2009.
 It comprises 60,000 color images at a resolution of 32×32 pixels, evenly distributed across 10 distinct classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
@@
 While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
-CIFAR-10H The CIFAR-10H dataset provides human-derived soft labels for the 10,000 images in the CIFAR-10 test set, capturing the variability in human annotation during image classification tasks.
-Developed by Peterson et al.
-[2019], the dataset comprises 511,400 annotations collected from 2,571 Amazon Mechanical Turk workers, with each image receiving approximately 51 labels.
+CIFAR-10H
+The CIFAR-10H dataset provides human-derived soft labels for the 10,000 images in the CIFAR-10 test set, capturing the variability in human annotation during image classification tasks.
+Developed by petersonHumanUncertainty2019, the dataset comprises 511,400 annotations collected from 2,571 Amazon Mechanical Turk workers, with each image receiving approximately 51 labels.
 Annotators classified images into one of the ten CIFAR-10 categories, enabling the construction of probability distributions over labels for each image.
 CIFAR-10H is publicly available under the Creative Commons BY-NC-SA 4.0 license.
-CIFAR-100 The CIFAR-100 dataset, introduced by Krizhevsky et al.
-[2009], comprises 60,000 color images at 32×32 resolution, divided into 100 classes with 600 images each.
+CIFAR-100
+The CIFAR-100 dataset, introduced by krizhevsky2009learning, comprises 60,000 color images at 32×32 resolution, divided into 100 classes with 600 images each.
 Each image has a “fine” label (specific class) and a “coarse” label (superclass), with the 100 classes grouped into 20 superclasses.
@@
 While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
-QualityMRI The QualityMRI dataset, introduced by Obuchowicz et al.
-[2020], is part of the Data-Centric Image Classification (DCIC) Benchmark, which aims to evaluate the impact of dataset curation on model performance.
+QualityMRI
+The QualityMRI dataset, introduced by qualityMRI, is part of the Data-Centric Image Classification (DCIC) Benchmark, which aims to evaluate the impact of dataset curation on model performance.
 The dataset contains 310 magnetic resonance (MRI) images spanning various quality levels and is designed to assess the MRI image quality.
 The dataset is publicly available under the Creative Commons BY-SA 4.0 license.
-SVHN The SVHN dataset, introduced by Netzer et al.
-[2011], consists of over 600,000 32×32 RGB images of digits (0–9) obtained from real-world house number images in Google Street View.
+SVHN
+The SVHN dataset, introduced by netzerReading2011, consists of over 600,000 32×32 RGB images of digits (0–9) obtained from real-world house number images in Google Street View.
 It includes three subsets: 73,257 training images, 26,032 test images, and 531,131 additional images for extra training.
@@
 While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
-Places365 Places365, introduced by Zhou et al.
-[2018], is a large-scale scene recognition dataset containing 1.8 million training images across 365 scene categories.
+Places365
+Places365, introduced by zhouPlaces2018, is a large-scale scene recognition dataset containing 1.8 million training images across 365 scene categories.
 The validation set includes 50 images per category, and the test set has 900 images per category.
@@
 While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
-FMNIST Fashion-MNIST (FMNIST), introduced by Xiao et al.
-[2017], is a dataset of Zalando’s article images, comprising 70,000 28×28 grayscale images labeled across 10 classes, such as T-shirt/top, Trouser, and Sneaker.
+FMNIST
+Fashion-MNIST (FMNIST), introduced by xiaoFashionMNIST2017, is a dataset of Zalando’s article images, comprising 70,000 28×28 grayscale images labeled across 10 classes, such as T-shirt/top, Trouser, and Sneaker.
 It includes 60,000 training and 10,000 test images and serves as a direct replacement for the original MNIST dataset for benchmarking machine learning algorithms.
 FMNIST is publicly available under the MIT License.
-ImageNet ImageNet, introduced by Deng et al.
-[2009], is a large-scale image database organized according to the WordNet hierarchy, containing over 14 million images across more than 20,000 categories.
+ImageNet
+ImageNet, introduced by deng2009imagenet, is a large-scale image database organized according to the WordNet hierarchy, containing over 14 million images across more than 20,000 categories.
 The ILSVRC subset (ImageNet-1K) includes 1,281,167 training images, 50,000 validation images, and 100,000 test images across 1,000 classes.
@@
 B.2 Models
-Fully-Connected Network We train fully-connected neural networks on the ChaosNLI dataset.
-The network consists of 4 linear layers with $[768-256-64-16-3]$ units with ReLU activations, except for the last layer, which has Softmax transformation to transform the logits into probabilities.
+Fully-Connected Network
+We train fully-connected neural networks on the ChaosNLI dataset.
+The network consists of 4 linear layers with [76825664163][768-256-64-16-3] units with ReLU activations, except for the last layer, which has Softmax transformation to transform the logits into probabilities.
 We use the hyperparameters (cf.
-Table 2) similar to the optimal parameters found in Javanmardi et al.
-[2024].
-ResNet18 For experiments on the CIFAR-10 dataset, we use the PyTorch ResNet-18 implementation and hyperparameters provided by https://github.com/kuangliu/pytorch-cifar.
+Table˜2) similar to the optimal parameters found in javanmardiConformalizedCredal2024.
+ResNet18
+For experiments on the CIFAR-10 dataset, we use the PyTorch ResNet-18 implementation and hyperparameters provided by https://github.com/kuangliu/pytorch-cifar.
 This model is specifically optimized for CIFAR-10 and is trained from scratch, without any pretraining on ImageNet.
@@
 no pretrained weights.
-Hyperparameters Each dataset is trained using a dedicated set of hyperparameters as presented in Table 2.
+Hyperparameters
+Each dataset is trained using a dedicated set of hyperparameters as presented in Table˜2.
 We evaluated multiple configurations and selected the best-performing ones for each dataset.
 To ensure fair and consistent comparisons, all models trained on a given dataset, both our approach and the baselines, use the same hyperparameter settings.
-The only exception is the CreBNN, which requires a KL-divergence penalty of $71e-$ 7 and zero weight decay when using the Adam optimizer [Kingma and Ba, 2015].
-When we apply the SGD optimizer with a learning rate scheduler, namely Cosine Annealing [Loshchilov and Hutter, 2017], CreBNN requires additionally a momentum of $0.9$ to enable effective learning.
+The only exception is the CreBNN, which requires a KL-divergence penalty of 1e71e-7 and zero weight decay when using the Adam optimizer [kingmaAdam2015].
+When we apply the SGD optimizer with a learning rate scheduler, namely Cosine Annealing [loshchilovSGDR2017], CreBNN requires additionally a momentum of $0.9$ to enable effective learning.
 Table 2: Hyperparameters used for each dataset.
@@
 B.3 Out-of-Distribution Detection
-We use the SVHN [Netzer et al., 2011], Places365 [Zhou et al., 2018], CIFAR-100 [Krizhevsky et al., 2009], Fashion-MNIST [Xiao et al., 2017], and ImageNet [Le and Yang, 2015] datasets for Out-of-Distribution detection.
+We use the SVHN [netzerReading2011], Places365 [zhouPlaces2018], CIFAR-100 [krizhevsky2009learning], Fashion-MNIST [xiaoFashionMNIST2017], and ImageNet [leTinyImagenet2015] datasets for Out-of-Distribution detection.
 The Out-of-Distribution task is treated as a binary classification task where the epistemic uncertainty is used as the classification criterion.
@@
 B.4 Computing Uncertainty
-Computing the total and aleatoric uncertainty involves solving a per instance optimization problem, after which the epistemic uncertainty is obtained as their difference (see Equation 5).
+Computing the total and aleatoric uncertainty involves solving a per instance optimization problem, after which the epistemic uncertainty is obtained as their difference (see Equation˜5).
 This optimization is performed using SciPy’s minimize function with the SLSQP solver and the default parameters.
-Interval-Based For the interval-based credal sets, the optimization is initialized with the mean of the predicted distributions, bounded between the lower and upper probabilities for each class, and constrained to ensure that the solution forms a valid probability distribution, namely, the class probabilities must sum to $1$.
-Convex Hull For the credal sets based on the convex hull, the optimization is done on the weights of convex combination, instead of the probability distribution.
-Uniform weights are used as the initial value, the weights are bounded between 00 and $1$, and constrained to sum to $1$.
-Estimated Computing Time Here, we provide an estimated upper bound of the computation time for computing the lower and upper entropy of a credal set based on the example of the largest possible credal set, namely the full probability simplex for a $10$-class problem, such as CIFAR-10.
+Interval-Based
+For the interval-based credal sets, the optimization is initialized with the mean of the predicted distributions, bounded between the lower and upper probabilities for each class, and constrained to ensure that the solution forms a valid probability distribution, namely, the class probabilities must sum to $1$.
+Convex Hull
+For the credal sets based on the convex hull, the optimization is done on the weights of convex combination, instead of the probability distribution.
+Uniform weights are used as the initial value, the weights are bounded between 0 and $1$, and constrained to sum to $1$.
+Estimated Computing Time
+Here, we provide an estimated upper bound of the computation time for computing the lower and upper entropy of a credal set based on the example of the largest possible credal set, namely the full probability simplex for a $10$-class problem, such as CIFAR-10.
 For this largest set, computing the upper entropy using an interval-based credal set takes on average $0.03$ seconds, while computing the lower entropy takes about $0.02$ seconds.
@@
 We list all details regarding the implementations of the baselines that were used in the paper. In general, all baselines were implemented in our own code base.
-Credal Wrapper (CreWra) The Credal Wrapper was initially implemented in TensorFlow, but we reimplemented it in PyTorch to ensure compatibility with our framework.
+Credal Wrapper (CreWra)
+The Credal Wrapper was initially implemented in TensorFlow, but we reimplemented it in PyTorch to ensure compatibility with our framework.
 It follows a standard ensemble learning approach, training multiple models independently.
 Like our method, the Credal Wrapper constructs credal sets using class-wise upper and lower probability bounds, making it well-aligned with our implementation.
-Overall, we closely follow both the original paper and their available implementation [Wang et al., 2024b].
-Credal Ensembling ($\text{CreEns}_{\alpha}$) Since no official code was available for neural network implementations of Credal Ensembling, we reimplemented the method ourselves.
-Our implementation closely follows all details provided in Nguyen et al.
-[2025].
+Overall, we closely follow both the original paper and their available implementation [wangCredalWrapper2024].
+Credal Ensembling ($\text{CreEns}_{\alpha}$)
+Since no official code was available for neural network implementations of Credal Ensembling, we reimplemented the method ourselves.
+Our implementation closely follows all details provided in nguyenCredalEnsembling2025.
 The approach builds on standard ensemble training procedures, with inference adapted according to their proposed method of sorting predictions based on a distance measure and selecting only $\alpha\%$ closest predictions to construct credal sets.
 In our experiments, we use the Euclidean distance measure and evaluate several values of $\alpha$.
-Credal Deep Ensembles (CreNet) As the official implementation of Credal Deep Ensembles is only available in TensorFlow, we reimplemented the method in PyTorch to ensure compatibility with our codebase.
+Credal Deep Ensembles (CreNet)
+As the official implementation of Credal Deep Ensembles is only available in TensorFlow, we reimplemented the method in PyTorch to ensure compatibility with our codebase.
 Our implementation closely mirrors the original TensorFlow code, particularly in adapting the model architecture and loss function.
 Specifically, we replace each model’s final linear layer with a head comprising a linear layer outputting $2\times\text{classes}$ (representing upper and lower probability bounds), followed by a batch normalization layer and the custom IntSoftmax layer.
-We also reimplemented the proposed loss function, which computes a cross-entropy loss for the upper bounds and selectively backpropagates the lower-bound loss only for the $\delta\%$ of samples with the highest loss values, as described in Wang et al.
-[2024a].
-In our experiments, we use $\delta=0.5$, as suggested in Wang et al.
-[2024a].
-Credal Bayesian Deep Learning (CreBNN) No code or implementation details for Credal Bayesian Deep Learning (CreBNN) were made publicly available, and despite multiple attempts to contact the authors, we received no response or further clarification.
+We also reimplemented the proposed loss function, which computes a cross-entropy loss for the upper bounds and selectively backpropagates the lower-bound loss only for the $\delta\%$ of samples with the highest loss values, as described in wangCredalDeep2024.
+In our experiments, we use $\delta=0.5$, as suggested in wangCredalDeep2024.
+Credal Bayesian Deep Learning (CreBNN)
+No code or implementation details for Credal Bayesian Deep Learning (CreBNN) were made publicly available, and despite multiple attempts to contact the authors, we received no response or further clarification.
 As a result, we reimplemented the method ourselves based solely on the high-level description provided in the paper.
-In our implementation, each ensemble member is a Bayesian neural network (BNN) trained with variational inference using different priors, with prior means $\mu$ sampled from $[-1,1]$ and standard deviations $\sigma$ from $[0.1,2]$ to form a diverse prior set.
+In our implementation, each ensemble member is a Bayesian neural network (BNN) trained with variational inference using different priors, with prior means $\mu$ sampled from [1,1][-1,1] and standard deviations $\sigma$ from $[0.1,2]$ to form a diverse prior set.
 During inference, we draw one sample from each BNN to obtain a finite set of probability distributions, and construct the credal set as the convex hull of these predictions.
 B.7 Computing Resources
-To run the experiments presented in this work, we utilized the computing resources detailed in Table 3. The total estimated GPU usage amounts to approximately $750$ hours.
-Table 3: Specifications of Computing Resources
+To run the experiments presented in this work, we utilized the computing resources detailed in Table˜3. The total estimated GPU usage amounts to approximately $750$ hours.
+Table 3: Specifications of Computing Resources.
 Component Specification
@@
 C.1 Out-of-Distribution detection
-In addition to the OoD experiments in Section 5, we evaluate a broad range of values $\alpha$ for the CreEnsα approach.
+In addition to the OoD experiments in Section˜5, we evaluate a broad range of values $\alpha$ for the CreEns approach.
 Table 4: Out-of-Distribution detection based on epistemic uncertainty.
@@
 $\text{CreEns}_{0.4}$ $0.947\scriptstyle{\pm 0.001}$ $0.906\scriptstyle{\pm 0.001}$ $0.908\scriptstyle{\pm 0.000}$ $0.940\scriptstyle{\pm 0.001}$ $0.883\scriptstyle{\pm 0.001}$
-$\text{CreEns}_{0.2}$ $0.950\scriptstyle{\pm 0.001}$ $0.909\scriptstyle{\pm 0.001}$ $0.911\scriptstyle{\pm 0.000}$ $0.946\scriptstyle{\pm 0.001}$ $0.885\scriptstyle{\pm 0.001}$
+$\text{CreEns}_{0.2}$ $50\scriptstyle{\pm 0.001} 0.909$$09\scriptstyle{\pm 0.001}$ $0.911\scriptstyle{\pm 0.000}$ $0.946\scriptstyle{\pm 0.001}$ $0.885\scriptstyle{\pm 0.001}$
 $\text{CreEns}_{0.0}$ $0.955\scriptstyle{\pm 0.001}$ $0.913\scriptstyle{\pm 0.000}$ $0.914\scriptstyle{\pm 0.001}$ $0.949\scriptstyle{\pm 0.001}$ $0.888\scriptstyle{\pm 0.000}$
 CreNet $0.943\scriptstyle{\pm 0.003}$ $0.918\scriptstyle{\pm 0.000}$ $0.912\scriptstyle{\pm 0.000}$ $0.951\scriptstyle{\pm 0.002}$ $0.884\scriptstyle{\pm 0.001}$
-CreBNN $0.907\scriptstyle{\pm 0.006}$ $0.885\scriptstyle{\pm 0.002}$ $0.880\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.002}$ $0.859\scriptstyle{\pm 0.002}$
+CreBNN $0.907\scriptstyle{\pm 0.006}$ $85\scriptstyle{\pm 0.002} 0.880$$80\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.002}$ $0.859\scriptstyle{\pm 0.002}$
 $\text{CreRL}_{1.0}$ $0.948\scriptstyle{\pm 0.003}$ $\mathbf{0.918\scriptstyle{\pm 0.002}}$ $\mathbf{0.916\scriptstyle{\pm 0.001}}$ $\mathbf{0.957\scriptstyle{\pm 0.002}}$ $\mathbf{0.889\scriptstyle{\pm 0.002}}$
@@
 To do so, we vary the constant, taking values $\beta\in\{5,10,20,30,50,80,100,200,500\}$.
-The ensemble is then trained as proposed in Algorithm 1.
-The coverage and efficiency Pareto front is shown in Figure 5 with the mean and standard deviation over three runs.
+The ensemble is then trained as proposed in Algorithm˜1.
+The coverage and efficiency Pareto front is shown in Figure˜5 with the mean and standard deviation over three runs.
 We split the results into three Figures for better readability.
@@
 To do so, we vary the constant, taking values $M\in\{1,2,3,4,5,10,20,30,50\}$.
-The ensemble is then trained as proposed in Algorithm 1.
-The coverage and efficiency Pareto front is shown in Figure 6 with the mean and standard deviation over three runs.
+The ensemble is then trained as proposed in Algorithm˜1.
+The coverage and efficiency Pareto front is shown in Figure˜6 with the mean and standard deviation over three runs.
 We split the results into three Figures for better readability.
@@
 After this, the number of ensemble members does not have a significant impact on the resulting Pareto front anymore.
-Table 5: Average entropy of ground-truth distribution in the test sets of ChaosNLI, CIFAR-10 and QualityMRI.
+Table 5: Average entropy of ground-truth distributions in the test sets of ChaosNLI, CIFAR-10 and QualityMRI.
 Dataset Avg. Entropy
@@
 This is reflected, for example, in the degree of disagreement among the collected annotations, which serve as the ground-truth distributions in our evaluations.
-To provide further insight, we report the average entropy of these ground-truth distributions in Table 5.
+To provide further insight, we report the average entropy of these ground-truth distributions in Table˜5.
 C.4 Performance of Ensemble Members
-We visualized the trade-off between coverage and efficiency in Figure 4, and provide the corresponding numerical values in Table 6.
+We visualized the trade-off between coverage and efficiency in Figure˜4, and provide the corresponding numerical values in Table˜6.
 We also report the accuracies of the individual predictors within each ensemble for both the baselines and our proposed method to provide insight into their standalone performance (cf.
-Tables 7 and 8).
-For CreNet, whose ensemble members directly predict probability intervals, we use their intersection probability technique to derive pointwise predictions [Wang et al., 2024a].
+Tables˜7 and 8).
+For CreNet, whose ensemble members directly predict probability intervals, we use their intersection probability technique to derive pointwise predictions [wangCredalDeep2024].
 Table 6: Coverage and efficiency for different methods across datasets.
@@
 CreWra $0.453\scriptstyle{\pm 0.050}$ $0.607\scriptstyle{\pm 0.037}$ $0.450\scriptstyle{\pm 0.001}$ $0.963\scriptstyle{\pm 0.000}$ $0.355\scriptstyle{\pm 0.057}$ $0.608\scriptstyle{\pm 0.021}$
-CreNet $0.001\scriptstyle{\pm 0.002}$ $0.978\scriptstyle{\pm 0.008}$ $0.094\scriptstyle{\pm 0.010}$ $0.999\scriptstyle{\pm 0.000}$ $0.188\scriptstyle{\pm 0.020}$ $0.792\scriptstyle{\pm 0.003}$
-CreBNN $0.195\scriptstyle{\pm 0.019}$ $0.648\scriptstyle{\pm 0.019}$ $0.331\scriptstyle{\pm 0.004}$ $0.871\scriptstyle{\pm 0.003}$ $0.780\scriptstyle{\pm 0.312}$ $0.184\scriptstyle{\pm 0.255}$
-CreEns0.0 $0.298\scriptstyle{\pm 0.041}$ $0.607\scriptstyle{\pm 0.037}$ $0.000\scriptstyle{\pm 0.000}$ $0.963\scriptstyle{\pm 0.000}$ $0.329\scriptstyle{\pm 0.040}$ $0.611\scriptstyle{\pm 0.023}$
-CreEns0.2 $0.165\scriptstyle{\pm 0.007}$ $0.769\scriptstyle{\pm 0.010}$ $0.000\scriptstyle{\pm 0.000}$ $0.983\scriptstyle{\pm 0.000}$ $0.177\scriptstyle{\pm 0.060}$ $0.787\scriptstyle{\pm 0.020}$
-CreEns0.4 $0.096\scriptstyle{\pm 0.008}$ $0.834\scriptstyle{\pm 0.006}$ $0.000\scriptstyle{\pm 0.000}$ $0.986\scriptstyle{\pm 0.000}$ $0.118\scriptstyle{\pm 0.059}$ $0.844\scriptstyle{\pm 0.021}$
-CreEns0.6 $0.043\scriptstyle{\pm 0.003}$ $0.886\scriptstyle{\pm 0.005}$ $0.000\scriptstyle{\pm 0.000}$ $0.988\scriptstyle{\pm 0.000}$ $0.081\scriptstyle{\pm 0.057}$ $0.899\scriptstyle{\pm 0.012}$
-CreEns0.8 $0.012\scriptstyle{\pm 0.002}$ $0.941\scriptstyle{\pm 0.002}$ $0.000\scriptstyle{\pm 0.000}$ $0.993\scriptstyle{\pm 0.000}$ $0.054\scriptstyle{\pm 0.033}$ $0.950\scriptstyle{\pm 0.008}$
-CreEns0.9 $0.000\scriptstyle{\pm 0.000}$ $0.976\scriptstyle{\pm 0.001}$ $0.000\scriptstyle{\pm 0.000}$ $0.997\scriptstyle{\pm 0.000}$ $0.016\scriptstyle{\pm 0.023}$ $0.983\scriptstyle{\pm 0.002}$
+CreNet $0.001\scriptstyle{\pm 0.002}$ $978\scriptstyle{\pm 0.008} 0.0$$094\scriptstyle{\pm 0.010}$ $0.999\scriptstyle{\pm 0.000}$ $0.188\scriptstyle{\pm 0.020}$ $0.792\scriptstyle{\pm 0.003}$
+CreBNN $0.195\scriptstyle{\pm 0.018}$ $0.649\scriptstyle{\pm 0.017}$ $0.331\scriptstyle{\pm 0.005}$ $0.871\scriptstyle{\pm 0.003}$ $898\scriptstyle{\pm 0.133} 0.0$$090\scriptstyle{\pm 0.124}$
+CreEns0.0 $0.298\scriptstyle{\pm 0.041}$ $607\scriptstyle{\pm 0.037} 0.0$$000\scriptstyle{\pm 0.000}$ $0.963\scriptstyle{\pm 0.000}$ $0.329\scriptstyle{\pm 0.040}$ $0.611\scriptstyle{\pm 0.023}$
+CreEns0.2 $0.165\scriptstyle{\pm 0.007}$ $769\scriptstyle{\pm 0.010} 0.0$$000\scriptstyle{\pm 0.000}$ $0.983\scriptstyle{\pm 0.000}$ $0.177\scriptstyle{\pm 0.060}$ $0.787\scriptstyle{\pm 0.020}$
+CreEns0.4 $0.096\scriptstyle{\pm 0.008}$ $834\scriptstyle{\pm 0.006} 0.0$$000\scriptstyle{\pm 0.000}$ $0.986\scriptstyle{\pm 0.000}$ $0.118\scriptstyle{\pm 0.059}$ $0.844\scriptstyle{\pm 0.021}$
+CreEns0.6 $0.043\scriptstyle{\pm 0.003}$ $886\scriptstyle{\pm 0.005} 0.0$$000\scriptstyle{\pm 0.000}$ $988\scriptstyle{\pm 0.000} 0.0$$081\scriptstyle{\pm 0.057}$ $0.899\scriptstyle{\pm 0.012}$
+CreEns0.8 $0.012\scriptstyle{\pm 0.002}$ $941\scriptstyle{\pm 0.002} 0.0$$000\scriptstyle{\pm 0.000}$ $993\scriptstyle{\pm 0.000} 0.0$$054\scriptstyle{\pm 0.033}$ $0.950\scriptstyle{\pm 0.008}$
+CreEns0.9 $0.000\scriptstyle{\pm 0.000}$ $976\scriptstyle{\pm 0.001} 0.0$$000\scriptstyle{\pm 0.000}$ $997\scriptstyle{\pm 0.000} 0.0$$016\scriptstyle{\pm 0.023}$ $0.983\scriptstyle{\pm 0.002}$
 CreEns0.95 $0.000\scriptstyle{\pm 0.000}$ $1.000\scriptstyle{\pm 0.000}$ $0.000\scriptstyle{\pm 0.000}$ $1.000\scriptstyle{\pm 0.000}$ $0.000\scriptstyle{\pm 0.000}$ $1.000\scriptstyle{\pm 0.000}$
 Table 7: Accuracy per ensemble member of CreRLα for varying $\alpha$ values on ChaosNLI, CIFAR-10, and QualityMRI.
-ChaosNLI $\bm{\alpha}$ $\bm{0.0}$ $\bm{0.2}$ $\bm{0.4}$ $\bm{0.6}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{0.95}$ $\bm{1.0}$ Member $\bm{m}$ 1 $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ 2 $0.370\scriptstyle{\pm 0.012}$ $0.526\scriptstyle{\pm 0.075}$ $0.512\scriptstyle{\pm 0.063}$ $0.577\scriptstyle{\pm 0.011}$ $0.608\scriptstyle{\pm 0.055}$ $0.664\scriptstyle{\pm 0.028}$ $0.682\scriptstyle{\pm 0.016}$ $0.678\scriptstyle{\pm 0.019}$ 3 $0.453\scriptstyle{\pm 0.010}$ $0.549\scriptstyle{\pm 0.008}$ $0.407\scriptstyle{\pm 0.020}$ $0.604\scriptstyle{\pm 0.045}$ $0.634\scriptstyle{\pm 0.016}$ $0.668\scriptstyle{\pm 0.003}$ $0.663\scriptstyle{\pm 0.027}$ $0.657\scriptstyle{\pm 0.020}$ 4 $0.178\scriptstyle{\pm 0.014}$ $0.468\scriptstyle{\pm 0.022}$ $0.483\scriptstyle{\pm 0.045}$ $0.483\scriptstyle{\pm 0.046}$ $0.651\scriptstyle{\pm 0.016}$ $0.660\scriptstyle{\pm 0.012}$ $0.661\scriptstyle{\pm 0.010}$ $0.667\scriptstyle{\pm 0.006}$ 5 $0.370\scriptstyle{\pm 0.012}$ $0.568\scriptstyle{\pm 0.018}$ $0.425\scriptstyle{\pm 0.046}$ $0.560\scriptstyle{\pm 0.042}$ $0.624\scriptstyle{\pm 0.018}$ $0.656\scriptstyle{\pm 0.016}$ $0.677\scriptstyle{\pm 0.017}$ $0.686\scriptstyle{\pm 0.020}$ 6 $0.453\scriptstyle{\pm 0.010}$ $0.535\scriptstyle{\pm 0.015}$ $0.530\scriptstyle{\pm 0.023}$ $0.585\scriptstyle{\pm 0.025}$ $0.647\scriptstyle{\pm 0.017}$ $0.663\scriptstyle{\pm 0.012}$ $0.674\scriptstyle{\pm 0.007}$ $0.676\scriptstyle{\pm 0.006}$ 7 $0.178\scriptstyle{\pm 0.014}$ $0.453\scriptstyle{\pm 0.015}$ $0.506\scriptstyle{\pm 0.041}$ $0.594\scriptstyle{\pm 0.044}$ $0.637\scriptstyle{\pm 0.020}$ $0.647\scriptstyle{\pm 0.003}$ $0.664\scriptstyle{\pm 0.005}$ $0.675\scriptstyle{\pm 0.006}$ 8 $0.370\scriptstyle{\pm 0.012}$ $0.585\scriptstyle{\pm 0.027}$ $0.494\scriptstyle{\pm 0.119}$ $0.585\scriptstyle{\pm 0.039}$ $0.646\scriptstyle{\pm 0.038}$ $0.673\scriptstyle{\pm 0.022}$ $0.677\scriptstyle{\pm 0.024}$ $0.682\scriptstyle{\pm 0.014}$ 9 $0.453\scriptstyle{\pm 0.010}$ $0.493\scriptstyle{\pm 0.074}$ $0.595\scriptstyle{\pm 0.042}$ $0.604\scriptstyle{\pm 0.029}$ $0.663\scriptstyle{\pm 0.007}$ $0.668\scriptstyle{\pm 0.002}$ $0.661\scriptstyle{\pm 0.023}$ $0.670\scriptstyle{\pm 0.009}$ 10 $0.178\scriptstyle{\pm 0.014}$ $0.476\scriptstyle{\pm 0.022}$ $0.514\scriptstyle{\pm 0.050}$ $0.552\scriptstyle{\pm 0.054}$ $0.650\scriptstyle{\pm 0.004}$ $0.648\scriptstyle{\pm 0.008}$ $0.664\scriptstyle{\pm 0.013}$ $0.675\scriptstyle{\pm 0.008}$ 11 $0.370\scriptstyle{\pm 0.012}$ $0.533\scriptstyle{\pm 0.041}$ $0.584\scriptstyle{\pm 0.046}$ $0.587\scriptstyle{\pm 0.035}$ $0.676\scriptstyle{\pm 0.029}$ $0.684\scriptstyle{\pm 0.018}$ $0.684\scriptstyle{\pm 0.012}$ $0.683\scriptstyle{\pm 0.017}$ 12 $0.453\scriptstyle{\pm 0.010}$ $0.644\scriptstyle{\pm 0.011}$ $0.638\scriptstyle{\pm 0.015}$ $0.650\scriptstyle{\pm 0.011}$ $0.659\scriptstyle{\pm 0.011}$ $0.670\scriptstyle{\pm 0.013}$ $0.672\scriptstyle{\pm 0.005}$ $0.665\scriptstyle{\pm 0.019}$ 13 $0.178\scriptstyle{\pm 0.014}$ $0.510\scriptstyle{\pm 0.065}$ $0.584\scriptstyle{\pm 0.079}$ $0.633\scriptstyle{\pm 0.021}$ $0.652\scriptstyle{\pm 0.009}$ $0.664\scriptstyle{\pm 0.006}$ $0.668\scriptstyle{\pm 0.010}$ $0.670\scriptstyle{\pm 0.014}$ 14 $0.370\scriptstyle{\pm 0.012}$ $0.560\scriptstyle{\pm 0.040}$ $0.603\scriptstyle{\pm 0.015}$ $0.632\scriptstyle{\pm 0.031}$ $0.668\scriptstyle{\pm 0.020}$ $0.675\scriptstyle{\pm 0.014}$ $0.684\scriptstyle{\pm 0.014}$ $0.676\scriptstyle{\pm 0.011}$ 15 $0.453\scriptstyle{\pm 0.010}$ $0.644\scriptstyle{\pm 0.019}$ $0.643\scriptstyle{\pm 0.013}$ $0.662\scriptstyle{\pm 0.017}$ $0.679\scriptstyle{\pm 0.014}$ $0.685\scriptstyle{\pm 0.013}$ $0.677\scriptstyle{\pm 0.007}$ $0.670\scriptstyle{\pm 0.019}$ 16 $0.178\scriptstyle{\pm 0.014}$ $0.630\scriptstyle{\pm 0.021}$ $0.637\scriptstyle{\pm 0.001}$ $0.647\scriptstyle{\pm 0.013}$ $0.651\scriptstyle{\pm 0.005}$ $0.668\scriptstyle{\pm 0.017}$ $0.658\scriptstyle{\pm 0.007}$ $0.671\scriptstyle{\pm 0.002}$ 17 $0.370\scriptstyle{\pm 0.012}$ $0.623\scriptstyle{\pm 0.045}$ $0.647\scriptstyle{\pm 0.017}$ $0.655\scriptstyle{\pm 0.010}$ $0.684\scriptstyle{\pm 0.014}$ $0.688\scriptstyle{\pm 0.024}$ $0.685\scriptstyle{\pm 0.020}$ $0.687\scriptstyle{\pm 0.010}$ 18 $0.453\scriptstyle{\pm 0.010}$ $0.652\scriptstyle{\pm 0.005}$ $0.655\scriptstyle{\pm 0.009}$ $0.664\scriptstyle{\pm 0.014}$ $0.684\scriptstyle{\pm 0.008}$ $0.669\scriptstyle{\pm 0.011}$ $0.663\scriptstyle{\pm 0.009}$ $0.663\scriptstyle{\pm 0.005}$ 19 $0.178\scriptstyle{\pm 0.014}$ $0.650\scriptstyle{\pm 0.008}$ $0.659\scriptstyle{\pm 0.003}$ $0.663\scriptstyle{\pm 0.015}$ $0.673\scriptstyle{\pm 0.006}$ $0.676\scriptstyle{\pm 0.005}$ $0.667\scriptstyle{\pm 0.001}$ $0.667\scriptstyle{\pm 0.010}$ 20 $0.370\scriptstyle{\pm 0.012}$ $0.689\scriptstyle{\pm 0.017}$ $0.677\scriptstyle{\pm 0.014}$ $0.679\scriptstyle{\pm 0.024}$ $0.684\scriptstyle{\pm 0.029}$ $0.688\scriptstyle{\pm 0.024}$ $0.671\scriptstyle{\pm 0.026}$ $0.673\scriptstyle{\pm 0.033}$ CIFAR-10 $\bm{\alpha}$ $\bm{0.0}$ $\bm{0.2}$ $\bm{0.4}$ $\bm{0.6}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{0.95}$ $\bm{1.0}$ Member $\bm{m}$ 1 $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.943\scriptstyle{\pm 0.001}$ 2 $0.100\scriptstyle{\pm 0.000}$ $0.490\scriptstyle{\pm 0.027}$ $0.711\scriptstyle{\pm 0.030}$ $0.792\scriptstyle{\pm 0.014}$ $0.860\scriptstyle{\pm 0.003}$ $0.889\scriptstyle{\pm 0.003}$ $0.898\scriptstyle{\pm 0.005}$ $0.934\scriptstyle{\pm 0.001}$ 3 $0.100\scriptstyle{\pm 0.000}$ $0.582\scriptstyle{\pm 0.043}$ $0.725\scriptstyle{\pm 0.027}$ $0.809\scriptstyle{\pm 0.004}$ $0.873\scriptstyle{\pm 0.008}$ $0.891\scriptstyle{\pm 0.002}$ $0.898\scriptstyle{\pm 0.004}$ $0.934\scriptstyle{\pm 0.002}$ 4 $0.100\scriptstyle{\pm 0.000}$ $0.591\scriptstyle{\pm 0.009}$ $0.722\scriptstyle{\pm 0.008}$ $0.816\scriptstyle{\pm 0.009}$ $0.871\scriptstyle{\pm 0.008}$ $0.891\scriptstyle{\pm 0.000}$ $0.902\scriptstyle{\pm 0.001}$ $0.936\scriptstyle{\pm 0.001}$ 5 $0.099\scriptstyle{\pm 0.000}$ $0.645\scriptstyle{\pm 0.036}$ $0.754\scriptstyle{\pm 0.012}$ $0.837\scriptstyle{\pm 0.011}$ $0.874\scriptstyle{\pm 0.002}$ $0.886\scriptstyle{\pm 0.004}$ $0.906\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.002}$ 6 $0.098\scriptstyle{\pm 0.000}$ $0.704\scriptstyle{\pm 0.010}$ $0.775\scriptstyle{\pm 0.002}$ $0.843\scriptstyle{\pm 0.012}$ $0.879\scriptstyle{\pm 0.006}$ $0.891\scriptstyle{\pm 0.003}$ $0.903\scriptstyle{\pm 0.003}$ $0.934\scriptstyle{\pm 0.002}$ 7 $0.100\scriptstyle{\pm 0.000}$ $0.697\scriptstyle{\pm 0.012}$ $0.782\scriptstyle{\pm 0.004}$ $0.851\scriptstyle{\pm 0.008}$ $0.881\scriptstyle{\pm 0.003}$ $0.893\scriptstyle{\pm 0.006}$ $0.902\scriptstyle{\pm 0.002}$ $0.936\scriptstyle{\pm 0.001}$ 8 $0.100\scriptstyle{\pm 0.000}$ $0.748\scriptstyle{\pm 0.015}$ $0.801\scriptstyle{\pm 0.011}$ $0.850\scriptstyle{\pm 0.006}$ $0.880\scriptstyle{\pm 0.006}$ $0.897\scriptstyle{\pm 0.003}$ $0.906\scriptstyle{\pm 0.004}$ $0.935\scriptstyle{\pm 0.001}$ 9 $0.101\scriptstyle{\pm 0.000}$ $0.765\scriptstyle{\pm 0.025}$ $0.813\scriptstyle{\pm 0.020}$ $0.848\scriptstyle{\pm 0.004}$ $0.890\scriptstyle{\pm 0.004}$ $0.896\scriptstyle{\pm 0.005}$ $0.908\scriptstyle{\pm 0.001}$ $0.936\scriptstyle{\pm 0.002}$ 10 $0.100\scriptstyle{\pm 0.000}$ $0.779\scriptstyle{\pm 0.007}$ $0.830\scriptstyle{\pm 0.007}$ $0.862\scriptstyle{\pm 0.011}$ $0.886\scriptstyle{\pm 0.012}$ $0.895\scriptstyle{\pm 0.004}$ $0.904\scriptstyle{\pm 0.003}$ $0.936\scriptstyle{\pm 0.001}$ 11 $0.100\scriptstyle{\pm 0.000}$ $0.790\scriptstyle{\pm 0.003}$ $0.831\scriptstyle{\pm 0.003}$ $0.865\scriptstyle{\pm 0.005}$ $0.889\scriptstyle{\pm 0.007}$ $0.904\scriptstyle{\pm 0.001}$ $0.908\scriptstyle{\pm 0.004}$ $0.936\scriptstyle{\pm 0.001}$ 12 $0.100\scriptstyle{\pm 0.000}$ $0.821\scriptstyle{\pm 0.009}$ $0.854\scriptstyle{\pm 0.013}$ $0.863\scriptstyle{\pm 0.008}$ $0.890\scriptstyle{\pm 0.002}$ $0.902\scriptstyle{\pm 0.004}$ $0.907\scriptstyle{\pm 0.003}$ $0.934\scriptstyle{\pm 0.002}$ 13 $0.100\scriptstyle{\pm 0.000}$ $0.824\scriptstyle{\pm 0.001}$ $0.850\scriptstyle{\pm 0.002}$ $0.875\scriptstyle{\pm 0.003}$ $0.884\scriptstyle{\pm 0.009}$ $0.904\scriptstyle{\pm 0.003}$ $0.907\scriptstyle{\pm 0.002}$ $0.937\scriptstyle{\pm 0.002}$ 14 $0.100\scriptstyle{\pm 0.000}$ $0.840\scriptstyle{\pm 0.002}$ $0.860\scriptstyle{\pm 0.008}$ $0.874\scriptstyle{\pm 0.009}$ $0.891\scriptstyle{\pm 0.002}$ $0.906\scriptstyle{\pm 0.006}$ $0.912\scriptstyle{\pm 0.004}$ $0.934\scriptstyle{\pm 0.002}$ 15 $0.099\scriptstyle{\pm 0.000}$ $0.849\scriptstyle{\pm 0.007}$ $0.868\scriptstyle{\pm 0.006}$ $0.879\scriptstyle{\pm 0.004}$ $0.898\scriptstyle{\pm 0.003}$ $0.905\scriptstyle{\pm 0.001}$ $0.911\scriptstyle{\pm 0.000}$ $0.933\scriptstyle{\pm 0.002}$ 16 $0.098\scriptstyle{\pm 0.000}$ $0.868\scriptstyle{\pm 0.009}$ $0.873\scriptstyle{\pm 0.006}$ $0.887\scriptstyle{\pm 0.004}$ $0.904\scriptstyle{\pm 0.004}$ $0.910\scriptstyle{\pm 0.003}$ $0.911\scriptstyle{\pm 0.002}$ $0.936\scriptstyle{\pm 0.002}$ 17 $0.100\scriptstyle{\pm 0.000}$ $0.863\scriptstyle{\pm 0.003}$ $0.879\scriptstyle{\pm 0.003}$ $0.888\scriptstyle{\pm 0.004}$ $0.901\scriptstyle{\pm 0.005}$ $0.908\scriptstyle{\pm 0.002}$ $0.909\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.001}$ 18 $0.100\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.004}$ $0.892\scriptstyle{\pm 0.005}$ $0.898\scriptstyle{\pm 0.002}$ $0.902\scriptstyle{\pm 0.005}$ $0.913\scriptstyle{\pm 0.005}$ $0.916\scriptstyle{\pm 0.001}$ $0.934\scriptstyle{\pm 0.001}$ 19 $0.101\scriptstyle{\pm 0.000}$ $0.887\scriptstyle{\pm 0.007}$ $0.897\scriptstyle{\pm 0.004}$ $0.902\scriptstyle{\pm 0.001}$ $0.909\scriptstyle{\pm 0.002}$ $0.915\scriptstyle{\pm 0.002}$ $0.915\scriptstyle{\pm 0.001}$ $0.936\scriptstyle{\pm 0.001}$ 20 $0.100\scriptstyle{\pm 0.000}$ $0.903\scriptstyle{\pm 0.005}$ $0.904\scriptstyle{\pm 0.003}$ $0.908\scriptstyle{\pm 0.002}$ $0.910\scriptstyle{\pm 0.004}$ $0.920\scriptstyle{\pm 0.002}$ $0.921\scriptstyle{\pm 0.001}$ $0.935\scriptstyle{\pm 0.002}$ QualityMRI $\bm{\alpha}$ $\bm{0.0}$ $\bm{0.2}$ $\bm{0.4}$ $\bm{0.6}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{0.95}$ $\bm{1.0}$ Member $\bm{m}$ 1 $0.624\scriptstyle{\pm 0.033}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.624\scriptstyle{\pm 0.033}$ $0.532\scriptstyle{\pm 0.105}$ 2 $0.376\scriptstyle{\pm 0.008}$ $0.511\scriptstyle{\pm 0.050}$ $0.532\scriptstyle{\pm 0.023}$ $0.570\scriptstyle{\pm 0.046}$ $0.575\scriptstyle{\pm 0.020}$ $0.597\scriptstyle{\pm 0.057}$ $0.586\scriptstyle{\pm 0.055}$ $0.495\scriptstyle{\pm 0.119}$ 3 $0.624\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.015}$ $0.672\scriptstyle{\pm 0.027}$ $0.629\scriptstyle{\pm 0.026}$ $0.618\scriptstyle{\pm 0.015}$ $0.624\scriptstyle{\pm 0.055}$ $0.656\scriptstyle{\pm 0.042}$ $0.570\scriptstyle{\pm 0.075}$ 4 $0.376\scriptstyle{\pm 0.008}$ $0.586\scriptstyle{\pm 0.015}$ $0.548\scriptstyle{\pm 0.103}$ $0.543\scriptstyle{\pm 0.020}$ $0.581\scriptstyle{\pm 0.013}$ $0.618\scriptstyle{\pm 0.027}$ $0.618\scriptstyle{\pm 0.053}$ $0.511\scriptstyle{\pm 0.112}$ 5 $0.624\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.027}$ $0.645\scriptstyle{\pm 0.035}$ $0.602\scriptstyle{\pm 0.042}$ $0.597\scriptstyle{\pm 0.070}$ $0.629\scriptstyle{\pm 0.013}$ $0.651\scriptstyle{\pm 0.038}$ $0.570\scriptstyle{\pm 0.101}$ 6 $0.376\scriptstyle{\pm 0.008}$ $0.543\scriptstyle{\pm 0.020}$ $0.586\scriptstyle{\pm 0.040}$ $0.581\scriptstyle{\pm 0.023}$ $0.591\scriptstyle{\pm 0.020}$ $0.570\scriptstyle{\pm 0.008}$ $0.597\scriptstyle{\pm 0.035}$ $0.532\scriptstyle{\pm 0.103}$ 7 $0.624\scriptstyle{\pm 0.008}$ $0.618\scriptstyle{\pm 0.042}$ $0.624\scriptstyle{\pm 0.020}$ $0.608\scriptstyle{\pm 0.020}$ $0.618\scriptstyle{\pm 0.030}$ $0.640\scriptstyle{\pm 0.042}$ $0.629\scriptstyle{\pm 0.013}$ $0.511\scriptstyle{\pm 0.124}$ 8 $0.376\scriptstyle{\pm 0.008}$ $0.597\scriptstyle{\pm 0.099}$ $0.586\scriptstyle{\pm 0.027}$ $0.575\scriptstyle{\pm 0.046}$ $0.597\scriptstyle{\pm 0.060}$ $0.591\scriptstyle{\pm 0.027}$ $0.602\scriptstyle{\pm 0.020}$ $0.500\scriptstyle{\pm 0.115}$ 9 $0.624\scriptstyle{\pm 0.008}$ $0.640\scriptstyle{\pm 0.020}$ $0.651\scriptstyle{\pm 0.055}$ $0.683\scriptstyle{\pm 0.046}$ $0.613\scriptstyle{\pm 0.023}$ $0.645\scriptstyle{\pm 0.013}$ $0.629\scriptstyle{\pm 0.013}$ $0.522\scriptstyle{\pm 0.107}$ 10 $0.376\scriptstyle{\pm 0.008}$ $0.565\scriptstyle{\pm 0.035}$ $0.602\scriptstyle{\pm 0.008}$ $0.548\scriptstyle{\pm 0.035}$ $0.602\scriptstyle{\pm 0.059}$ $0.591\scriptstyle{\pm 0.038}$ $0.570\scriptstyle{\pm 0.008}$ $0.505\scriptstyle{\pm 0.095}$ 11 $0.624\scriptstyle{\pm 0.008}$ $0.629\scriptstyle{\pm 0.026}$ $0.624\scriptstyle{\pm 0.020}$ $0.608\scriptstyle{\pm 0.027}$ $0.591\scriptstyle{\pm 0.033}$ $0.613\scriptstyle{\pm 0.066}$ $0.581\scriptstyle{\pm 0.013}$ $0.565\scriptstyle{\pm 0.068}$ 12 $0.376\scriptstyle{\pm 0.008}$ $0.538\scriptstyle{\pm 0.015}$ $0.570\scriptstyle{\pm 0.055}$ $0.640\scriptstyle{\pm 0.008}$ $0.591\scriptstyle{\pm 0.008}$ $0.602\scriptstyle{\pm 0.038}$ $0.570\scriptstyle{\pm 0.030}$ $0.538\scriptstyle{\pm 0.107}$ 13 $0.624\scriptstyle{\pm 0.008}$ $0.688\scriptstyle{\pm 0.053}$ $0.613\scriptstyle{\pm 0.013}$ $0.677\scriptstyle{\pm 0.023}$ $0.656\scriptstyle{\pm 0.008}$ $0.602\scriptstyle{\pm 0.042}$ $0.618\scriptstyle{\pm 0.042}$ $0.559\scriptstyle{\pm 0.100}$ 14 $0.376\scriptstyle{\pm 0.008}$ $0.532\scriptstyle{\pm 0.035}$ $0.581\scriptstyle{\pm 0.035}$ $0.570\scriptstyle{\pm 0.046}$ $0.586\scriptstyle{\pm 0.020}$ $0.591\scriptstyle{\pm 0.020}$ $0.554\scriptstyle{\pm 0.033}$ $0.538\scriptstyle{\pm 0.127}$ 15 $0.624\scriptstyle{\pm 0.008}$ $0.634\scriptstyle{\pm 0.040}$ $0.651\scriptstyle{\pm 0.008}$ $0.586\scriptstyle{\pm 0.050}$ $0.651\scriptstyle{\pm 0.059}$ $0.629\scriptstyle{\pm 0.023}$ $0.618\scriptstyle{\pm 0.050}$ $0.559\scriptstyle{\pm 0.112}$ 16 $0.376\scriptstyle{\pm 0.008}$ $0.565\scriptstyle{\pm 0.023}$ $0.618\scriptstyle{\pm 0.040}$ $0.586\scriptstyle{\pm 0.062}$ $0.570\scriptstyle{\pm 0.030}$ $0.597\scriptstyle{\pm 0.026}$ $0.554\scriptstyle{\pm 0.008}$ $0.516\scriptstyle{\pm 0.092}$ 17 $0.624\scriptstyle{\pm 0.008}$ $0.624\scriptstyle{\pm 0.040}$ $0.677\scriptstyle{\pm 0.035}$ $0.629\scriptstyle{\pm 0.047}$ $0.618\scriptstyle{\pm 0.027}$ $0.591\scriptstyle{\pm 0.053}$ $0.624\scriptstyle{\pm 0.040}$ $0.554\scriptstyle{\pm 0.066}$ 18 $0.376\scriptstyle{\pm 0.008}$ $0.575\scriptstyle{\pm 0.055}$ $0.586\scriptstyle{\pm 0.008}$ $0.597\scriptstyle{\pm 0.013}$ $0.591\scriptstyle{\pm 0.020}$ $0.570\scriptstyle{\pm 0.008}$ $0.602\scriptstyle{\pm 0.027}$ $0.522\scriptstyle{\pm 0.119}$ 19 $0.624\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.050}$ $0.656\scriptstyle{\pm 0.046}$ $0.634\scriptstyle{\pm 0.055}$ $0.629\scriptstyle{\pm 0.040}$ $0.624\scriptstyle{\pm 0.038}$ $0.618\scriptstyle{\pm 0.040}$ $0.527\scriptstyle{\pm 0.077}$ 20 $0.376\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.027}$ $0.570\scriptstyle{\pm 0.065}$ $0.608\scriptstyle{\pm 0.020}$ $0.597\scriptstyle{\pm 0.026}$ $0.608\scriptstyle{\pm 0.020}$ $0.554\scriptstyle{\pm 0.020}$ $0.548\scriptstyle{\pm 0.115}$
+ChaosNLI $\bm{\alpha}$ $\bm{0.0}$ $\bm{0.2}$ $\bm{0.4}$ $\bm{0.6}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{0.95}$ $\bm{1.0}$ Member $\bm{m}$ 1 $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ $0.678\scriptstyle{\pm 0.017}$ 2 $0.370\scriptstyle{\pm 0.012}$ $0.526\scriptstyle{\pm 0.075}$ $0.512\scriptstyle{\pm 0.063}$ $0.577\scriptstyle{\pm 0.011}$ $0.608\scriptstyle{\pm 0.055}$ $0.664\scriptstyle{\pm 0.028}$ $0.682\scriptstyle{\pm 0.016}$ $0.678\scriptstyle{\pm 0.019}$ 3 $0.453\scriptstyle{\pm 0.010}$ $0.549\scriptstyle{\pm 0.008}$ $0.407\scriptstyle{\pm 0.020}$ $0.604\scriptstyle{\pm 0.045}$ $0.634\scriptstyle{\pm 0.016}$ $0.668\scriptstyle{\pm 0.003}$ $0.663\scriptstyle{\pm 0.027}$ $0.657\scriptstyle{\pm 0.020}$ 4 $0.178\scriptstyle{\pm 0.014}$ $0.468\scriptstyle{\pm 0.022}$ $0.483\scriptstyle{\pm 0.045}$ $0.483\scriptstyle{\pm 0.046}$ $51\scriptstyle{\pm 0.016} 0.660$$60\scriptstyle{\pm 0.012}$ $0.661\scriptstyle{\pm 0.010}$ $0.667\scriptstyle{\pm 0.006}$ 5 $0.370\scriptstyle{\pm 0.012}$ $0.568\scriptstyle{\pm 0.018}$ $0.425\scriptstyle{\pm 0.046}$ $0.560\scriptstyle{\pm 0.042}$ $0.624\scriptstyle{\pm 0.018}$ $0.656\scriptstyle{\pm 0.016}$ $0.677\scriptstyle{\pm 0.017}$ $0.686\scriptstyle{\pm 0.020}$ 6 $0.453\scriptstyle{\pm 0.010}$ $0.535\scriptstyle{\pm 0.015}$ $0.530\scriptstyle{\pm 0.023}$ $0.585\scriptstyle{\pm 0.025}$ $0.647\scriptstyle{\pm 0.017}$ $0.663\scriptstyle{\pm 0.012}$ $0.674\scriptstyle{\pm 0.007}$ $0.676\scriptstyle{\pm 0.006}$ 7 $0.178\scriptstyle{\pm 0.014}$ $0.453\scriptstyle{\pm 0.015}$ $0.506\scriptstyle{\pm 0.041}$ $0.594\scriptstyle{\pm 0.044}$ $0.637\scriptstyle{\pm 0.020}$ $0.647\scriptstyle{\pm 0.003}$ $0.664\scriptstyle{\pm 0.005}$ $0.675\scriptstyle{\pm 0.006}$ 8 $0.370\scriptstyle{\pm 0.012}$ $0.585\scriptstyle{\pm 0.027}$ $0.494\scriptstyle{\pm 0.119}$ $0.585\scriptstyle{\pm 0.039}$ $0.646\scriptstyle{\pm 0.038}$ $0.673\scriptstyle{\pm 0.022}$ $0.677\scriptstyle{\pm 0.024}$ $0.682\scriptstyle{\pm 0.014}$ 9 $0.453\scriptstyle{\pm 0.010}$ $0.493\scriptstyle{\pm 0.074}$ $0.595\scriptstyle{\pm 0.042}$ $0.604\scriptstyle{\pm 0.029}$ $0.663\scriptstyle{\pm 0.007}$ $0.668\scriptstyle{\pm 0.002}$ $0.661\scriptstyle{\pm 0.023}$ $0.670\scriptstyle{\pm 0.009}$ 10 $0.178\scriptstyle{\pm 0.014}$ $0.476\scriptstyle{\pm 0.022}$ $0.514\scriptstyle{\pm 0.050}$ $0.552\scriptstyle{\pm 0.054}$ $0.650\scriptstyle{\pm 0.004}$ $0.648\scriptstyle{\pm 0.008}$ $0.664\scriptstyle{\pm 0.013}$ $0.675\scriptstyle{\pm 0.008}$ 11 $0.370\scriptstyle{\pm 0.012}$ $0.533\scriptstyle{\pm 0.041}$ $0.584\scriptstyle{\pm 0.046}$ $0.587\scriptstyle{\pm 0.035}$ $0.676\scriptstyle{\pm 0.029}$ $0.684\scriptstyle{\pm 0.018}$ $0.684\scriptstyle{\pm 0.012}$ $0.683\scriptstyle{\pm 0.017}$ 12 $0.453\scriptstyle{\pm 0.010}$ $0.644\scriptstyle{\pm 0.011}$ $0.638\scriptstyle{\pm 0.015}$ $0.650\scriptstyle{\pm 0.011}$ $0.659\scriptstyle{\pm 0.011}$ $0.670\scriptstyle{\pm 0.013}$ $0.672\scriptstyle{\pm 0.005}$ $0.665\scriptstyle{\pm 0.019}$ 13 $0.178\scriptstyle{\pm 0.014}$ $0.510\scriptstyle{\pm 0.065}$ $0.584\scriptstyle{\pm 0.079}$ $0.633\scriptstyle{\pm 0.021}$ $0.652\scriptstyle{\pm 0.009}$ $0.664\scriptstyle{\pm 0.006}$ $0.668\scriptstyle{\pm 0.010}$ $0.670\scriptstyle{\pm 0.014}$ 14 $0.370\scriptstyle{\pm 0.012}$ $0.560\scriptstyle{\pm 0.040}$ $0.603\scriptstyle{\pm 0.015}$ $0.632\scriptstyle{\pm 0.031}$ $0.668\scriptstyle{\pm 0.020}$ $0.675\scriptstyle{\pm 0.014}$ $0.684\scriptstyle{\pm 0.014}$ $0.676\scriptstyle{\pm 0.011}$ 15 $0.453\scriptstyle{\pm 0.010}$ $0.644\scriptstyle{\pm 0.019}$ $0.643\scriptstyle{\pm 0.013}$ $0.662\scriptstyle{\pm 0.017}$ $0.679\scriptstyle{\pm 0.014}$ $0.685\scriptstyle{\pm 0.013}$ $0.677\scriptstyle{\pm 0.007}$ $0.670\scriptstyle{\pm 0.019}$ 16 $0.178\scriptstyle{\pm 0.014}$ $0.630\scriptstyle{\pm 0.021}$ $0.637\scriptstyle{\pm 0.001}$ $0.647\scriptstyle{\pm 0.013}$ $0.651\scriptstyle{\pm 0.005}$ $0.668\scriptstyle{\pm 0.017}$ $0.658\scriptstyle{\pm 0.007}$ $0.671\scriptstyle{\pm 0.002}$ 17 $0.370\scriptstyle{\pm 0.012}$ $0.623\scriptstyle{\pm 0.045}$ $0.647\scriptstyle{\pm 0.017}$ $0.655\scriptstyle{\pm 0.010}$ $0.684\scriptstyle{\pm 0.014}$ $0.688\scriptstyle{\pm 0.024}$ $0.685\scriptstyle{\pm 0.020}$ $0.687\scriptstyle{\pm 0.010}$ 18 $0.453\scriptstyle{\pm 0.010}$ $0.652\scriptstyle{\pm 0.005}$ $0.655\scriptstyle{\pm 0.009}$ $0.664\scriptstyle{\pm 0.014}$ $0.684\scriptstyle{\pm 0.008}$ $0.669\scriptstyle{\pm 0.011}$ $0.663\scriptstyle{\pm 0.009}$ $0.663\scriptstyle{\pm 0.005}$ 19 $0.178\scriptstyle{\pm 0.014}$ $0.650\scriptstyle{\pm 0.008}$ $0.659\scriptstyle{\pm 0.003}$ $0.663\scriptstyle{\pm 0.015}$ $0.673\scriptstyle{\pm 0.006}$ $0.676\scriptstyle{\pm 0.005}$ $0.667\scriptstyle{\pm 0.001}$ $0.667\scriptstyle{\pm 0.010}$ 20 $0.370\scriptstyle{\pm 0.012}$ $0.689\scriptstyle{\pm 0.017}$ $0.677\scriptstyle{\pm 0.014}$ $0.679\scriptstyle{\pm 0.024}$ $0.684\scriptstyle{\pm 0.029}$ $0.688\scriptstyle{\pm 0.024}$ $0.671\scriptstyle{\pm 0.026}$ $0.673\scriptstyle{\pm 0.033}$ CIFAR-10 $\bm{\alpha}$ $\bm{0.0}$ $\bm{0.2}$ $\bm{0.4}$ $\bm{0.6}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{0.95}$ $\bm{1.0}$ Member $\bm{m}$ 1 $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.003}$ $0.943\scriptstyle{\pm 0.001}$ 2 $0.100\scriptstyle{\pm 0.000}$ $0.490\scriptstyle{\pm 0.027}$ $0.711\scriptstyle{\pm 0.030}$ $0.792\scriptstyle{\pm 0.014}$ $0.860\scriptstyle{\pm 0.003}$ $0.889\scriptstyle{\pm 0.003}$ $0.898\scriptstyle{\pm 0.005}$ $0.934\scriptstyle{\pm 0.001}$ 3 $0.100\scriptstyle{\pm 0.000}$ $0.582\scriptstyle{\pm 0.043}$ $0.725\scriptstyle{\pm 0.027}$ $0.809\scriptstyle{\pm 0.004}$ $0.873\scriptstyle{\pm 0.008}$ $0.891\scriptstyle{\pm 0.002}$ $0.898\scriptstyle{\pm 0.004}$ $0.934\scriptstyle{\pm 0.002}$ 4 $0.100\scriptstyle{\pm 0.000}$ $0.591\scriptstyle{\pm 0.009}$ $0.722\scriptstyle{\pm 0.008}$ $0.816\scriptstyle{\pm 0.009}$ $0.871\scriptstyle{\pm 0.008}$ $0.891\scriptstyle{\pm 0.000}$ $0.902\scriptstyle{\pm 0.001}$ $0.936\scriptstyle{\pm 0.001}$ 5 $0.099\scriptstyle{\pm 0.000}$ $0.645\scriptstyle{\pm 0.036}$ $0.754\scriptstyle{\pm 0.012}$ $0.837\scriptstyle{\pm 0.011}$ $0.874\scriptstyle{\pm 0.002}$ $0.886\scriptstyle{\pm 0.004}$ $0.906\scriptstyle{\pm 0.003}$ $0.935\scriptstyle{\pm 0.002}$ 6 $0.098\scriptstyle{\pm 0.000}$ $0.704\scriptstyle{\pm 0.010}$ $0.775\scriptstyle{\pm 0.002}$ $0.843\scriptstyle{\pm 0.012}$ $0.879\scriptstyle{\pm 0.006}$ $0.891\scriptstyle{\pm 0.003}$ $0.903\scriptstyle{\pm 0.003}$ $0.934\scriptstyle{\pm 0.002}$ 7 $0.100\scriptstyle{\pm 0.000}$ $0.697\scriptstyle{\pm 0.012}$ $0.782\scriptstyle{\pm 0.004}$ $0.851\scriptstyle{\pm 0.008}$ $0.881\scriptstyle{\pm 0.003}$ $0.893\scriptstyle{\pm 0.006}$ $0.902\scriptstyle{\pm 0.002}$ $0.936\scriptstyle{\pm 0.001}$ 8 $0.100\scriptstyle{\pm 0.000}$ $0.748\scriptstyle{\pm 0.015}$ $0.801\scriptstyle{\pm 0.011}$ $50\scriptstyle{\pm 0.006} 0.880$$80\scriptstyle{\pm 0.006}$ $0.897\scriptstyle{\pm 0.003}$ $0.906\scriptstyle{\pm 0.004}$ $0.935\scriptstyle{\pm 0.001}$ 9 $0.101\scriptstyle{\pm 0.000}$ $0.765\scriptstyle{\pm 0.025}$ $0.813\scriptstyle{\pm 0.020}$ $0.848\scriptstyle{\pm 0.004}$ $0.890\scriptstyle{\pm 0.004}$ $0.896\scriptstyle{\pm 0.005}$ $0.908\scriptstyle{\pm 0.001}$ $0.936\scriptstyle{\pm 0.002}$ 10 $0.100\scriptstyle{\pm 0.000}$ $0.779\scriptstyle{\pm 0.007}$ $0.830\scriptstyle{\pm 0.007}$ $0.862\scriptstyle{\pm 0.011}$ $0.886\scriptstyle{\pm 0.012}$ $0.895\scriptstyle{\pm 0.004}$ $0.904\scriptstyle{\pm 0.003}$ $0.936\scriptstyle{\pm 0.001}$ 11 $0.100\scriptstyle{\pm 0.000}$ $0.790\scriptstyle{\pm 0.003}$ $0.831\scriptstyle{\pm 0.003}$ $0.865\scriptstyle{\pm 0.005}$ $0.889\scriptstyle{\pm 0.007}$ $0.904\scriptstyle{\pm 0.001}$ $0.908\scriptstyle{\pm 0.004}$ $0.936\scriptstyle{\pm 0.001}$ 12 $0.100\scriptstyle{\pm 0.000}$ $0.821\scriptstyle{\pm 0.009}$ $0.854\scriptstyle{\pm 0.013}$ $0.863\scriptstyle{\pm 0.008}$ $0.890\scriptstyle{\pm 0.002}$ $0.902\scriptstyle{\pm 0.004}$ $0.907\scriptstyle{\pm 0.003}$ $0.934\scriptstyle{\pm 0.002}$ 13 $0.100\scriptstyle{\pm 0.000}$ $0.824\scriptstyle{\pm 0.001}$ $0.850\scriptstyle{\pm 0.002}$ $0.875\scriptstyle{\pm 0.003}$ $0.884\scriptstyle{\pm 0.009}$ $0.904\scriptstyle{\pm 0.003}$ $0.907\scriptstyle{\pm 0.002}$ $0.937\scriptstyle{\pm 0.002}$ 14 $0.100\scriptstyle{\pm 0.000}$ $0.840\scriptstyle{\pm 0.002}$ $0.860\scriptstyle{\pm 0.008}$ $0.874\scriptstyle{\pm 0.009}$ $0.891\scriptstyle{\pm 0.002}$ $0.906\scriptstyle{\pm 0.006}$ $0.912\scriptstyle{\pm 0.004}$ $0.934\scriptstyle{\pm 0.002}$ 15 $0.099\scriptstyle{\pm 0.000}$ $0.849\scriptstyle{\pm 0.007}$ $0.868\scriptstyle{\pm 0.006}$ $0.879\scriptstyle{\pm 0.004}$ $0.898\scriptstyle{\pm 0.003}$ $0.905\scriptstyle{\pm 0.001}$ $0.911\scriptstyle{\pm 0.000}$ $0.933\scriptstyle{\pm 0.002}$ 16 $0.098\scriptstyle{\pm 0.000}$ $0.868\scriptstyle{\pm 0.009}$ $0.873\scriptstyle{\pm 0.006}$ $0.887\scriptstyle{\pm 0.004}$ $0.904\scriptstyle{\pm 0.004}$ $0.910\scriptstyle{\pm 0.003}$ $0.911\scriptstyle{\pm 0.002}$ $0.936\scriptstyle{\pm 0.002}$ 17 $0.100\scriptstyle{\pm 0.000}$ $0.863\scriptstyle{\pm 0.003}$ $0.879\scriptstyle{\pm 0.003}$ $0.888\scriptstyle{\pm 0.004}$ $0.901\scriptstyle{\pm 0.005}$ $8\scriptstyle{\pm 0.002} 0.909±0.0$$9\scriptstyle{\pm 0.002}$ $0.935\scriptstyle{\pm 0.001}$ 18 $0.100\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.004}$ $0.892\scriptstyle{\pm 0.005}$ $0.898\scriptstyle{\pm 0.002}$ $0.902\scriptstyle{\pm 0.005}$ $0.913\scriptstyle{\pm 0.005}$ $0.916\scriptstyle{\pm 0.001}$ $0.934\scriptstyle{\pm 0.001}$ 19 $0.101\scriptstyle{\pm 0.000}$ $0.887\scriptstyle{\pm 0.007}$ $0.897\scriptstyle{\pm 0.004}$ $2\scriptstyle{\pm 0.001} 0.909±0.0$$9\scriptstyle{\pm 0.002}$ $0.915\scriptstyle{\pm 0.002}$ $0.915\scriptstyle{\pm 0.001}$ $0.936\scriptstyle{\pm 0.001}$ 20 $0.100\scriptstyle{\pm 0.000}$ $0.903\scriptstyle{\pm 0.005}$ $0.904\scriptstyle{\pm 0.003}$ $0.908\scriptstyle{\pm 0.002}$ $0.910\scriptstyle{\pm 0.004}$ $0.920\scriptstyle{\pm 0.002}$ $0.921\scriptstyle{\pm 0.001}$ $0.935\scriptstyle{\pm 0.002}$ QualityMRI $\bm{\alpha}$ $\bm{0.0}$ $\bm{0.2}$ $\bm{0.4}$ $\bm{0.6}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{0.95}$ $\bm{1.0}$ Member $\bm{m}$ 1 $0.624\scriptstyle{\pm 0.033}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.602\scriptstyle{\pm 0.020}$ $0.624\scriptstyle{\pm 0.033}$ $0.532\scriptstyle{\pm 0.105}$ 2 $0.376\scriptstyle{\pm 0.008}$ $0.511\scriptstyle{\pm 0.050}$ $0.532\scriptstyle{\pm 0.023}$ $0.570\scriptstyle{\pm 0.046}$ $0.575\scriptstyle{\pm 0.020}$ $0.597\scriptstyle{\pm 0.057}$ $0.586\scriptstyle{\pm 0.055}$ $0.495\scriptstyle{\pm 0.119}$ 3 $0.624\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.015}$ $0.672\scriptstyle{\pm 0.027}$ $0.629\scriptstyle{\pm 0.026}$ $0.618\scriptstyle{\pm 0.015}$ $0.624\scriptstyle{\pm 0.055}$ $0.656\scriptstyle{\pm 0.042}$ $0.570\scriptstyle{\pm 0.075}$ 4 $0.376\scriptstyle{\pm 0.008}$ $0.586\scriptstyle{\pm 0.015}$ $0.548\scriptstyle{\pm 0.103}$ $0.543\scriptstyle{\pm 0.020}$ $0.581\scriptstyle{\pm 0.013}$ $0.618\scriptstyle{\pm 0.027}$ $0.618\scriptstyle{\pm 0.053}$ $0.511\scriptstyle{\pm 0.112}$ 5 $0.624\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.027}$ $0.645\scriptstyle{\pm 0.035}$ $0.602\scriptstyle{\pm 0.042}$ $0.597\scriptstyle{\pm 0.070}$ $0.629\scriptstyle{\pm 0.013}$ $0.651\scriptstyle{\pm 0.038}$ $0.570\scriptstyle{\pm 0.101}$ 6 $0.376\scriptstyle{\pm 0.008}$ $0.543\scriptstyle{\pm 0.020}$ $0.586\scriptstyle{\pm 0.040}$ $0.581\scriptstyle{\pm 0.023}$ $0.591\scriptstyle{\pm 0.020}$ $0.570\scriptstyle{\pm 0.008}$ $0.597\scriptstyle{\pm 0.035}$ $0.532\scriptstyle{\pm 0.103}$ 7 $0.624\scriptstyle{\pm 0.008}$ $0.618\scriptstyle{\pm 0.042}$ $0.624\scriptstyle{\pm 0.020}$ $0.608\scriptstyle{\pm 0.020}$ $0.618\scriptstyle{\pm 0.030}$ $0.640\scriptstyle{\pm 0.042}$ $0.629\scriptstyle{\pm 0.013}$ $0.511\scriptstyle{\pm 0.124}$ 8 $0.376\scriptstyle{\pm 0.008}$ $0.597\scriptstyle{\pm 0.099}$ $0.586\scriptstyle{\pm 0.027}$ $0.575\scriptstyle{\pm 0.046}$ $0.597\scriptstyle{\pm 0.060}$ $0.591\scriptstyle{\pm 0.027}$ $0.602\scriptstyle{\pm 0.020}$ $0.500\scriptstyle{\pm 0.115}$ 9 $0.624\scriptstyle{\pm 0.008}$ $0.640\scriptstyle{\pm 0.020}$ $0.651\scriptstyle{\pm 0.055}$ $0.683\scriptstyle{\pm 0.046}$ $0.613\scriptstyle{\pm 0.023}$ $0.645\scriptstyle{\pm 0.013}$ $0.629\scriptstyle{\pm 0.013}$ $0.522\scriptstyle{\pm 0.107}$ 10 $0.376\scriptstyle{\pm 0.008}$ $0.565\scriptstyle{\pm 0.035}$ $0.602\scriptstyle{\pm 0.008}$ $0.548\scriptstyle{\pm 0.035}$ $0.602\scriptstyle{\pm 0.059}$ $0.591\scriptstyle{\pm 0.038}$ $70\scriptstyle{\pm 0.008} 0.505$$05\scriptstyle{\pm 0.095}$ 11 $0.624\scriptstyle{\pm 0.008}$ $0.629\scriptstyle{\pm 0.026}$ $0.624\scriptstyle{\pm 0.020}$ $0.608\scriptstyle{\pm 0.027}$ $0.591\scriptstyle{\pm 0.033}$ $0.613\scriptstyle{\pm 0.066}$ $0.581\scriptstyle{\pm 0.013}$ $0.565\scriptstyle{\pm 0.068}$ 12 $0.376\scriptstyle{\pm 0.008}$ $0.538\scriptstyle{\pm 0.015}$ $0.570\scriptstyle{\pm 0.055}$ $0.640\scriptstyle{\pm 0.008}$ $0.591\scriptstyle{\pm 0.008}$ $0.602\scriptstyle{\pm 0.038}$ $0.570\scriptstyle{\pm 0.030}$ $0.538\scriptstyle{\pm 0.107}$ 13 $0.624\scriptstyle{\pm 0.008}$ $0.688\scriptstyle{\pm 0.053}$ $0.613\scriptstyle{\pm 0.013}$ $0.677\scriptstyle{\pm 0.023}$ $0.656\scriptstyle{\pm 0.008}$ $0.602\scriptstyle{\pm 0.042}$ $0.618\scriptstyle{\pm 0.042}$ $0.559\scriptstyle{\pm 0.100}$ 14 $0.376\scriptstyle{\pm 0.008}$ $0.532\scriptstyle{\pm 0.035}$ $0.581\scriptstyle{\pm 0.035}$ $0.570\scriptstyle{\pm 0.046}$ $0.586\scriptstyle{\pm 0.020}$ $0.591\scriptstyle{\pm 0.020}$ $0.554\scriptstyle{\pm 0.033}$ $0.538\scriptstyle{\pm 0.127}$ 15 $0.624\scriptstyle{\pm 0.008}$ $0.634\scriptstyle{\pm 0.040}$ $0.651\scriptstyle{\pm 0.008}$ $0.586\scriptstyle{\pm 0.050}$ $0.651\scriptstyle{\pm 0.059}$ $0.629\scriptstyle{\pm 0.023}$ $0.618\scriptstyle{\pm 0.050}$ $0.559\scriptstyle{\pm 0.112}$ 16 $0.376\scriptstyle{\pm 0.008}$ $0.565\scriptstyle{\pm 0.023}$ $0.618\scriptstyle{\pm 0.040}$ $0.586\scriptstyle{\pm 0.062}$ $0.570\scriptstyle{\pm 0.030}$ $0.597\scriptstyle{\pm 0.026}$ $0.554\scriptstyle{\pm 0.008}$ $0.516\scriptstyle{\pm 0.092}$ 17 $0.624\scriptstyle{\pm 0.008}$ $0.624\scriptstyle{\pm 0.040}$ $0.677\scriptstyle{\pm 0.035}$ $0.629\scriptstyle{\pm 0.047}$ $0.618\scriptstyle{\pm 0.027}$ $0.591\scriptstyle{\pm 0.053}$ $0.624\scriptstyle{\pm 0.040}$ $0.554\scriptstyle{\pm 0.066}$ 18 $0.376\scriptstyle{\pm 0.008}$ $0.575\scriptstyle{\pm 0.055}$ $0.586\scriptstyle{\pm 0.008}$ $0.597\scriptstyle{\pm 0.013}$ $0.591\scriptstyle{\pm 0.020}$ $0.570\scriptstyle{\pm 0.008}$ $0.602\scriptstyle{\pm 0.027}$ $0.522\scriptstyle{\pm 0.119}$ 19 $0.624\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.050}$ $0.656\scriptstyle{\pm 0.046}$ $0.634\scriptstyle{\pm 0.055}$ $0.629\scriptstyle{\pm 0.040}$ $0.624\scriptstyle{\pm 0.038}$ $0.618\scriptstyle{\pm 0.040}$ $0.527\scriptstyle{\pm 0.077}$ 20 $0.376\scriptstyle{\pm 0.008}$ $0.608\scriptstyle{\pm 0.027}$ $0.570\scriptstyle{\pm 0.065}$ $0.608\scriptstyle{\pm 0.020}$ $0.597\scriptstyle{\pm 0.026}$ $0.608\scriptstyle{\pm 0.020}$ $0.554\scriptstyle{\pm 0.020}$ $0.548\scriptstyle{\pm 0.115}$
 Table 8: Accuracy per ensemble member of baselines on ChaosNLI, CIFAR-10, and QualityMRI.
-ChaosNLI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.658\scriptstyle{\pm 0.013}$ $0.671\scriptstyle{\pm 0.011}$ $0.512\scriptstyle{\pm 0.083}$ $0.525\scriptstyle{\pm 0.051}$ 2 $0.648\scriptstyle{\pm 0.012}$ $0.676\scriptstyle{\pm 0.013}$ $0.453\scriptstyle{\pm 0.010}$ $0.488\scriptstyle{\pm 0.098}$ 3 $0.641\scriptstyle{\pm 0.020}$ $0.676\scriptstyle{\pm 0.013}$ $0.506\scriptstyle{\pm 0.085}$ $0.542\scriptstyle{\pm 0.030}$ 4 $0.636\scriptstyle{\pm 0.012}$ $0.666\scriptstyle{\pm 0.008}$ $0.511\scriptstyle{\pm 0.092}$ $0.485\scriptstyle{\pm 0.037}$ 5 $0.651\scriptstyle{\pm 0.032}$ $0.667\scriptstyle{\pm 0.011}$ $0.532\scriptstyle{\pm 0.111}$ $0.554\scriptstyle{\pm 0.051}$ 6 $0.660\scriptstyle{\pm 0.016}$ $0.674\scriptstyle{\pm 0.021}$ $0.503\scriptstyle{\pm 0.081}$ $0.470\scriptstyle{\pm 0.115}$ 7 $0.637\scriptstyle{\pm 0.014}$ $0.677\scriptstyle{\pm 0.009}$ $0.453\scriptstyle{\pm 0.010}$ $0.518\scriptstyle{\pm 0.036}$ 8 $0.667\scriptstyle{\pm 0.016}$ $0.670\scriptstyle{\pm 0.010}$ $0.453\scriptstyle{\pm 0.010}$ $0.446\scriptstyle{\pm 0.030}$ 9 $0.584\scriptstyle{\pm 0.086}$ $0.663\scriptstyle{\pm 0.005}$ $0.453\scriptstyle{\pm 0.010}$ $0.462\scriptstyle{\pm 0.113}$ 10 $0.663\scriptstyle{\pm 0.035}$ $0.674\scriptstyle{\pm 0.029}$ $0.664\scriptstyle{\pm 0.012}$ $0.531\scriptstyle{\pm 0.064}$ 11 $0.648\scriptstyle{\pm 0.017}$ $0.653\scriptstyle{\pm 0.015}$ $0.591\scriptstyle{\pm 0.099}$ $0.484\scriptstyle{\pm 0.047}$ 12 $0.660\scriptstyle{\pm 0.016}$ $0.662\scriptstyle{\pm 0.016}$ $0.587\scriptstyle{\pm 0.097}$ $0.549\scriptstyle{\pm 0.071}$ 13 $0.664\scriptstyle{\pm 0.014}$ $0.655\scriptstyle{\pm 0.017}$ $0.581\scriptstyle{\pm 0.099}$ $0.487\scriptstyle{\pm 0.040}$ 14 $0.650\scriptstyle{\pm 0.011}$ $0.654\scriptstyle{\pm 0.006}$ $0.655\scriptstyle{\pm 0.007}$ $0.505\scriptstyle{\pm 0.039}$ 15 $0.635\scriptstyle{\pm 0.038}$ $0.651\scriptstyle{\pm 0.018}$ $0.602\scriptstyle{\pm 0.097}$ $0.458\scriptstyle{\pm 0.143}$ 16 $0.652\scriptstyle{\pm 0.026}$ $0.654\scriptstyle{\pm 0.007}$ $0.629\scriptstyle{\pm 0.046}$ $0.486\scriptstyle{\pm 0.012}$ 17 $0.628\scriptstyle{\pm 0.019}$ $0.633\scriptstyle{\pm 0.015}$ $0.587\scriptstyle{\pm 0.086}$ $0.518\scriptstyle{\pm 0.057}$ 18 $0.652\scriptstyle{\pm 0.009}$ $0.605\scriptstyle{\pm 0.021}$ $0.502\scriptstyle{\pm 0.079}$ $0.519\scriptstyle{\pm 0.031}$ 19 $0.660\scriptstyle{\pm 0.012}$ $0.581\scriptstyle{\pm 0.037}$ $0.584\scriptstyle{\pm 0.095}$ $0.487\scriptstyle{\pm 0.050}$ 20 $0.651\scriptstyle{\pm 0.013}$ $0.480\scriptstyle{\pm 0.074}$ $0.586\scriptstyle{\pm 0.095}$ $0.560\scriptstyle{\pm 0.043}$ CIFAR-10 CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.942\scriptstyle{\pm 0.001}$ $0.955\scriptstyle{\pm 0.000}$ $0.875\scriptstyle{\pm 0.004}$ $0.941\scriptstyle{\pm 0.000}$ 2 $0.944\scriptstyle{\pm 0.001}$ $0.954\scriptstyle{\pm 0.000}$ $0.877\scriptstyle{\pm 0.005}$ $0.943\scriptstyle{\pm 0.000}$ 3 $0.944\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.867\scriptstyle{\pm 0.015}$ $0.944\scriptstyle{\pm 0.001}$ 4 $0.943\scriptstyle{\pm 0.002}$ $0.952\scriptstyle{\pm 0.000}$ $0.881\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.001}$ 5 $0.943\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.004}$ $0.942\scriptstyle{\pm 0.001}$ 6 $0.943\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.000}$ $0.869\scriptstyle{\pm 0.006}$ $0.942\scriptstyle{\pm 0.002}$ 7 $0.941\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.877\scriptstyle{\pm 0.001}$ $0.943\scriptstyle{\pm 0.001}$ 8 $0.942\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.001}$ $0.880\scriptstyle{\pm 0.006}$ $0.943\scriptstyle{\pm 0.002}$ 9 $0.943\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.005}$ $0.944\scriptstyle{\pm 0.001}$ 10 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.873\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.003}$ 11 $0.942\scriptstyle{\pm 0.001}$ $0.953\scriptstyle{\pm 0.001}$ $0.882\scriptstyle{\pm 0.004}$ $0.943\scriptstyle{\pm 0.000}$ 12 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.000}$ 13 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.001}$ $0.874\scriptstyle{\pm 0.000}$ $0.943\scriptstyle{\pm 0.002}$ 14 $0.945\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.006}$ $0.941\scriptstyle{\pm 0.001}$ 15 $0.942\scriptstyle{\pm 0.002}$ $0.948\scriptstyle{\pm 0.000}$ $0.856\scriptstyle{\pm 0.019}$ $0.942\scriptstyle{\pm 0.001}$ 16 $0.942\scriptstyle{\pm 0.000}$ $0.941\scriptstyle{\pm 0.000}$ $0.870\scriptstyle{\pm 0.002}$ $0.942\scriptstyle{\pm 0.001}$ 17 $0.942\scriptstyle{\pm 0.001}$ $0.930\scriptstyle{\pm 0.001}$ $0.854\scriptstyle{\pm 0.035}$ $0.942\scriptstyle{\pm 0.002}$ 18 $0.944\scriptstyle{\pm 0.001}$ $0.921\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.002}$ 19 $0.943\scriptstyle{\pm 0.001}$ $0.906\scriptstyle{\pm 0.002}$ $0.873\scriptstyle{\pm 0.001}$ $0.942\scriptstyle{\pm 0.001}$ 20 $0.942\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.002}$ $0.860\scriptstyle{\pm 0.012}$ $0.942\scriptstyle{\pm 0.001}$ QualityMRI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.457\scriptstyle{\pm 0.099}$ $0.430\scriptstyle{\pm 0.170}$ $0.548\scriptstyle{\pm 0.115}$ $0.581\scriptstyle{\pm 0.137}$ 2 $0.452\scriptstyle{\pm 0.139}$ $0.409\scriptstyle{\pm 0.163}$ $0.484\scriptstyle{\pm 0.115}$ $0.559\scriptstyle{\pm 0.112}$ 3 $0.425\scriptstyle{\pm 0.124}$ $0.419\scriptstyle{\pm 0.160}$ $0.468\scriptstyle{\pm 0.126}$ $0.570\scriptstyle{\pm 0.110}$ 4 $0.419\scriptstyle{\pm 0.126}$ $0.446\scriptstyle{\pm 0.187}$ $0.548\scriptstyle{\pm 0.115}$ $0.575\scriptstyle{\pm 0.102}$ 5 $0.398\scriptstyle{\pm 0.135}$ $0.387\scriptstyle{\pm 0.139}$ $0.468\scriptstyle{\pm 0.126}$ $0.548\scriptstyle{\pm 0.126}$ 6 $0.403\scriptstyle{\pm 0.117}$ $0.398\scriptstyle{\pm 0.158}$ $0.554\scriptstyle{\pm 0.118}$ $0.554\scriptstyle{\pm 0.084}$ 7 $0.409\scriptstyle{\pm 0.175}$ $0.414\scriptstyle{\pm 0.161}$ $0.446\scriptstyle{\pm 0.107}$ $0.554\scriptstyle{\pm 0.095}$ 8 $0.430\scriptstyle{\pm 0.119}$ $0.441\scriptstyle{\pm 0.182}$ $0.554\scriptstyle{\pm 0.118}$ $0.559\scriptstyle{\pm 0.133}$ 9 $0.392\scriptstyle{\pm 0.137}$ $0.419\scriptstyle{\pm 0.181}$ $0.473\scriptstyle{\pm 0.110}$ $0.586\scriptstyle{\pm 0.141}$ 10 $0.446\scriptstyle{\pm 0.177}$ $0.430\scriptstyle{\pm 0.140}$ $0.462\scriptstyle{\pm 0.130}$ $0.570\scriptstyle{\pm 0.118}$ 11 $0.425\scriptstyle{\pm 0.137}$ $0.414\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.602\scriptstyle{\pm 0.107}$ 12 $0.398\scriptstyle{\pm 0.122}$ $0.430\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.103}$ 13 $0.398\scriptstyle{\pm 0.130}$ $0.430\scriptstyle{\pm 0.140}$ $0.446\scriptstyle{\pm 0.107}$ $0.575\scriptstyle{\pm 0.090}$ 14 $0.409\scriptstyle{\pm 0.066}$ $0.430\scriptstyle{\pm 0.142}$ $0.457\scriptstyle{\pm 0.122}$ $0.581\scriptstyle{\pm 0.070}$ 15 $0.409\scriptstyle{\pm 0.146}$ $0.414\scriptstyle{\pm 0.132}$ $0.468\scriptstyle{\pm 0.126}$ $0.624\scriptstyle{\pm 0.133}$ 16 $0.414\scriptstyle{\pm 0.107}$ $0.425\scriptstyle{\pm 0.167}$ $0.554\scriptstyle{\pm 0.118}$ $0.608\scriptstyle{\pm 0.100}$ 17 $0.419\scriptstyle{\pm 0.139}$ $0.414\scriptstyle{\pm 0.119}$ $0.554\scriptstyle{\pm 0.118}$ $0.581\scriptstyle{\pm 0.091}$ 18 $0.403\scriptstyle{\pm 0.168}$ $0.414\scriptstyle{\pm 0.129}$ $0.468\scriptstyle{\pm 0.126}$ $0.581\scriptstyle{\pm 0.117}$ 19 $0.425\scriptstyle{\pm 0.151}$ $0.414\scriptstyle{\pm 0.073}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.142}$ 20 $0.430\scriptstyle{\pm 0.154}$ $0.446\scriptstyle{\pm 0.068}$ $0.548\scriptstyle{\pm 0.117}$ $0.581\scriptstyle{\pm 0.103}$
+ChaosNLI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.658\scriptstyle{\pm 0.013}$ $0.671\scriptstyle{\pm 0.011}$ $0.680\scriptstyle{\pm 0.013}$ $0.525\scriptstyle{\pm 0.051}$ 2 $0.648\scriptstyle{\pm 0.012}$ $0.676\scriptstyle{\pm 0.013}$ $0.674\scriptstyle{\pm 0.031}$ $0.488\scriptstyle{\pm 0.098}$ 3 $0.641\scriptstyle{\pm 0.020}$ $0.676\scriptstyle{\pm 0.013}$ $0.517\scriptstyle{\pm 0.079}$ $0.542\scriptstyle{\pm 0.030}$ 4 $0.636\scriptstyle{\pm 0.012}$ $0.666\scriptstyle{\pm 0.008}$ $0.683\scriptstyle{\pm 0.030}$ $0.485\scriptstyle{\pm 0.037}$ 5 $0.651\scriptstyle{\pm 0.032}$ $0.667\scriptstyle{\pm 0.011}$ $0.533\scriptstyle{\pm 0.112}$ $0.554\scriptstyle{\pm 0.051}$ 6 $0.660\scriptstyle{\pm 0.016}$ $0.674\scriptstyle{\pm 0.021}$ $0.667\scriptstyle{\pm 0.008}$ $0.470\scriptstyle{\pm 0.115}$ 7 $0.637\scriptstyle{\pm 0.014}$ $0.677\scriptstyle{\pm 0.009}$ $0.675\scriptstyle{\pm 0.012}$ $0.518\scriptstyle{\pm 0.036}$ 8 $0.667\scriptstyle{\pm 0.016}$ $0.670\scriptstyle{\pm 0.010}$ $0.675\scriptstyle{\pm 0.029}$ $0.446\scriptstyle{\pm 0.030}$ 9 $0.584\scriptstyle{\pm 0.086}$ $0.663\scriptstyle{\pm 0.005}$ $0.610\scriptstyle{\pm 0.104}$ $0.462\scriptstyle{\pm 0.113}$ 10 $0.663\scriptstyle{\pm 0.035}$ $0.674\scriptstyle{\pm 0.029}$ $0.668\scriptstyle{\pm 0.017}$ $0.531\scriptstyle{\pm 0.064}$ 11 $0.648\scriptstyle{\pm 0.017}$ $0.653\scriptstyle{\pm 0.015}$ $0.529\scriptstyle{\pm 0.096}$ $0.484\scriptstyle{\pm 0.047}$ 12 $0.660\scriptstyle{\pm 0.016}$ $0.662\scriptstyle{\pm 0.016}$ $0.684\scriptstyle{\pm 0.019}$ $0.549\scriptstyle{\pm 0.071}$ 13 $0.664\scriptstyle{\pm 0.014}$ $0.655\scriptstyle{\pm 0.017}$ $0.526\scriptstyle{\pm 0.105}$ $0.487\scriptstyle{\pm 0.040}$ 14 $0.650\scriptstyle{\pm 0.011}$ $0.654\scriptstyle{\pm 0.006}$ $0.668\scriptstyle{\pm 0.030}$ $0.505\scriptstyle{\pm 0.039}$ 15 $0.635\scriptstyle{\pm 0.038}$ $0.651\scriptstyle{\pm 0.018}$ $0.673\scriptstyle{\pm 0.009}$ $0.458\scriptstyle{\pm 0.143}$ 16 $0.652\scriptstyle{\pm 0.026}$ $0.654\scriptstyle{\pm 0.007}$ $0.616\scriptstyle{\pm 0.108}$ $0.486\scriptstyle{\pm 0.012}$ 17 $0.628\scriptstyle{\pm 0.019}$ $0.633\scriptstyle{\pm 0.015}$ $0.675\scriptstyle{\pm 0.030}$ $0.518\scriptstyle{\pm 0.057}$ 18 $0.652\scriptstyle{\pm 0.009}$ $0.605\scriptstyle{\pm 0.021}$ $0.598\scriptstyle{\pm 0.109}$ $0.519\scriptstyle{\pm 0.031}$ 19 $0.660\scriptstyle{\pm 0.012}$ $0.581\scriptstyle{\pm 0.037}$ $0.628\scriptstyle{\pm 0.116}$ $0.487\scriptstyle{\pm 0.050}$ 20 $0.651\scriptstyle{\pm 0.013}$ $0.480\scriptstyle{\pm 0.074}$ $0.536\scriptstyle{\pm 0.116}$ $0.560\scriptstyle{\pm 0.043}$ CIFAR-10 CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.942\scriptstyle{\pm 0.001}$ $0.955\scriptstyle{\pm 0.000}$ $0.875\scriptstyle{\pm 0.004}$ $0.941\scriptstyle{\pm 0.000}$ 2 $0.944\scriptstyle{\pm 0.001}$ $0.954\scriptstyle{\pm 0.000}$ $0.877\scriptstyle{\pm 0.005}$ $0.943\scriptstyle{\pm 0.000}$ 3 $0.944\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.867\scriptstyle{\pm 0.015}$ $0.944\scriptstyle{\pm 0.001}$ 4 $0.943\scriptstyle{\pm 0.002}$ $0.952\scriptstyle{\pm 0.000}$ $0.881\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.001}$ 5 $0.943\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.004}$ $0.942\scriptstyle{\pm 0.001}$ 6 $0.943\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.000}$ $0.869\scriptstyle{\pm 0.006}$ $0.942\scriptstyle{\pm 0.002}$ 7 $0.941\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.877\scriptstyle{\pm 0.001}$ $0.943\scriptstyle{\pm 0.001}$ 8 $0.942\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.001}$ $0.880\scriptstyle{\pm 0.006}$ $0.943\scriptstyle{\pm 0.002}$ 9 $0.943\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.005}$ $0.944\scriptstyle{\pm 0.001}$ 10 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.873\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.003}$ 11 $0.942\scriptstyle{\pm 0.001}$ $0.953\scriptstyle{\pm 0.001}$ $0.882\scriptstyle{\pm 0.004}$ $0.943\scriptstyle{\pm 0.000}$ 12 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.000}$ 13 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.001}$ $0.874\scriptstyle{\pm 0.000}$ $0.943\scriptstyle{\pm 0.002}$ 14 $0.945\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.006}$ $0.941\scriptstyle{\pm 0.001}$ 15 $0.942\scriptstyle{\pm 0.002}$ $0.948\scriptstyle{\pm 0.000}$ $0.856\scriptstyle{\pm 0.019}$ $0.942\scriptstyle{\pm 0.001}$ 16 $0.942\scriptstyle{\pm 0.000}$ $0.941\scriptstyle{\pm 0.000}$ $0.870\scriptstyle{\pm 0.002}$ $0.942\scriptstyle{\pm 0.001}$ 17 $0.942\scriptstyle{\pm 0.001}$ $0.930\scriptstyle{\pm 0.001}$ $0.854\scriptstyle{\pm 0.035}$ $0.942\scriptstyle{\pm 0.002}$ 18 $0.944\scriptstyle{\pm 0.001}$ $0.921\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.002}$ 19 $0.943\scriptstyle{\pm 0.001}$ $0.906\scriptstyle{\pm 0.002}$ $0.873\scriptstyle{\pm 0.001}$ $0.942\scriptstyle{\pm 0.001}$ 20 $0.942\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.002}$ $0.860\scriptstyle{\pm 0.012}$ $0.942\scriptstyle{\pm 0.001}$ QualityMRI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.457\scriptstyle{\pm 0.099}$ $0.430\scriptstyle{\pm 0.170}$ $0.548\scriptstyle{\pm 0.115}$ $0.581\scriptstyle{\pm 0.137}$ 2 $0.452\scriptstyle{\pm 0.139}$ $0.409\scriptstyle{\pm 0.163}$ $0.484\scriptstyle{\pm 0.115}$ $0.559\scriptstyle{\pm 0.112}$ 3 $0.425\scriptstyle{\pm 0.124}$ $0.419\scriptstyle{\pm 0.160}$ $0.468\scriptstyle{\pm 0.126}$ $0.570\scriptstyle{\pm 0.110}$ 4 $0.419\scriptstyle{\pm 0.126}$ $0.446\scriptstyle{\pm 0.187}$ $0.548\scriptstyle{\pm 0.115}$ $0.575\scriptstyle{\pm 0.102}$ 5 $0.398\scriptstyle{\pm 0.135}$ $0.387\scriptstyle{\pm 0.139}$ $0.468\scriptstyle{\pm 0.126}$ $0.548\scriptstyle{\pm 0.126}$ 6 $0.403\scriptstyle{\pm 0.117}$ $0.398\scriptstyle{\pm 0.158}$ $0.554\scriptstyle{\pm 0.118}$ $0.554\scriptstyle{\pm 0.084}$ 7 $0.409\scriptstyle{\pm 0.175}$ $0.414\scriptstyle{\pm 0.161}$ $0.446\scriptstyle{\pm 0.107}$ $0.554\scriptstyle{\pm 0.095}$ 8 $0.430\scriptstyle{\pm 0.119}$ $0.441\scriptstyle{\pm 0.182}$ $0.554\scriptstyle{\pm 0.118}$ $0.559\scriptstyle{\pm 0.133}$ 9 $0.392\scriptstyle{\pm 0.137}$ $0.419\scriptstyle{\pm 0.181}$ $0.473\scriptstyle{\pm 0.110}$ $0.586\scriptstyle{\pm 0.141}$ 10 $0.446\scriptstyle{\pm 0.177}$ $0.430\scriptstyle{\pm 0.140}$ $0.462\scriptstyle{\pm 0.130}$ $0.570\scriptstyle{\pm 0.118}$ 11 $0.425\scriptstyle{\pm 0.137}$ $0.414\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.602\scriptstyle{\pm 0.107}$ 12 $0.398\scriptstyle{\pm 0.122}$ $0.430\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.103}$ 13 $0.398\scriptstyle{\pm 0.130}$ $0.430\scriptstyle{\pm 0.140}$ $0.446\scriptstyle{\pm 0.107}$ $0.575\scriptstyle{\pm 0.090}$ 14 $0.409\scriptstyle{\pm 0.066}$ $0.430\scriptstyle{\pm 0.142}$ $0.457\scriptstyle{\pm 0.122}$ $0.581\scriptstyle{\pm 0.070}$ 15 $0.409\scriptstyle{\pm 0.146}$ $0.414\scriptstyle{\pm 0.132}$ $0.468\scriptstyle{\pm 0.126}$ $0.624\scriptstyle{\pm 0.133}$ 16 $0.414\scriptstyle{\pm 0.107}$ $0.425\scriptstyle{\pm 0.167}$ $0.554\scriptstyle{\pm 0.118}$ $0.608\scriptstyle{\pm 0.100}$ 17 $9\scriptstyle{\pm 0.139} 0.414±0.1$$4\scriptstyle{\pm 0.119}$ $0.554\scriptstyle{\pm 0.118}$ $0.581\scriptstyle{\pm 0.091}$ 18 $0.403\scriptstyle{\pm 0.168}$ $0.414\scriptstyle{\pm 0.129}$ $0.468\scriptstyle{\pm 0.126}$ $0.581\scriptstyle{\pm 0.117}$ 19 $0.425\scriptstyle{\pm 0.151}$ $0.414\scriptstyle{\pm 0.073}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.142}$ 20 $0.430\scriptstyle{\pm 0.154}$ $0.446\scriptstyle{\pm 0.068}$ $0.548\scriptstyle{\pm 0.117}$ $0.581\scriptstyle{\pm 0.103}$